### v2

# The Obstacle Problem Arising from the American Chooser Option

Abstract.
We study the obstacle problem associated with the American chooser option. The obstacle is given by the maximum of an American call option and an American put option, which, in turn, can be expressed as the maximum of the solutions to the corresponding obstacle problems. This structure makes the obstacle problem particularly challenging and non-trivial. Using theoretical analysis, we overcome these difficulties and establish the existence and uniqueness of a strong solution. Furthermore, we rigorously prove the monotonicity and smoothness of the free boundary arising from the obstacle problem.
Key words and phrases:
American chooser option, Free boundary problem, Optimal stopping, Parabolic obstacle problem
2020 Mathematics Subject Classification:
35R35, 35K85, 60G40, 91G80
1. Introduction
In this paper, our primary objective is to analyze the following (lower) obstacle problem derived from the American chooser option:
(1.1) $\begin{cases}\partial_{t}V^{\rm ch}(t,s)+\mathfrak{L}V^{\rm ch}(t,s)\leqslant 0\quad\text{for }\ (t,s)\in D_{T}\ \text{ with }\ V^{\rm ch}(t,s)=\max\{C^{A}(t,s),P^{A}(t,s)\},\\ \partial_{t}V^{\rm ch}(t,s)+\mathfrak{L}V^{\rm ch}(t,s)=0\quad\text{for }\ (t,s)\in D_{T}\ \text{ with }\ V^{\rm ch}(t,s)>\max\{C^{A}(t,s),P^{A}(t,s)\},\\ V^{\rm ch}(T,s)=\max\{C^{A}(T,s),P^{A}(T,s)\}\quad\text{for }\ 0<s<\infty.\end{cases}$
where the differential operator $\mathfrak{L}$ is defined as
$\mathfrak{L}:=\frac{\sigma^{2}}{2}s^{2}\partial_{ss}+(r-q)s\partial_{s}-r,\quad\text{with}\quad r,\;\sigma>0,\;q\geqslant 0,$
[…]
(1.2) $\begin{cases}\partial_{t}C^{A}(t,s)+\mathfrak{L}C^{A}(t,s)\leqslant 0\quad\text{for }\ (t,s)\in D_{T_{c}}\ \text{ with }\ C^{A}(t,s)=(s-K_{c})^{+},\\ \partial_{t}C^{A}(t,s)+\mathfrak{L}C^{A}(t,s)=0\quad\text{for }\ (t,s)\in D_{T_{c}}\ \text{ with }\ C^{A}(t,s)>(s-K_{c})^{+},\\ C^{A}(T,s)=(s-K_{c})^{+}\quad\text{for }\ 0<s<\infty.\end{cases}$
(1.3) $\begin{cases}\partial_{t}P^{A}(t,s)+\mathfrak{L}P^{A}(t,s)\leqslant 0\quad\text{for }\ (t,s)\in D_{T_{p}}\ \text{ with }\ P^{A}(t,s)=(K_{p}-s)^{+},\\ \partial_{t}P^{A}(t,s)+\mathfrak{L}P^{A}(t,s)=0\quad\text{for }\ (t,s)\in D_{T_{p}}\ \text{ with }\ P^{A}(t,s)>(K_{p}-s)^{+},\\ P^{A}(T,s)=(K_{p}-s)^{+}\quad\text{for }\ 0<s<\infty.\end{cases}$
To ensure well-posedness, we assume:
$0<T<\min\{T_{c},T_{p}\},\quad 0<K_{p}<K_{c}.$
The existence and uniqueness of strong solutions to (1.2) and (1.3) are well-established, as shown in [27, 15].
The American chooser option allows the holder to exercise it as either an American call or an American put option, depending on which is more valuable at the time of exercise. This flexibility makes it particularly useful in uncertain market conditions where large price movements are anticipated. Consequently, it shares similarities with the American strangle option and includes it as a special case.
From a mathematical perspective, pricing an American chooser option leads to a parabolic obstacle problem, where the obstacle function is determined as the maximum of two separate solutions—one corresponding to the American call and the other to the American put. Unlike standard American options, in which the obstacle function is explicitly defined and well-behaved, the American chooser option introduces additional complexity due to the implicit and non-standard nature of the obstacle. This fundamental distinction makes the problem significantly more challenging to analyze within a mathematical framework.
Standard American options have been extensively studied in the literature (e.g., [24, 26, 28, 15]). Most of these studies however rely on obstacle problems where the obstacle function has an explicit, closed-form representation. In contrast, the American chooser option involves an obstacle function that depends on the interaction between two separate obstacle problems. This interaction gives rise to a new level of mathematical difficulty, as the obstacle is no longer explicitly known but rather emerges as the maximum of two distinct solutions.
Earlier studies on the American chooser option, such as [9, 22], adopted probabilistic approaches. In contrast, we present a rigorous analysis within a PDE framework, which, to the best of our knowledge, has not been previously explored. Our work focuses on the associated obstacle problem and develops a detailed PDE-based methodology. To address the challenges mentioned above, we carefully analyze the structure and behavior of the solutions to the two underlying obstacle problems, and as a result, we are able to construct a framework to study the American chooser option.
Specifically, following the methodology in [15], we introduce a penalized problem to approximate the original obstacle problem. Given the irregular and implicit nature of the obstacle function in this setting, a regularized version of the problem is considered. By leveraging the properties of the obstacle function and employing advanced PDE techniques, including the comparison principle, we demonstrate that the penalized problem yields a uniformly bounded solution. Importantly, while the obstacle in [15] is explicitly defined, the obstacle in our case depends on two interacting obstacle problems. Despite this complexity, we show that strong solutions can still be constructed, enabling the application of the comparison principle and rigorous analysis.
A crucial aspect of our analysis is the structure of the free boundaries. Due to the nature of the American chooser option, the associated obstacle problem features two time-dependent free boundaries. Mathematically, we observe that the obstacle’s components align with solutions in separate, disjoint regions, as in [15]. However, unlike [15], the exact values of the obstacle at each point in the domain are unknown. The key insight is that these regions lie within the respective exercise regions of the American put and call options. This guarantees the existence of two distinct free boundaries provided that the exercise region of the American chooser option is nonempty. We further establish their monotonicity and smoothness, which are crucial for understanding the solution’s behavior.
Our paper is structured as follows. Section 2 provides the financial background and formal definition of the American chooser option and demonstrates the properties of the Amercian call and put options. Section 3 establishes the existence and uniqueness of the obstacle problem solution using the penalty method. Section 4 examines the monotonicity and smoothness of the two free boundaries that arise in this framework.
2. Preliminaries
2.1. Notations
Throughout the paper, for each $\epsilon>0$, $\varphi_{\epsilon}$ is a smooth function in $\mathbb{R}$ satisfying that
(2.1) $\begin{cases}\varphi_{\varepsilon}\geqslant 0,\quad 0\leqslant\varphi^{\prime}_{\varepsilon}\leqslant 1,\quad\varphi^{\prime\prime}_{\varepsilon}\geqslant 0,\\ \varphi_{\varepsilon}(t)=t\ \ \text{if}\ \ t\geqslant\varepsilon,\quad\varphi_{\varepsilon}(t)=0\ \ \text{if}\ \ t\leqslant-\varepsilon,\\ \varphi_{\varepsilon}(t)\leqslant(t+\varepsilon)^{+}\ \ \text{for}\ \ t\in\mathbb{R},\\ \displaystyle{\lim_{\varepsilon\to 0}\varphi_{\varepsilon}(t)}=t^{+}\ \ \text{uniformly for }\ t\in\mathbb{R}.\end{cases}$
Let $\Omega_{T}^{n}:=(0,T)\times(-n,n)$.
  • $C^{k+\frac{\alpha}{2},2k+\alpha}(\Omega_{T}^{n})$, $\alpha\in(0,1)$, $k\in\mathbb{N}$ is the Banach space under the following norm for $V$:
    $\begin{split}&\|V\|_{C^{k+\frac{\alpha}{2},2k+\alpha}(\Omega_{T}^{n})}\\ &\quad:=\sum_{i+2j=0,2,\cdots,2k}\limits\sup_{(t_{1},x_{1}),(t_{2},x_{2})\in\Omega_{T}^{n}}\{|D^{i}_{x}D^{j}_{t}V(t_{1},x_{1})|+\frac{|D^{i}_{x}D^{j}_{t}V(t_{1},x_{1})-D^{i}_{x}D^{j}_{t}V(t_{2},x_{2})|}{{|t_{1}-t_{2}|}^{\frac{\alpha}{2}}+{|x_{1}-x_{2}|}^{\alpha}}\}.\end{split}$
  • $L^{p}(\Omega_{T}^{n}),p\geqslant 1$ is the completion of $C^{\infty}(\Omega_{T}^{n})$ under the following norm for $V$:
[…]
  • $W^{1,2}_{p}(\Omega_{T}^{n}),$ $p\geqslant 1$ is the completion of $C^{\infty}(\Omega_{T}^{n})$ under the following norm for $V$:
[…]
Let ${G}$ be a parabolic domain, $(\tau_{0},x_{0})\in G$ and $\varepsilon>0$.
  • $Q((\tau_{0},x_{0}),\varepsilon)$ is the cylinder such that
    $Q((\tau_{0},x_{0}),\varepsilon):=\{(\tau,x)\in\Omega_{T}:\text{max}\{|x-x_{0}|,|\tau-\tau_{0}|^{\frac{1}{2}}\}<\varepsilon,\tau<\tau_{0}\}.$
  • $\mathcal{P}{G}$ is the parabolic boundary of ${G}$ which is defined to be the set of all points $(\tau_{0},x_{0})\in\partial{G}$ such that for any $\varepsilon>0,$ the cylinder $Q((\tau_{0},x_{0}),\varepsilon)$ contains points not in ${G}$.
2.2. Financial Background: American Chooser Option
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a filtered probability space with a filtration $\{\mathcal{F}_{t}\}_{t\geqslant 0}$ satisfying the usual conditions. Under the risk-neutral measure $\mathbb{Q}$, the stock price $S_{t}$ follows a geometric Brownian motion (GBM) given by
$dS_{t}=(r-q)S_{t}dt+\sigma S_{t}dW_{t}^{\mathbb{Q}},\quad S_{0}>0,$
where $r>0$ is the constant risk-free interest rate, $q\geqslant 0$ is the continuous dividend rate, $\sigma>0$ is the volatility of the stock, and $W_{t}^{\mathbb{Q}}$ is a standard Brownian motion under the risk-neutral measure $\mathbb{Q}$. Throughout this paper, we assume $q>0$ for simplicity. The case $q=0$ leads to a degenerate situation with only one free boundary, but similar results can be obtained (see [15]).
To define the American chooser option, we first introduce the American call and put options. These two options serve as the fundamental building blocks for the American chooser option. An American call option with strike price $K_{c}$ and expiration time $T_{c}$ gives the holder the right to buy the underlying asset at price $K_{c}$ at any time $\tau\in\mathcal{U}_{t,T_{c}}$, where $\mathcal{U}_{t,T}$ is the set of all $\mathcal{F}$-stopping times taking values in $[t,T]$. The value $C^{A}(t,S_{t})$ of the American call option at time $t$ is given by
(2.2) $C^{A}(t,S_{t})=\sup_{\tau\in\mathcal{U}_{t,T_{c}}}\mathbb{E}^{\mathbb{Q}}\left[e^{-r(\tau-t)}(S_{\tau}-K_{c})^{+}\mid\mathcal{F}_{t}\right].$
Similarly, an American put option with the strike price $K_{p}<K_{c}$ and expiration time $T_{p}$ grants the holder the right to sell the underlying asset at price $K_{p}$ at any time $\tau\in[0,T_{p}]$. The value $P^{A}(t,S_{t})$ of the American put option at time $t$ is given by
(2.3) $P^{A}(t,S_{t})=\sup_{\tau\in\mathcal{U}_{t,T_{p}}}\mathbb{E}^{\mathbb{Q}}\left[e^{-r(\tau-t)}(K_{p}-S_{\tau})^{+}\mid\mathcal{F}_{t}\right].$
If $C^{A}$ and $P^{A}$ are strong solutions to the obstacle problems (1.2) and (1.3), respectively, then one can show, by applying Itô’s lemma for Sobolev space (see [17]), that they solve the corresponding optimal stopping problems (2.2) and (2.3); see, for instance, [3, Appendix B].
Under these circumstances, the value of the American chooser option, denoted by $V^{\rm ch}$, is defined as the solution to the following optimal stopping problem:
(2.4) $V^{\rm ch}(t,S_{t})=\sup_{\tau\in\mathcal{U}_{t,T}}\mathbb{E}\left[e^{-r(\tau-t)}\max\left\{C^{A}(\tau,S_{\tau}),P^{A}(\tau,S_{\tau})\right\}\mid\mathcal{F}_{t}\right].$
Using dynamic programming and Itô’s lemma, we can easily derive the obstacle problem (1.1) from (2.4).
2.3. Change of variables
Since the obstacle problems in (1.2) and (1.3) are backward degenerate problems, we transform them into forward non-degenerate problems and denote the solutions by $\widehat{C}^{A}$ and $\widehat{P}^{A}$, respectively. More precisely, we introduce the following transformation:
$\widehat{C}^{A}(\zeta,x)=C^{A}(T_{c}-\zeta,e^{x})\quad\text{and}\quad\widehat{P}^{A}(\zeta,x)=P^{A}(T_{p}-\zeta,e^{x}).$
Moreover, we define the domain $\Omega_{\eta}$ for a given constant $\eta>0$ as
$\Omega_{\eta}:=\{(\zeta,x)\mid 0<\zeta<\eta,\;x\in\mathbb{R}\}$
Then, $\widehat{C}^{A}$ and $\widehat{P}^{A}$ satisfy the following obstacle problems, respectively:
(2.5) $\begin{cases}\partial_{\zeta}\widehat{C}^{A}(\zeta,x)-\mathcal{L}\widehat{C}^{A}(\zeta,x)\geqslant 0\quad\text{for }\ (\zeta,x)\in\Omega_{T_{c}}\ \text{ with }\ \widehat{C}^{A}(\zeta,x)=(e^{x}-K_{c})^{+},\\ \partial_{\zeta}\widehat{C}^{A}(\zeta,x)-\mathcal{L}\widehat{C}^{A}(\zeta,x)=0\quad\text{for }\ (\zeta,x)\in\Omega_{T_{c}}\ \text{ with }\ \widehat{C}^{A}(\zeta,x)>(e^{x}-K_{c})^{+},\\ \widehat{C}^{A}(0,x)=(e^{x}-K_{c})^{+}\quad\text{for }\ x\in\mathbb{R},\end{cases}$
(2.6) $\begin{cases}\partial_{\zeta}\widehat{P}^{A}(\zeta,x)-\mathcal{L}\widehat{P}^{A}(\zeta,x)\geqslant 0\quad\text{for }\ (\zeta,x)\in\Omega_{T_{p}}\ \text{ with }\ \widehat{P}^{A}(\zeta,x)=(K_{p}-e^{x})^{+},\\ \partial_{\zeta}\widehat{P}^{A}(\zeta,x)-\mathcal{L}\widehat{P}^{A}(\zeta,x)=0\quad\text{for }\ (\zeta,x)\in\Omega_{T_{p}}\ \text{ with }\ \widehat{P}^{A}(\zeta,x)>(K_{p}-e^{x})^{+},\\ \widehat{P}^{A}(0,x)=(K_{p}-e^{x})^{+}\quad\text{for }\ x\in\mathbb{R},\end{cases}$
where the differential operator $\mathcal{L}$ is given by
$\mathcal{L}:=\frac{\sigma^{2}}{2}\partial_{xx}+(r-q-\frac{\sigma^{2}}{2})\partial_{x}-r.$
According to the vast literature on American options (see, for instance, [1, 15, 16, 20, 25, 27]), the obstacle problems (2.5) and (2.6) have unique strong solutions, $\widehat{C}^{A}\in W^{1,2}_{p\textup{,}\,\text{loc}}(\Omega_{T_{c}})\cap C(\overline{\Omega_{T_{c}}})$ and $\widehat{P}^{A}\in W^{1,2}_{p\textup{,}\,\text{loc}}(\Omega_{T_{p}})\cap C(\overline{\Omega_{T_{p}}})$, respectively.
We again perform the change of variables by setting
$V(\tau,x):=V^{\rm ch}(T-\tau,e^{x}).$
Then, $V(\tau,x)$ satisfies the following forward non-degenerate obstacle problems:
(2.7) $\begin{cases}\partial_{\tau}V(\tau,x)-\mathcal{L}V(\tau,x)\geqslant 0\quad\text{for }\ (\tau,x)\in\Omega_{T}\ \text{ with }\ V(\tau,x)=\max\{C(\tau,x),\;P(\tau,x)\},\\ \partial_{\tau}V(\tau,x)-\mathcal{L}V(\tau,x)=0\quad\text{for }\ (\tau,x)\in\Omega_{T}\ \text{ with }\ V(\tau,x)>\max\{C(\tau,x),\;P(\tau,x)\},\\ V(0,x)=\max\{C(0,x),P(0,x)\}\quad\text{for }\ x\in\mathbb{R},\end{cases}$
where
(2.8) $\begin{cases}C(\tau,x):=\widehat{C}^{A}(T_{c}-T+\tau,x)=C^{A}(T-\tau,e^{x}),\\ P(\tau,x):=\widehat{P}^{A}(T_{p}-T+\tau,x)=P^{A}(T-\tau,e^{x}),\end{cases}\quad\tau\in[0,T].$
2.4. Properties of American Call and Put Options
The payoff function of the American chooser option takes the maximum form of the American call option $C^{A}$ and the American put option $P^{A}$. Therefore, the properties of $C^{A}$ and $P^{A}$ are essential for the analysis of the American chooser option. Since the properties of $C^{A}$ and $P^{A}$ naturally extend to $\widehat{C}^{A}$ and $\widehat{P}^{A}$, we briefly summarize the well-known properties of the American call and American put options in terms of $\widehat{C}^{A}$ and $\widehat{P}^{A}$ for convenience.
In each obstacle problem in (2.5) and (2.6), we define the exercise regions, denoted by $\mathcal{E}_{C}$ and $\mathcal{E}_{P}$, and the continuation regions, denoted by $\mathcal{C}_{C}$ and $\mathcal{C}_{P}$, as follows:
$\mathcal{E}_{C}:=\{(\zeta,x)\in\Omega_{T_{c}}\,:\,\widehat{C}^{A}(\zeta,x)=e^{x}-K_{c}\},\quad\mathcal{C}_{C}:=\{(\zeta,x)\in\Omega_{T_{c}}\,:\,\widehat{C}^{A}(\zeta,x)>(e^{x}-K_{c})^{+}\},$
$\mathcal{E}_{P}:=\{(\zeta,x)\in\Omega_{T_{p}}\,:\,\widehat{P}^{A}(\zeta,x)=K_{p}-e^{x}\},\quad\mathcal{C}_{P}:=\{(\zeta,x)\in\Omega_{T_{p}}\,:\,\widehat{P}^{A}(\zeta,x)>(K_{p}-e^{x})^{+}\}.$
Then, the free boundaries $\hat{x}_{c}(\tau)$ for the American call and $\hat{x}_{p}(\tau)$ for the American put are well-defined as follows:
$\hat{x}_{c}(\tau):=\partial\mathcal{E}_{c}=\inf\{x\in\mathbb{R}\,:\,(\tau,x)\in\mathcal{E}_{c}\}\quad\text{and}\quad\hat{x}_{p}(\tau):=\partial\mathcal{E}_{p}=\sup\{x\in\mathbb{R}\,:\,(\tau,x)\in\mathcal{E}_{p}\}.$
It is well known that $\hat{x}_{c}(\tau)$ and $\hat{x}_{p}(\tau)$ satisfy the following properties (see [2, 4, 8, 21, 27]):
  • $\hat{x}_{p}(\tau)$ is a smooth and strictly decreasing function for $\tau\in(0,T_{p}]$.
  • $\hat{x}_{c}(\tau)$ is a smooth and strictly increasing function for $\tau\in(0,T_{c}]$.
  • As $\tau$ approaches 0, the limiting behaviors of $\hat{x}_{c}(\tau)$ and $\hat{x}_{p}(\tau)$ are given by
    $\lim_{\tau\to 0^{+}}\hat{x}_{c}(\tau)=\ln{\left(\max\left\{1,\tfrac{r}{q}\right\}K_{c}\right)}\quad\text{and}\quad\lim_{\tau\to 0^{+}}\hat{x}_{p}(\tau)=\ln{\left(\min\left\{1,\tfrac{r}{q}\right\}K_{p}\right)},$
    where $q>0$. If $q=0$, there does not exist the free boundary $\hat{x}_{c}(\tau)$ and $\lim_{\tau\to 0^{+}}\hat{x}_{p}(\tau)=\ln{K_{p}}$.
Additionally, the following property can also be obtained by [22].
  • There exists a unique $\bar{x}>0$ such that
    (2.9) $\widehat{C}^{A}(T_{c}-T,\bar{x})=\widehat{P}^{A}(T_{p}-T,\bar{x}),\quad\text{or equivalently}\quad C(0,\bar{x})=P(0,\bar{x}).$
According to the properties of the free boundaries of the American call and put options described above, since $K_{c}>K_{p}$, the following holds, regardless of the relationship between $T_{c}$ and $T_{p}$:
$\hat{x}_{c}(\tau)>\hat{x}_{p}(\tau),\quad\quad\forall\;\tau\geqslant 0.$
In other words, $\mathcal{E}_{C}$ and $\mathcal{E}_{P}$ are always disjoint. Figure 1 illustrates the behaviors of the free boundaries $\hat{x}_{c}(\tau)$ and $\hat{x}_{p}(\tau)$.
(a) the free boundary $\hat{x}_{c}(\tau)$
(b) the free boundary $\hat{x}_{p}(\tau)$
Figure 1. The free boundaries $\hat{x}_{c}(\tau)$ and $\hat{x}_{p}(\tau)$.
In terms of $C$ and $P$, we also define $x_{c}(\tau)$ and $x_{p}(\tau)$ as follows:
$x_{c}(\tau):=\hat{x}_{c}(T_{c}-T+\tau),\quad\text{and}\quad x_{p}(\tau):=\hat{x}_{p}(T_{p}-T+\tau).$
For further discussion, it is necessary to verify various properties of ${C}$ and ${P}$. Since ${C}$ and ${P}$ are restricted functions of $\widehat{C}^{A}$ and $\widehat{P}^{A}$, it is natural to analyze $\widehat{C}^{A}$ and $\widehat{P}^{A}$ over their original whole domains and then conclude that the same properties hold for ${C}$ and ${P}$, respectively. The following results are essentially well-known and their proofs are provided in the Appendix.
Lemma 2.1.
Let $\widehat{C}^{A}$ and $\widehat{P}^{A}$ be the solution to (2.5) and (2.6), respectively.
  • (i)
    $20\leqslant\widehat{C}^{A}(\zeta,x)\leqslant e^{x}+$ 2 in $\Omega_{T_{c}}$, and $20\leqslant\widehat{P}^{A}(\zeta,x)\leqslant K_{p}+$ 2 in $\Omega_{T_{p}}$. Hence, by (2.8), we have that for all $(\tau,x)\in\Omega_{T}$,
    $20\leqslant C(\tau,x)\leqslant e^{x}+2\quad\text{and}\quad 0\leqslant P(\tau,x)\leqslant K_{p}+$ 2 .
  • (ii)
    $0\leqslant\partial_{x}\widehat{C}^{A}(\zeta,x)\leqslant e^{x}$ in $\Omega_{T_{c}}$, and $-e^{x}\leqslant\partial_{x}\widehat{P}^{A}(\zeta,x)\leqslant 0$ in $\Omega_{T_{p}}$. Hence, by (2.8), we have that for all $(\tau,x)\in\Omega_{T}$,
    $0\leqslant\partial_{x}C(\tau,x)\leqslant e^{x}\quad\text{and}\quad-e^{x}\leqslant\partial_{x}P(\tau,x)\leqslant 0$
  • (iii)
    $\partial_{\tau}\widehat{C}^{A}\geqslant 0$ in $\Omega_{T_{c}}$ and $\partial_{\tau}\widehat{P}^{A}\geqslant 0$ in $\Omega_{T_{p}}$. Hence, by (2.8), we have that for all $(\tau,x)\in\Omega_{T}$,
    $\partial_{\tau}C(\tau,x)\geqslant 0\quad\text{and}\quad\partial_{\tau}P(\tau,x)\geqslant 0.$
  • (iv)
    For each $\tau\in(0,T]$, there exists $x_{\tau}\in\mathbb{R}$ such that $C(\tau,x)>P(\tau,x)$ for all $x\geqslant x_{\tau}$.
Proof.
See the Appendix. ∎
3. Existence and uniqueness of solution
In this section, we aim to prove the existence and uniqueness of the solution to the obstacle problem (2.7).
3.1. Existence and uniqueness in a bounded region
We first consider the following obstacle problem in the bounded region $\Omega_{T}^{n}:=(0,T)\times(-n,n)$ for each $n\in\mathbb{N}$ with Neumann boundary condition:
(3.1) $\begin{cases}\partial_{\tau}V_{n}(\tau,x)-\mathcal{L}V_{n}(\tau,x)\geqslant 0\quad\text{for}\ \ (\tau,x)\in\Omega_{T}^{n}\ \ \text{with}\ \ V_{n}(\tau,x)=J(\tau,x),\\ \partial_{\tau}V_{n}(\tau,x)-\mathcal{L}V_{n}(\tau,x)=0\quad\text{for}\ \ (\tau,x)\in\Omega_{T}^{n}\ \ \text{with}\ \ V_{n}(\tau,x)>J(\tau,x),\\ \partial_{x}V_{n}(\tau,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}V_{n}(\tau,n)=e^{n}\quad\text{for }\ \tau\in[0,T],\\ V_{n}(0,x)=J(0,x)\quad\text{for }\ x\in(-n,n),\end{cases}$
where
$J(\tau,x):=\max\{C(\tau,x),P(\tau,x)\}\quad\text{ for\ all }\ (\tau,x)\in\Omega_{T}^{n}.$
We prove the existence and uniqueness of the solution to (3.1) in Theorem 3.3. To this end, we use the so called penalty method.
Define a penalty function $\beta_{\varepsilon}(t)\in C^{\infty}(\mathbb{R})$ with $\varepsilon>0$ satisfying
(3.2) $\begin{cases}\ \beta_{\varepsilon}(t)\leqslant 0,\ \beta^{\prime}_{\varepsilon}(t)\geqslant 0\ \ \text{and}\ \ \beta^{\prime\prime}_{\varepsilon}(t)\leqslant 0\ \ \text{for \ all}\ \ t\in\mathbb{R},\\ \ \beta_{\varepsilon}(t)=0\ \ \text{if}\ \ t\geqslant\varepsilon,\ \beta_{\varepsilon}(0)=-\mathcal{K}_{0}\ \ \text{where}\ \ \mathcal{K}_{0}:=2\{(q+r)e^{n}+2rK_{c}+5r\},\\ \ \displaystyle{\lim_{\varepsilon\to 0}\beta_{\varepsilon}(t)}=0\ \ \text{if}\ \ t>0,\quad\displaystyle{\lim_{\varepsilon\to 0}\beta_{\varepsilon}(t)}=-\infty\ \ \text{if}\ \ t<0\end{cases}$
and $\varphi_{\varepsilon}(t)\in C^{{\infty}}(\mathbb{R})$ satisfying (2.1). We then consider the following penalized problem;
(3.3) $\begin{cases}\partial_{\tau}V_{n,\varepsilon}-\mathcal{L}V_{n,\varepsilon}+\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})=0\quad\text{in }\ \Omega^{n}_{T},\\ \partial_{x}V_{n,\varepsilon}(\tau,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}V_{n,\varepsilon}(\tau,n)=e^{n}\quad\text{for }\ \tau\in[0,T],\\ V_{n,\varepsilon}(0,x)=J_{\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n),\end{cases}$
where
(3.4) $J_{\varepsilon}:=\varphi_{\varepsilon}(C-P)+P.$
Theorem 3.1.
For each fixed $n\in\mathbb{N}\setminus\{0\}$ with $n>\textup{max}\{|{x_{c}}(0)|,|{x_{p}}(0)|\}$, there exists a unique solution $V_{n,\varepsilon}\in W^{1,2}_{p}(\Omega_{T}^{n})\cap C(\overline{\Omega_{T}^{n}})$ to the problem (3.3), where $1<p<\infty.$
Proof.
We apply the Schauder fixed point theorem [10, 280p] to the following setting. Set $\mathcal{B}:=C(\overline{\Omega_{T}^{n}})$ and $\mathcal{D}:=\{w\in\mathcal{B}:w\geqslant 0\}$. Then $\mathcal{D}$ is a closed convex subset of the Banach space $\mathcal{B}.$ For each $w\in\mathcal{D}$, let $u\in W^{1,2}_{p}(\Omega_{T}^{n})$ be the unique solution to
(3.5) $\begin{cases}\ \partial_{\tau}u-\mathcal{L}u+\beta_{\varepsilon}(w-J_{\varepsilon})=0\quad\text{in }\ \Omega^{n}_{T},\\ \ \partial_{x}u(\tau,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}u(\tau,n)=e^{n}\quad\text{for }\ \tau\in[0,T],\\ \ u(0,x)=J_{\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
Note that the well-posedness of the solution to (3.5) in $W^{1,2}_{p}$ can be found in [19], and the relevant $W^{1,2}_{p}(\Omega_{T}^{n})$ estimate implies
$\|u\|_{W^{1,2}_{p}(\Omega_{T}^{n})}\leqslant\mathcal{K}(\|e^{n}+e^{-n}\|_{W^{1}_{p}([0,T))}+\|J_{\varepsilon}\|_{W^{2}_{p}((-n,n))}+\|\beta_{\varepsilon}(w-J_{\varepsilon})\|_{L^{p}(\Omega_{T}^{n})})$
for some constant $\mathcal{K}>0$. Moreover, the parabolic Sobolev embedding theorem [19] yields that $u$ is Hölder continuous in $\overline{\Omega_{T}^{n}}$. Thus $u\in\mathcal{B}.$ Now we define operator $\mathcal{F}:\mathcal{D}\rightarrow\mathcal{B}$ by $\mathcal{F}(w):=u$. In order to use the Schauder fixed point theorem in $\mathcal{B},$ it suffices to show the following three properties:
  • (1)
    $\mathcal{F}(\mathcal{D})\subset\mathcal{D}$;
  • (2)
    $\mathcal{F}$ is continuous;
  • (3)
    $\mathcal{F}(\mathcal{D})$ is precompact in $\mathcal{B}$.
We shall prove it in Appendix. Utilizing the Schauder fixed point theorem, we obtain the solution $V_{n,\varepsilon}$ of the problem (3.3). In particular, $V_{n,\varepsilon}\in W^{1,2}_{p}(\Omega^{n}_{T})\cap C(\overline{\Omega^{n}_{T}})$ for each $1<p<\infty.$
Once we have obtained the solution $V_{n,\varepsilon}$ of the problem (3.3), we focus on the convergence of $V_{n,\varepsilon}$ as $\varepsilon\rightarrow 0^{+}$. Our claim is that the limit exists and is in fact, a solution to the obstacle problem (3.1).
Lemma 3.2.
Let $V_{n,\varepsilon}$ be the solution of (3.1). Then for each $n\in\mathbb{N},$ there exists some constant $\mathcal{K}>0$ independent of $\varepsilon\in(0,1)$ such that
$-\mathcal{K}\leqslant\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\leqslant 0\quad\text{and}\quad 0\leqslant V_{n,\varepsilon}\leqslant\mathcal{K}\quad\text{in }\ \Omega_{T}^{n}.$
Proof.
From Lemma 2.1 (ii), we note
$0\leqslant\partial_{x}C\leqslant e^{x}\ \ \text{and}\ \ -e^{x}\leqslant\partial_{x}P\leqslant 0\quad\text{in}\ \ \Omega^{n}_{T}.$
Next, observe that $\partial_{x}J_{\varepsilon}=(1-\varphi^{\prime}_{\varepsilon}(C-P))\partial_{x}P+\varphi^{\prime}_{\varepsilon}(C-P)\partial_{x}C$ and $0\leqslant\varphi^{\prime}_{\varepsilon}\leqslant 1$. Combining these estimates, we deduce
(3.6) $-e^{x}\leqslant\partial_{x}J_{\varepsilon}\leqslant e^{x}\quad\text{in}\ \ \Omega^{n}_{T}.$
Moreover, by Lemma 2.1 (i), we have
$|C-P|\leqslant C+P\leqslant e^{n}+K_{p}+4\quad\text{in}\ \ \Omega_{T}^{n}$
and since $C,P\in W^{1,2}_{p,\text{loc}}(\Omega_{T})$, it follows that almost everywhere in $\Omega_{T}^{n}$,
[…]
By direct computation together with the choice of $\mathcal{K}_{0}$, $J_{\varepsilon}$ satisfies that in $\Omega^{n}_{T}$,
$\partial_{\tau}J_{\varepsilon}-\mathcal{L}J_{\varepsilon}$ $=\varphi^{\prime}_{\varepsilon}(C-P)\big{[}\partial_{\tau}C-\mathcal{L}C\big{]}+\big{\{}1-\varphi_{\varepsilon}^{\prime}(C-P)\big{\}}\big{[}\partial_{\tau}P-\mathcal{L}P\big{]}$
$\ \ -\frac{\sigma^{2}}{2}\varphi^{\prime\prime}_{\varepsilon}(C-P)\big{[}\partial_{x}(C-P)\big{]}^{2}-r\varphi^{\prime}(C-P)\big{[}C-P\big{]}+r\varphi_{\varepsilon}(C-P)$
$\leqslant\ \varphi^{\prime}(\cdot)\big{|}qe^{n}-rK_{c}\big{|}+\big{\{}1-\varphi^{\prime}(\cdot)\big{\}}\big{|}-qe^{n}+rK_{p}\big{|}+2r\big{[}\max\limits_{\Omega^{n}_{T}}\big{|}C-P\big{|}+1\big{]}$
$\leqslant 2(qe^{n}+rK_{c})+2r(e^{n}+K_{p}+5)$
$\leqslant 2(q+r)e^{n}+4rK_{c}+10r$
$=\mathcal{K}_{0}=-\beta_{\varepsilon}(0)$
for sufficiently small $10<\varepsilon<$ 1. Consequently, we deduce
$\begin{cases}\partial_{\tau}J_{\varepsilon}-\mathcal{L}J_{\varepsilon}+\beta_{\varepsilon}(0)\leqslant 0\quad\text{in }\ \Omega^{n}_{T},\\ \partial_{x}J_{\varepsilon}(\tau,-n)\geqslant-e^{-n}=\partial_{x}V_{n,\varepsilon}(\tau,-n)\quad\text{and}\quad\partial_{x}J_{\varepsilon}(\tau,n)\leqslant e^{n}=\partial_{x}V_{n,\varepsilon}(\tau,n)\quad\text{for }\ \tau\in[0,T],\\ J_{\varepsilon}(0,x)=V_{n,\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
Hence, $J_{\varepsilon}$ is a subsolution to (3.3) and by the comparison principle [6, 53p]
$V_{n,\varepsilon}\geqslant J_{\varepsilon}\geqslant 0\quad\text{in }\ \Omega_{T}^{n}.$
Moreover, the monotonicity of $\beta_{\varepsilon}$ implies
$-\mathcal{K}_{0}=\beta_{\varepsilon}(0)\leqslant\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\leqslant 0.$
Next, our claim is that $V_{n,\varepsilon}$ is bounded. To this end, define $h(\tau,x):=e^{x}+e^{-n}x^{2}+\mathcal{K}_{1}$ where $\mathcal{K}_{1}$ is a constant satisfying $\mathcal{K}_{1}\geqslant\max\{(1-\frac{q}{r}-\frac{\sigma^{2}}{2r})^{2}+\frac{\sigma^{2}}{r},K_{p}+6\}.$ This implies $h\geqslant J_{\varepsilon}+\varepsilon$ and hence $\beta_{\varepsilon}(h-J_{\varepsilon})=0$ in $\Omega_{T}^{n}.$ Moreover in $\Omega_{T}^{n}$,
$\partial_{\tau}h-\mathcal{L}h=qe^{x}+e^{-n}\{rx^{2}-(2r-2q-\sigma^{2})x-\sigma^{2}\}+r\mathcal{K}_{1}\geqslant 0.$
Thus, $h$ satisfies
$\begin{cases}\partial_{\tau}h-\mathcal{L}h+\beta_{\varepsilon}(h-J_{\varepsilon})\geqslant 0\quad\text{in }\ \Omega^{n}_{T},\\ \partial_{x}h(\tau,-n)=(1-2n)e^{-n}\leqslant\partial_{x}V_{n,\varepsilon}(\tau,-n)\quad\text{and}\\ \partial_{x}h(\tau,n)=2ne^{-n}+e^{n}\geqslant\partial_{x}V_{n,\varepsilon}(\tau,n)\quad\text{for }\ \tau\in[0,T],\\ h(0,x)\geqslant V_{n,\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
By the comparison principle [6, 53p],
$V_{n,\varepsilon}\leqslant h\leqslant e^{n}+e^{-n}n^{2}+\mathcal{K}_{1}\quad\text{in}\ \ \Omega^{n}_{T}.\qed$
[…]
Theorem 3.3.
For each fixed $n\in\mathbb{N}\setminus\{0\}$ with $n>\textup{max}\{|{x_{c}}(0)|,|{x_{p}}(0)|\}$, there exists a unique solution $V_{n}\in C(\overline{\Omega^{n}_{T}})\cap W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))$ to the problem (3.1), where $1<p<\infty$ and $\varrho>0$.
Proof.
For the solution $V_{n,\varepsilon}$ to we consider the $W^{1,2}_{p}$ estimate in the domain $\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}),$ where $1<p<\infty$ and $\varrho>0$ and $C^{\frac{\alpha}{2},\alpha}$ estimate for $V_{n,\varepsilon}$ [19, 14]:
$\|V_{n,\varepsilon}\|_{W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))}$ $\leqslant\mathcal{K}\Big{(}\|V_{n,\varepsilon}\|_{L^{\infty}(\Omega^{n}_{T})}++\|\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\|_{L^{p}(\Omega^{n}_{T})}+\|e^{-n}+e^{n}\|_{W^{1}_{p}([0,T))}$
$\quad\quad+\|\varphi_{\epsilon}(C-P)+P\|_{W^{2}_{p}((-n,-n)\setminus(\bar{x}-\delta/2,\bar{x}+\delta/2))}\Big{)}$
$\leqslant\mathcal{K}^{\prime},$
and
$\|V_{n,\varepsilon}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega^{n}_{T}})}\leqslant\mathcal{K}^{\prime\prime}.$
Note that the constants $\mathcal{K}^{\prime}$ and $\mathcal{K}^{\prime\prime}$ are independent of $\varepsilon\in(0,\delta/2)$ due to Lemma 3.2 and the property of $\varphi_{\varepsilon}$ in (2.1). Therefore, for each large $n\in\mathbb{N}$, $V_{n,\varepsilon}$ is bounded in $C(\overline{\Omega^{n}_{T}})\cap W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,x))$ for $\varepsilon\in(0,\delta/2)$, and hence there exists $V_{n}\in C(\overline{\Omega^{n}_{T}})\cap W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))$ for all $1<p<\infty$ and $\varrho>0$ such that
$V_{n,\varepsilon}\rightarrow V_{n}\quad$ $\text{in}\quad C(\overline{\Omega^{n}_{T}}),$
(3.7) $V_{n,\varepsilon}\rightharpoonup V_{n}\quad$ $\text{in}\quad W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})),$
up to a subsequence, as $\varepsilon\rightarrow 0^{+}$. Moreover, the Sobolev embedding theorem yields
$\partial_{x}V_{n,\varepsilon}\rightarrow\partial_{x}V_{n}\quad$ $\text{in}\quad C(\overline{\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})}).$
Now, let us verify that $V_{n}$ is a solution to (3.1). Fix any $\psi\in C_{c}^{\infty}(\{V_{n}>J\}\cap\Omega^{n}_{T})$. Then there exist small $\varrho,\delta>0$ such that
$\text{supp}(\psi)\subset\{V_{n}>J\}\cap\Omega^{n}_{T}\setminus B_{\varrho}(0,x_{0})\quad\text{and}\quad\min\limits_{\text{supp}(\psi)}(V_{n}-J)>2\delta.$
Choose $\varepsilon_{0}>0$ small so that for every $\varepsilon\in(0,\varepsilon_{0}]$,
$|(V_{n,\varepsilon}-J_{\varepsilon})-(V_{n}-J)|<\delta\quad\text{and}\quad\varepsilon<\delta/2.$
Then for such small $\varepsilon>0$
$V_{n,\varepsilon}-J_{\varepsilon}>\delta>\varepsilon,$
and this implies
$\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})=0\ \ \text{in}\ \text{supp}(\psi).$
Therefore, we obtain
[…]
Thus, by (3.7), taking $\varepsilon\rightarrow 0^{+}$ yields $\partial_{\tau}V_{n}-\mathcal{L}V_{n}=0$ a.e. in $\{V_{n}>J\}\cap\Omega^{n}_{T}.$ We next show that $V_{n}$ is a supersolution in $\Omega^{n}_{T}$. Fix any non-negative $\psi\in C_{c}^{\infty}(\Omega^{n}_{T})$. Then Lemma 3.2 implies that for all $\varepsilon\in(0,1),$
[…]
We refer to (3.7) and hence taking $\varepsilon\rightarrow 0^{+}$ yields
[…]
Therefore, $V_{n}$ is a solution to (3.1).
Now, we prove the uniqueness of $V_{n}.$ Let $V_{n}$ and $\tilde{V}_{n}$ be two solutions of (3.1). Then, define $\mathcal{N}:=\{(\tau,x)\in\overline{\Omega^{n}_{T}}:{\tilde{V}_{n}}(\tau,x)>V_{n}(\tau,x)\}$ and suppose that $\mathcal{N}$ is nonempty. Notice that $V_{n},\tilde{V}_{n}$ are continuous functions. Hence the set $\mathcal{N}$ is open, contains a ball inside. In the set $\mathcal{N},$ we have $\tilde{V}_{n}>J$ so
$\partial_{\tau}\tilde{V}_{n}-\mathcal{L}\tilde{V}_{n}=0\ \ \text{in}\ \ \mathcal{N}.$
Next, we define an operator $\mathcal{M}$ by
$\mathcal{M}u:=-u+\beta\cdot Du$
where $\beta$ is the inward pointing unit vector field with its $\tau$-direction $\beta^{\tau}=0$ on $\mathcal{P}\mathcal{N}$. Since both $V_{n},\tilde{V}_{n}$ and each of their partial derivative with respect to $x$ are continuous by the Sobolev embedding, $\mathcal{M}(\tilde{V}_{n}-V_{n})=0$ on the parabolic boundary $\mathcal{P}\mathcal{N}$. Thus,
$\begin{cases}\partial_{\tau}{(\tilde{V}_{n}-V_{n})}-\mathcal{L}{(\tilde{V}_{n}-V_{n})}\leqslant 0\quad&\text{in }\ \mathcal{N},\\ \mathcal{M}{(\tilde{V}_{n}-V_{n})}=0\quad&\text{on}\ \ \mathcal{P}\mathcal{N}.\end{cases}$
By the comparison principle [14, Corollary 7.4], it follows that $\tilde{V}_{n}-V_{n}\leqslant 0$ and this contradicts the definition of $\mathcal{N}$ and hence $\mathcal{N}=\emptyset$. Applying the same procedure as above to $V_{n}-\tilde{V}_{n}$ in the set $\{V_{n}>\tilde{V}_{n}\}$ gives $\{V_{n}>\tilde{V}_{n}\}=\emptyset$. This proves $\tilde{V}_{n}\equiv V_{n}$ in $\Omega_{T}^{n}.$
3.2. Solution to the obstacle problem
Finally, we prove the existence and uniqueness of the solution to the obstacle problem (2.7).
Theorem 3.4.
There exists a unique solution $V\in W^{1,2}_{p,\mathrm{loc}}(\Omega_{T})\cap C(\overline{\Omega_{T}})$ to (2.7), where $1<p<\infty$.
Proof.
Let $V_{n}\in W^{1,2}_{p,\mathrm{loc}}(\Omega_{T}^{n})\cap C(\overline{\Omega^{n}_{T}})$ be the unique solution to the obstacle problem (3.1). Then it satisfies that
$\begin{cases}\partial_{\tau}V_{n}-\mathcal{L}V_{n}=f_{n}\quad\text{in }\ \Omega_{T}^{n},\\ \partial_{x}V_{n}(\tau,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}V_{n}(\tau,-n)=e^{n}\quad\text{for }\ \tau\in[0,T],\\ V_{n}(0,x)=\max\{C(0,x),P(0,x)\}\quad\text{for }\ x\in(-n,n),\end{cases}$
[…]
$f_{n}(\tau,x)\leqslant M\quad\text{in}\ \ \Omega^{R}_{T}:=(0,T)\times(-R,R)$
for some constant $M$ depending on $R$ but independent of $n$. Then it follows from $W^{1,2}_{p}$ estimates in [19, 355p] that for each small $\varrho>0,$
$\|V_{n}\|_{W^{1,2}_{p}(\Omega^{R}_{T}\setminus B_{\varrho}(0,\bar{x}))}$ $\leqslant\mathcal{K}(\|V_{n}\|_{L^{{}^{\infty}}(\Omega^{2R}_{T})}+\|C(0,\cdot)\|_{W^{2}_{p}([-2R,2R])}$
$\quad+\|P(0,\cdot)\|_{W^{2}_{p}([-2R,2R])}+\|f_{n}\|_{L^{{}^{\infty}}(\Omega^{2R}_{T})})$
$\leqslant\mathcal{K}^{\prime}$
for some constant $\mathcal{K}^{\prime}$ independent of $n$. Letting $n\rightarrow\infty,$ we deduce that up to a subsequence
$V_{n}\rightharpoonup V^{R}$ $\text{in}\ \ W^{1,2}_{p}(\Omega^{R}_{T}\setminus B_{\varrho}(0,\bar{x}))\quad\text{for\ \ all }\ 1<p<\infty\ \ \text{and}\ \ \varrho>0.$
Moreover, by a standard diagonal argument, we can find a subsequence $\{V_{n_{k}}\}_{k=1}^{\infty}$ such that, as $k\rightarrow\infty$,
$V_{n_{k}}\rightarrow V^{m}$ $\text{in}\ \ C(\overline{\Omega^{m}_{T}}),$
$V_{n_{k}}\rightharpoonup V^{m}$ $\text{in}\ \ W^{1,2}_{p}(\Omega^{m}_{T}\setminus B_{\varrho}(0,\bar{x}))\quad\text{for\ \ all }\ 1<p<\infty\ \ \text{and}\ \ \varrho>0$
for all large $m\in\mathbb{N}.$ Thus, $V^{m+1}=V^{m}$ in $\Omega^{m}_{T}$ so we can define $V:=V^{m}$ in $\overline{\Omega_{T}^{m}}$ for all large $m\in\mathbb{N}$. In addition, the Sobolev embedding theorem [19, Chapter II. Lemma 3.3] yields
$\partial_{x}V_{n_{k}}\rightarrow\partial_{x}V^{m}$ $\text{in}\ \ C(\overline{\Omega^{m}_{T}\setminus B_{\varrho}(0,\bar{x})}).$
Now let us prove that $V^{m}$ satisfies the following obstacle problem:
$\begin{cases}\partial_{\tau}V^{m}-\mathcal{L}V^{m}\geqslant 0\quad\ \text{for }\ (\tau,x)\in\Omega^{m}_{T}\ \ \text{with}\ \ V^{m}(\tau,x)=J(\tau,x),\\ \partial_{\tau}V^{m}-\mathcal{L}V^{m}=0\quad\ \text{for }\ (\tau,x)\in\Omega^{m}_{T}\ \ \text{with}\ \ V^{m}(\tau,x)>J(\tau,x),\\ V(0,x)=J(0,x)\quad\text{for }\ x\in(-m,m).\end{cases}$
Since $\partial_{\tau}V_{n}-\mathcal{L}V_{n}\geqslant 0$ in $\Omega_{T}^{n}$, by the weak convergence we see that $\partial_{\tau}V^{m}-\mathcal{L}V^{m}\geqslant 0$ and $V^{m}\geqslant J$ in $\Omega^{m}_{T}$. Thus, it remains to show that $\partial_{\tau}V^{m}-\mathcal{L}V^{m}=0$ in the set $\{V^{m}>J\}$. Fix any $\psi\in C_{c}^{\infty}(\{V^{m}>J\}\cap\Omega^{m}_{T})$ and $\varrho>0$ such that $\text{supp}(\psi)\subset\{V^{m}>J\}\cap\Omega_{T}^{m}\setminus B_{\varrho}(0,\bar{x})$. Then, there exists $\delta>0$ such that
$\min\limits_{\text{supp}(\psi)}(V^{m}-J)>\delta.$
Let us fix $n\in\mathbb{N}$ and define $S_{1,{n}}$ and $S_{2,{n}}$ by
$S_{1,{n}}:=\{(\tau,x)\in\Omega^{m}_{T}:|(V_{n}-J)-(V^{m}-J)|<\delta/2\},$
$S_{2,{n}}:=\{(\tau,x)\in\Omega^{m}_{T}:|(V_{n}-J)-(V^{m}-J)|\geqslant\delta/2\}.$
Notice that $V_{n}-J>\frac{\delta}{2}$ in $S_{1,{n}}$, which implies $S_{1,{n}}\subset\{(\tau,x)\in\Omega_{T}^{n}:V_{n}(\tau,x)>J(\tau,x)\}\cap\Omega^{m}_{T}.$ Next, recall that $\partial_{\tau}V_{n}-\mathcal{L}V_{n}=0$ holds almost everywhere in $\{(\tau,x)\in\Omega_{T}^{n}:V_{n}(\tau,x)>J(\tau,x)\}$. This leads to the conclusion that
$\iint_{S_{1,n}}({\partial_{\tau}V_{n}-\mathcal{L}V_{n}})\psi=0.$
Note also that the uniform convergence of $V_{n}-J$ to $V^{m}-J$ implies that $V_{n}-J$ converges to $V^{m}-J$ in measure as $n\rightarrow\infty$. In other words, the measure of the set $S_{2,n}$ approaches 00 as $n\rightarrow\infty$. Together with $\|V_{n_{k}}\|_{W^{1,2}_{p}(\Omega^{m}_{T}\setminus B_{\varrho}(0,\bar{x}))}\leqslant M$, we deduce that
$\displaystyle{\lim_{k\to\infty}}\iint_{S_{2,n_{k}}}({\partial_{\tau}V_{n_{k}}-\mathcal{L}V_{n_{k}}})\psi=0.$
Combining the above two identities, we obtain that
$\iint_{\Omega^{m}_{T}}({\partial_{\tau}V^{m}-\mathcal{L}V^{m}})\psi$ $=\displaystyle{\lim_{k\to\infty}}\iint_{S_{1,n_{k}}\cup S_{2,n_{k}}}({\partial_{\tau}V_{n_{k}}-\mathcal{L}V_{n_{k}}})\psi=0,$
and hence $\partial_{\tau}V^{m}-\mathcal{L}V^{m}=0$ in $\{V^{m}>J\}$.
Therefore, since $V=V^{m}$ for all $m\in\mathbb{N}$, $V$ is a solution to (2.7). For the uniqueness of a solution $V$ to (2.7), we notice that by [12, Remark 3], since $|e^{x}-K|<\max\{K,1\}e^{|x|}$ for $K>0$ and $x\in\mathbb{R}$, $|\max\{C,P\}|\leqslant M_{1}e^{|x|}$ for some $M_{1}>0$ and hence $|V|\leqslant M_{2}e^{|x|}$ for some $M_{2}>0$. Therefore, by [12, Theorem 3] we obtain the uniqueness of $V$. ∎
3.3. Estimate for the derivatives of V
Lemma 3.5.
Let $V$ be a solution to (2.7). Then the following properties hold:
  • (1)
    $\partial_{\tau}V\geqslant 0$,
  • (2)
    $-e^{x}\leqslant\partial_{x}V\leqslant e^{x}$.
Proof.
(1) Fix any $\delta\in(0,T)$ and consider $V_{n,\varepsilon}(\tau+\delta,x)$ for all $(\tau,x)\in\Omega^{n}_{T-\delta}.$ Observe that $V_{n,\varepsilon}(\tau+\delta,x)$ satisfies
(3.8) $\begin{cases}\partial_{\tau}V_{n,\varepsilon}(\tau+\delta,x)-\mathcal{L}V_{n,\varepsilon}(\tau+\delta,x)\\ \quad\quad+\beta_{\varepsilon}(V_{n,\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau+\delta,x))=0\quad\text{for }\ (\tau,x)\in\Omega^{n}_{T-\delta},\\ \partial_{x}V_{n,\varepsilon}(\tau+\delta,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}V_{n,\varepsilon}(\tau+\delta,n)=e^{n}\quad\text{for }\ \tau\in[0,T-\delta],\end{cases}$
Subtracting (3.3) from (3.8) and the mean value theorem yield
$\begin{cases}\partial_{\tau}\{V_{n,\varepsilon}(\tau+\delta,x)-V_{n,\varepsilon}(\tau,x)\}-\mathcal{L}\{V_{n,\varepsilon}(\tau+\delta,x)-V_{n,\varepsilon}(\tau,x)\}\\ \ \ +\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})\{V_{n,\varepsilon}(\tau+\delta,x)-V_{n,\varepsilon}(\tau,x)\}=\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})\{J_{\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau,x)\}\ \ \text{for }(\tau,x)\in\Omega^{n}_{T-\delta},\\ \partial_{x}\{V_{n,\varepsilon}(\tau+\delta,-n)-V_{n,\varepsilon}(\tau,-n)\}=0\quad\text{and}\\ \qquad\qquad\qquad\partial_{x}\{V_{n,\varepsilon}(\tau+\delta,n)-V_{n,\varepsilon}(\tau,x)\}=0\quad\text{for }\ \tau\in[0,T-\delta],\\ V_{n,\varepsilon}(\delta,x)-V_{n,\varepsilon}(0,x)=V_{n,\varepsilon}(\delta,x)-J_{\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n),\end{cases}$
where $\gamma_{\tau,x}>0$ is a positive number in between $V_{n,\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau+\delta,x)\geqslant 0$ and $V_{n,\varepsilon}(\tau,x)-J_{\varepsilon}(\tau,x)\geqslant 0$, determined by the mean value theorem. In the case that the latter two values are equal, we choose $\gamma_{\tau,x}$ to be large so that $\beta^{\prime}_{\epsilon}(\gamma_{\tau,x})=0$ (see (3.2)). Then from definition (3.2) it follows that $\beta^{\prime}_{\epsilon}(\gamma_{\tau,x})$ is nonnegative and bounded for $(\tau,x)\in\Omega^{n}_{T-\delta}$. In addition, by (3.4), (2.1) and Lemma 2.1 (iii), we have $(J_{\epsilon})_{\tau}=\varphi^{\prime}_{\epsilon}(C-P)(C_{\tau}-P_{\tau})+P_{\tau}\geqslant 0$. Hence the right hand side of the first equation $\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})\{J_{\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau,x)\}\geqslant 0$ for all $(\tau,x)\in\Omega^{n}_{T-\delta}$. Moreover, since $V_{n,\varepsilon}(\delta,x)\geqslant J_{\varepsilon}(\delta,x)$ and $V_{n,\varepsilon}(0,x)=J_{\varepsilon}(0,x)$ for all $x\in(-n,n)$, we deduce
$V_{n,\epsilon}(\delta,x)-V_{n,\epsilon}(0,x)\geqslant J_{\epsilon}(\delta,x)-J_{\epsilon}(0,x)\geqslant 0\quad\text{for }\ x\in(-n,n).$
Therefore, by the comparison principle,
$V_{n,\varepsilon}(\tau+\delta,x)\geqslant V_{n,\varepsilon}(\tau,x)\quad\text{for all}\ \ (\tau,x)\in\Omega^{n}_{T-\delta}.$
Take $\varepsilon\rightarrow 0$ and $n\rightarrow\infty$ to get the desired inequality.
(2) From the $W^{1,2}_{p}$ estimate and the Sobolev embedding theorem for the solution $V_{n,\varepsilon}$ of (3.3), we obtain that $V_{n,\varepsilon}\in C^{\frac{\alpha}{2},\alpha}(\overline{\Omega^{n}_{T}})$ for some $\alpha\in(0,1)$. Thus, the right hand side $\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\in C^{\frac{\alpha}{2},\alpha}(\Omega^{n}_{T})$. By the Schauder theory, we deduce that $V_{n,\varepsilon}\in C^{1+\frac{\alpha}{2},2+\alpha}(\Omega_{T}^{n})$. This increases the regularity of the right hand side and thus, applying the Schauder theory again yields $V_{n,\varepsilon}\in C^{2+\frac{\alpha}{2},4+\alpha}(\Omega_{T}^{n})$. By following such bootstrap argument, we conclude that $V_{n,\varepsilon}$ is smooth in $\Omega_{T}^{n}$. Differentiating the PDE (3.3) with respect to $x$, we observe that $\partial_{x}V_{n,\varepsilon}$ satisfies
$\begin{cases}\partial_{\tau}(\partial_{x}V_{n,\varepsilon})-\mathcal{L}(\partial_{x}V_{n,\varepsilon})+\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})(\partial_{x}V_{n,\varepsilon})=\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\partial_{x}J_{\varepsilon}\quad\text{in }\ \Omega_{T}^{n},\\ \partial_{x}V_{n,\varepsilon}(\tau,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}V_{n,\varepsilon}(\tau,n)=e^{n}\quad\text{for }\ \tau\in[0,T],\\ \partial_{x}V_{n,\varepsilon}(0,x)=\partial_{x}J_{\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
Define a linear operator $\tilde{\mathcal{L}}$ by
~$\tilde{\mathcal{L}}:=\mathcal{L}-\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})=\frac{\sigma^{2}}{2}\partial_{xx}+(r-q-\frac{\sigma^{2}}{2})\partial_{x}-(r+\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon}))$
Note that the zeroth-order coefficient of $\tilde{\mathcal{L}}$ being nonnegative in $\Omega^{n}_{T}$. Then, $\partial_{x}V_{n,\varepsilon}$ satisfies the equation $(\partial_{\tau}-\tilde{\mathcal{L}})(\partial_{x}V_{n,\varepsilon})=\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\partial_{x}J_{\varepsilon}$ in $\Omega^{n}_{T}$. Since $\mathcal{L}e^{x}=-qe^{x}$, with (3.6), we see that $\partial_{x}V_{n,\varepsilon}-e^{x}$ satisfies
$\begin{cases}\partial_{\tau}(\partial_{x}V_{n,\varepsilon}-e^{x})-\tilde{\mathcal{L}}(\partial_{x}V_{n,\varepsilon}-e^{x})=\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})(\partial_{x}J_{\varepsilon}-e^{x})-qe^{x}\leqslant 0\quad\text{in }\ \Omega_{T}^{n},\\ \partial_{x}V_{n,\varepsilon}(\tau,-n)-e^{-n}=-2e^{-n}\quad\text{and}\quad\partial_{x}V_{n,\varepsilon}(\tau,n)-e^{n}=0\quad\text{for }\ \tau\in[0,T],\\ \partial_{x}V_{n,\varepsilon}(0,x)-e^{x}=\partial_{x}J_{\varepsilon}(0,x)-e^{x}\leqslant 0\quad\text{for }\ x\in(-n,n).\end{cases}$
Thus, by the maximum principle, we deduce $\partial_{x}V_{n,\varepsilon}\leqslant e^{x}$ in $\Omega_{T}^{n}$. Similiarly, $\partial_{x}V_{n,\varepsilon}+e^{x}$ satisfies
$\begin{cases}\partial_{\tau}(\partial_{x}V_{n,\varepsilon}+e^{x})-\tilde{\mathcal{L}}(\partial_{x}V_{n,\varepsilon}+e^{x})=\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})(\partial_{x}J_{\varepsilon}+e^{x})+qe^{x}\geqslant 0\quad\text{in }\ \Omega_{T}^{n},\\ \partial_{x}V_{n,\varepsilon}(\tau,-n)+e^{-n}=0\quad\text{and}\quad\partial_{x}V_{n,\varepsilon}(\tau,n)+e^{n}=2e^{n}\quad\text{for }\ \tau\in[0,T],\\ \partial_{x}V_{n,\varepsilon}(0,x)+e^{x}=\partial_{x}J_{\varepsilon}(0,x)+e^{x}\geqslant 0\quad\text{for }\ x\in(-n,n)\end{cases}$
and by the maximum principle again, $\partial_{x}V_{n,\varepsilon}\geqslant-e^{x}$ in $\Omega^{n}_{T}$. Take $\varepsilon\rightarrow 0^{+}$ and $n\rightarrow\infty$ to get the desired inequality.
4. Analysis of the free boundary
4.1. Representation of the free boundary
Let us define the exercise region $\mathcal{E}$ and the continuation region $\mathcal{C}$ for the solution $V$ of (2.7) as
$\mathcal{E}:=$$\{(\tau,x)\in\Omega_{T}:V(\tau,x)=J(\tau,x)\}\quad\text{and}\quad\mathcal{C}:=\{(\tau,x)\in\Omega_{T}:V(\tau,x)>J(\tau,x)\}.$
Furthermore, let us define $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ as
$\mathcal{E}_{p}^{\rm ch}:=$$\{(\tau,x)\in\Omega_{T}:V(\tau,x)=P(\tau,x)\}\quad\text{and}\quad\mathcal{E}_{c}^{\rm ch}:=$$\{(\tau,x)\in\Omega_{T}:V(\tau,x)=C(\tau,x)\}.$
Note that $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are closed, and since $V\geqslant J$, we have
$\mathcal{E}_{p}^{\rm ch}\cup\mathcal{E}_{c}^{\rm ch}=\mathcal{E}.$
In the next lemma, we will show that $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are disjoint, hence $\partial\mathcal{E}$ is the disjoint union of $\partial\mathcal{E}_{p}^{\rm ch}$ and $\partial\mathcal{E}_{c}^{\rm ch}$. Therefore, examining $\partial\mathcal{E}_{p}^{\rm ch}$ and $\partial\mathcal{E}_{c}^{\rm ch}$ separately is sufficient to analyze $\partial\mathcal{E}$.
We start with proving that $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are nonempty.
Lemma 4.1.
Let $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ be defined as above. Then, $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are nonempty.
Proof.
We first observe from the increasing property of $x_{c}(\tau)$ and the inequality $x_{p}(\tau)<x_{c}(\tau)$ that $C(\tau,x)>e^{x}-K_{c}$ for all $(\tau,x)\in(0,T)\times(-\infty,x_{p}(0))$. Suppose that $\mathcal{E}_{p}^{\rm ch}$ is empty. Choose $x_{0}\in\mathbb{R}$ and $\delta>0$ such that $x_{0}+3\delta<\min\{x_{p}(0),\bar{x}\}$ and set $Q:=(0,\delta)\times(x_{0}-2\delta,x_{0}-\delta)$. Then $V$ satisfies $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$ since
$\partial_{\tau}V-\mathcal{L}V=\begin{cases}\partial_{\tau}C-\mathcal{L}C\quad&\text{in }\ \mathcal{E}\cap Q,\\ 0\quad&\text{in }\ \mathcal{C}\cap Q,\end{cases}$
nd $Q\subset\{C>e^{x}-K_{C}\}$. Therefore,
$(\partial_{\tau}-\mathcal{L})\{V-(K_{p}-e^{x})\}=qe^{x}-rK_{p}\leqslant rK_{p}(e^{-\delta}-1)<0\quad\text{in }\ Q.$
Let $\ell_{Q}:=\{0\}\times(x_{0}-2\delta,x_{0}-\delta)$. By the boundary $C^{\frac{\alpha}{2},\alpha}$ estimate [14, Theorem 6.33] for $\partial^{2}_{x}V$ on $Q\cup\mathcal{P}Q$, we have
$\partial_{\tau}V=\partial_{\tau}V-\mathcal{L}\{V-(K_{p}-e^{x})\}\leqslant rK_{p}(e^{-\delta}-1)<0\quad\text{on }\ \ell_{Q}$
since $V=K_{p}-e^{x}$ on $\ell_{Q}$. This contradicts Lemma 3.5 (1). Thus $\mathcal{E}_{p}^{\rm ch}\neq\emptyset.$ Similarly, we can obtain $\mathcal{E}_{c}^{\rm ch}\neq\emptyset.$
Lemma 4.2.
Let $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ be as in above. Then,
(4.1) $\mathcal{E}_{p}^{\rm ch}\subset\{P=K_{p}-e^{x}\}\quad\text{and}\quad\mathcal{E}_{c}^{\rm ch}\subset\{C=e^{x}-K_{c}\}.$
In particular, we have that
(4.2) $\mathcal{E}_{p}^{\rm ch}=\{V=K_{p}-e^{x}\}\quad\text{and}\quad\mathcal{E}_{c}^{\rm ch}=\{V=e^{x}-K_{c}\},$
and $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are disjoint.
Proof.
Suppose that there exists some point $(\tau_{0},x_{0})\in\{P>K_{p}-e^{x}\}\cap\mathcal{E}_{p}^{\rm ch}$. Since $\{P>K_{p}-e^{x}\}$ is open, there exists some $\delta>0$ such that $Q:=(\tau_{0}-\delta,\tau_{0}]\times(x_{0}-\delta,\infty)\subset\{P>K_{p}-e^{x}\}$ (see Figure 1 (B)). Note that
$\partial_{\tau}(V-P)-\mathcal{L}(V-P)\geqslant 0\quad\text{in }\ Q$
since $\partial_{\tau}P-\mathcal{L}P=0$ in $Q$ and $V-P$ attains its minimum value zero at $(\tau_{0},x_{0})\in Q.$ Thus, by the strong maximum principle [11, Corollary 2.4], $V\equiv P$ in $Q$. This is a contradiction since $C(\tau_{0},x)>P(\tau_{0},x)$ if $x$ is sufficiently large, see Lemma 2.1 (iii), and $V\geqslant C$. Therefore, $\{P>K_{p}-e^{x}\}\cap\mathcal{E}_{p}^{\rm ch}=\emptyset$, i.e., $\mathcal{E}_{p}^{\rm ch}\subset\{P=K_{p}-e^{x}\}$. Using a similar argument, we can obtain $\mathcal{E}_{c}^{\rm ch}\subset\{C=e^{x}-K_{c}\}$.
Since $V\geqslant P\geqslant K_{p}-e^{x}$ and $V\geqslant C\geqslant e^{x}-K_{c}$, it is obvious that $\mathcal{E}_{p}^{\rm ch}\supset\{V=K_{p}-e^{x}\}$ and $\mathcal{E}_{c}^{\rm ch}\supset\{C=e^{x}-K_{c}\}$. This together with (4.1) implies (4.2). Moreover, since $\mathcal{E}_{P}=\{P=K_{p}-e^{x}\}$ and $\mathcal{E}_{C}=\{C=e^{x}-K_{c}\}$ are disjoint (see Subsection 2.4), we conclude that $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are also disjoint. ∎
To proceed further, we note that the $x$-coordinate of the points in $\mathcal{E}_{C}$ and $\mathcal{E}_{P}$ is greater than $\ln K_{c}$ and less than $\ln K_{p}$ resepctively (see subsection 2.4). Combined with the above Lemma 4.2, we deduce that the same holds for the $x$-coordinate of each points in $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ respectively. Thus, we take into account Lemma 3.5 and Lemma 4.2 to write the $x$-coordinate of $\partial\mathcal{E}_{p}^{\rm ch}$ and $\partial\mathcal{E}_{c}^{\rm ch}$ by
(4.3) $x_{p}^{\rm ch}(\tau)$ $:=\sup\limits_{x<\ln K_{p}}\ \{x:V(\tau,x)=K_{p}-e^{x}\}\quad\tau\in(0,T),$
(4.4) $x_{c}^{\rm ch}(\tau)$ $:=\inf\limits_{x>\ln K_{c}}\ \{x:V(\tau,x)=e^{x}-K_{c}\}\quad\tau\in(0,T).$
Remark.
Since $\mathcal{E}_{p}^{\rm ch}$ and $\mathcal{E}_{c}^{\rm ch}$ are nonempty, the set $\{x:V(\tau_{1},x)=K_{p}-e^{x}\}$ and $\{x:V(\tau_{2},x)=e^{x}-K_{c}\}$ are nonempty for some $\tau_{1},\ \tau_{2}\in(0,T)$. Moreover, Lemma 3.5 (1) implies that $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$ are monotone increasing and decreasing respectively.
Figure 2. solution $V(\tau,x)$ and obstacle $\max\{C(\tau,x),P(\tau,x)\}.$
Figure 2 illustrates the solution $V(\tau,x)$ of (2.7) and the obstacle $J(\tau,x)=\max\{C(\tau,x),P(\tau,x)\}$. From the figure, we observe that $V(\tau,x)$ and the obstacle meet in the regions where $C(\tau,x)=e^{x}-K_{c}$ and $P(\tau,x)=K_{p}-e^{x}$, which was proven in Lemma 4.2. Therefore, this confirms that
(4.5) $x_{p}^{\rm ch}(\tau)\leqslant x_{p}(\tau)<x_{c}(\tau)\leqslant x_{c}^{\rm ch}(\tau),\quad\quad\forall\;\tau\in[0,T].$
4.2. Properties of the free boundary
At the end of the previous subsection, we represented each free boundary by viewing their $x$-coordinate as the supremum and infimum of the points that touch the obstacle, respectively. To see that such representations correspond to the actual free boundary, we need to establish that the parametrizations in (4.3) and (4.4) are continuous.
Lemma 4.3.
Let $x_{p}^{\rm ch}(\tau),$ $x_{c}^{\rm ch}(\tau):(0,T)\rightarrow\mathbb{R}$ be defined as in (4.3) and (4.4) respectively. Then $x_{p}^{\rm ch}$ and $x_{c}^{\rm ch}$ are continuous on $(0,T)$.
Proof.
Suppose that $x_{p}^{\rm ch}$ is not continuous at some point $\tau\in(\tau_{0},x_{p}^{\rm ch}(\tau_{0})).$ Then by the monotonicity of $x_{p}^{\rm ch}(\tau),$ either ${\lim_{\tau\to{\tau_{0}}^{+}}x_{p}^{\rm ch}(\tau)}>{x_{p}^{\rm ch}}(\tau_{0})$ or ${\lim_{\tau\to{\tau_{0}}^{-}}x_{p}^{\rm ch}(\tau)}<{x_{p}^{\rm ch}}(\tau_{0})$. Let us consider the first case. Then, there exist some $\varepsilon>0$ and $\delta>0$ such that for all $\tau$ in the interval $(\tau_{0},\tau_{0}+\delta)$ we have
$x_{p}^{\rm ch}(\tau_{0})-x_{p}^{\rm ch}(\tau)\geqslant\varepsilon.$
Note that the cylinder $Q:=(\tau_{0},\tau_{0}+\delta)\times(x_{p}^{\rm ch}(\tau_{0})-\frac{\varepsilon}{2},x_{p}^{\rm ch}(\tau_{0}))$ is contained in the set $\{V>K_{p}-e^{x}\}$ and not in the set $\mathcal{E}_{c}^{\rm ch}=\{V=C\}$. Hence, $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$ since
$\partial_{\tau}V-\mathcal{L}V=\begin{cases}\partial_{\tau}P-\mathcal{L}P\quad&\text{in}\ \ Q\cap\mathcal{E},\\ 0\quad&\text{in}\ \ Q\cap\mathcal{C},\end{cases}$
and $\partial_{\tau}P-\mathcal{L}P=0$ in $Q\cap\mathcal{E}\subset\{P>K_{p}-e^{x}\}.$ We recall that $x_{p}^{\rm ch}(\tau_{0})<x_{p}^{\rm ch}(0)<\ln\displaystyle{(\frac{r}{q}K_{p})}$ and this yields
$(\partial_{\tau}-\mathcal{L})\{V-(K_{p}-e^{x})\}=qe^{x}-rK_{p}\leqslant qe^{x_{p}^{\rm ch}(\tau_{0})}-rK_{p}<0\quad\text{in}\ \ Q.$
Note also that $V=K_{p}-e^{x}$ on the line segment $\ell_{Q}:=\{\tau_{0}\}\times(x_{p}^{\rm ch}(\tau_{0})-\frac{\varepsilon}{2},x_{p}^{\rm ch}(\tau_{0}))$. Hence, the boundary $C^{\frac{\alpha}{2},\alpha}$ estimate [14, Theorem 6.33] for $\partial^{2}_{x}V$ yields
$\partial_{\tau}V\leqslant qe^{x_{p}^{\rm ch}(\tau_{0})}-rK_{c}+\mathcal{L}\{V-(K_{p}-e^{x})\}<0\quad\text{on}\ \ \ell_{Q}.$
This contradicts Lemma 3.5 (1).
By following the same procedure for the other discontinuity case ${\lim_{\tau\to{\tau_{0}}^{-}}x_{p}^{\rm ch}(\tau)}<{x_{p}^{\rm ch}}(\tau_{0})$, it can be reached at contradiction. Therefore, $x_{p}^{\rm ch}$ is continuous on $(0,T)$. The continuity of $x_{c}^{\rm ch}$ follows analogously. ∎
In the following theorems, we will explore additional important properties of $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$.
Theorem 4.4.
Let $x_{p}^{\rm ch}(\tau):(0,T)\rightarrow\mathbb{R}$ be defined as in (4.3).
  • i)
    $x_{p}^{\rm ch}$ is strictly decreasing on $(0,T)$.
  • ii)
    ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}=\min\{\bar{x},{x_{p}}(0)\}$, where $\bar{x}$ is the point stated in (2.9).
Proof.
(i) From Lemma 3.5, we infer that $x_{p}^{\rm ch}(\tau)$ is monotone decreasing. Assume, for contradiction, that $x_{p}^{\rm ch}(\tau)$ is not strictly decreasing. Then there exists some $\tau_{1},\tau_{2}>0$ with $\tau_{1}>\tau_{2}$ such that $x_{p}^{\rm ch}(\tau_{1})=x_{p}^{\rm ch}(\tau_{2})=:x_{0}.$ Note that the cylinder $Q:=(\tau_{1},\tau_{2}-\delta)\times(x_{0},\ln{K_{c}})$ is contained in the continuation region $\mathcal{C}$. This implies $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$. Define for each small $\delta>0,$
$V_{\delta}(\tau,x):=V(\tau+\delta,x)-V(\tau,x)\ \ \text{for \ all }\ (\tau,x)\in\Omega_{T-\delta}.$
Then $V_{\delta}$ satisfies
$\partial_{\tau}V_{\delta}-\mathcal{L}V_{\delta}=0\quad$ in $\text{in }\ Q,$
$V_{\delta}(\tau,x_{0})=0\quad$ for  all $\text{for \ all }\ \tau\in(\tau_{1},\tau_{2}-\delta).$
Note that there exists a point $(\tau^{*},x^{*})$ such that $(\tau^{*},x^{*})\in Q$ and $V_{\delta}(\tau^{*},x^{*})=0.$ If not, from the parabolic Hopf lemma we infer that
$\partial_{x}V_{\delta}(\tau,x_{0})>0\quad\text{for \ all}\ \ \tau\in(\tau_{1},\tau_{2}-\delta).$
But $V(\tau,x_{0})=K_{p}-e^{x_{0}}$ implies $\partial_{x}V_{\delta}(\tau,x_{0})=0$ for all $\tau\in(\tau_{1},\tau_{2}-\delta)$, which is a contradiction. From Lemma 3.5, $V_{\delta}(\tau,x)\geqslant 0$ for all $(\tau,x)\in\Omega_{T-\delta}$. Thus, $V_{\delta}$ attains its minimum as zero at $(\tau^{*},x^{*})$. Note also that, by the monotonicity of $x_{c}^{\rm ch}(\tau)$ and $x_{p}^{\rm ch}(\tau)$, we can further obtain $\partial_{\tau}V_{\delta}-\mathcal{L}V_{\delta}=0$ in $\mathcal{C}\cap\Omega_{T-\delta}$. Thus, by the Schauder theory it follows that $V_{\delta}$ is smooth in the region $\mathcal{C}\cap\Omega_{T-\delta}$ and applying the strong maximum principle [14, Theorem 2.7] to $V_{\delta}$ yields
(4.6) $V_{\delta}\equiv 0\quad\text{in }\ \mathcal{C}\cap\Omega_{T-\delta}.$
Set a cylinder $Q^{\prime}:=(0,T-\delta)\times(\ln K_{p},\ln K_{c})$. We see (4.3), (4.4) and the Lemma 3.5 to find that the cylinder $Q^{\prime}$ is contained in $\mathcal{C}\cap\Omega_{T-\delta}$. This implies
$V(\tau,x)>\max\{C(\tau,x),P(\tau,x)\}$
for all $(\tau,x)\in Q^{\prime}$. Moreover, from (4.6) and that $\delta>0$ is arbitrary, we deduce that $V$ is constant with respect to $\tau$ in $Q^{\prime}$. Together with $V=\max\{C(0,\cdot),P(0,\cdot)\}$ at the initial boundary of $Q^{\prime}$ and $V\in C(\overline{\Omega_{T}})$, it follows that
$V(\tau,x)=\max\{C(0,x),P(0,x)\}$
for all $(\tau,x)\in Q^{\prime}$. Combining all, we obtain $\max\{C(0,x),P(0,x)\}>\max\{C(\tau,x),P(\tau,x)\}$ for all $(\tau,x)\in Q^{\prime}$, which contradicts $\partial_{\tau}C$ and $\partial_{\tau}P\geqslant 0$ in Lemma 2.1 iii).
(ii) Define $x_{p}^{\rm ch}(0)$ by
$x_{p}^{\rm ch}(0):=\sup\limits\{x:V(0,x)=K_{p}-e^{x}\}.$
Then, since $V(0,x)=\max\{C(0,x),P(0,x)\}$ for all $x\in\mathbb{R}$, it can be verified that $x_{p}^{\rm ch}(0)=\min\{\bar{x},{x_{p}}(0)\}$. We first note that ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}$ exists since $x_{p}^{\rm ch}(\tau)$ is monotone decreasing and bounded above. Suppose ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}>x_{p}^{\rm ch}(0)$. Then there exists some $\tau_{p}>0$ such that $x_{p}^{\rm ch}(\tau_{p})>x_{p}^{\rm ch}(0)$. This implies
$K_{p}-e^{x_{p}^{\rm ch}(\tau_{p})}$ $=V(\tau_{p},x_{p}^{\rm ch}(\tau_{p}))\geqslant J(\tau_{p},x_{p}^{\rm ch}(\tau_{p}))$
$\geqslant J(0,x_{p}^{\rm ch}(\tau_{p}))>K_{p}-e^{x_{p}^{\rm ch}(\tau_{p})},$
where the second inequality is due to $\partial_{\tau}C\geqslant 0$ and $\partial_{\tau}P\geqslant 0$ (see Lemma 2.1 iii)). This is a contradiction.
On the other hand, if ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}$ is strictly less than $x_{p}^{\rm ch}(0)$, there exist some $\epsilon>0$ and $\delta>0$ such that for all $\tau\in(0,\delta)$,
$x^{\rm ch}_{p}(0)-x^{\rm ch}_{p}(\tau)\geqslant\varepsilon.$
Now, consider a cylinder $Q:=(0,\delta)\times(x^{\rm ch}_{p}(0)-\frac{\varepsilon}{2},x^{\rm ch}_{p}(0))$. Since $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$ and $x^{\rm ch}_{p}(0)<\ln(\displaystyle{\frac{r}{q}}K_{p})$,
$(\partial_{\tau}-\mathcal{L})(V-(K_{p}-e^{x}))\leqslant qe^{x^{\rm ch}_{p}(0)}-rK_{p}<0\quad\text{in}\ \ Q.$
Note that $V=K_{p}-e^{x}$ on the line segment $\ell_{Q}:=\{0\}\times(x^{\rm ch}_{p}(0)-\frac{\varepsilon}{2},x^{\rm ch}_{p}(0))$. By the boundary $C^{\frac{\alpha}{2},\alpha}$ estimate [14, Theorem 6.33] for $\partial^{2}_{x}V$, we obtain
$(\partial_{\tau}-\mathcal{L})(V-(K_{p}-e^{x}))=\partial_{\tau}V<0\quad\text{on}\ \ell_{Q}.$
This contradicts the result in Lemma 3.5. Therefore, we conclude ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}=x_{p}^{\rm ch}(0)$.
By applying the same argument, analogous results for $x_{c}^{\rm ch}$ can be obtained. Hence we omit the proof.
Theorem 4.5.
Let $x_{c}^{\rm ch}(\tau):(0,T)\rightarrow\mathbb{R}$ be defined as in (4.4).
  • i)
    $x_{c}^{\rm ch}$ is strictly increasing on $(0,T)$.
  • ii)
    ${\lim_{\tau\to 0^{+}}x_{c}^{\rm ch}(\tau)}=\max\{\bar{x},{x_{c}}(0)\}$, where $\bar{x}$ is the point stated in (2.9).
Figure 3. The free boundaries $x_{c}^{\rm ch}(\tau)$ and ${x}_{p}^{\rm ch}(\tau)$.
Figure 3 illustrates the behavior of the two free boundaries, $x_{c}^{\rm ch}(\tau)$ and $x_{p}^{\rm ch}(\tau)$, for $V(\tau,x)$. In this figure, we can observe that, as shown in Theorems 4.4 and 4.5, $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$ are strictly decreasing and strictly increasing in $\tau\in[0,T]$, respectively. Moreover, we can also verify the inequality (4.5), which was discussed earlier.
Remark.
By the method introduced in [18] we can obtain that the free boundaries are Lipschitz continuous, i.e. $x_{c}^{\rm ch}$ and $x_{p}^{\rm ch}$ are Lipschitz continuous on $(0,T)$. Applying the boundary Harnack inequality [13] in the Lipschitz domain $\mathcal{C}$ as in [5, Prop 5.37] , we deduce that the free boundaries are of the class $C^{1,\alpha}$ for some $\alpha\in(0,1)$. Applying the higher order boundary Harnack inequality [23] further, we conclude that the free boundary is smooth.
Appendix
In Appendix, we provide the proof of Lemma 2.1, and show that $\mathcal{F}$ in the proof of Theorem 3.1 satisfies the conditions of the Schauder fixed point theorem.
Proof of Lemma 2.1
Similar results and approaches used in the proof can be found in previous research; see, e.g., [27, 28]. Therefore, we shall omit the details of approximations. We first prove Part (i). Let $\tilde{C}^{A}_{n,\varepsilon}$ and $\tilde{P}^{A}_{n,\varepsilon}$ be the unique solutions of the following equations with Dirichlet boundary conditions:
(4.7) $\begin{cases}\partial_{\zeta}\tilde{C}^{A}_{n,\varepsilon}-\mathcal{L}\tilde{C}^{A}_{n,\varepsilon}+\beta_{c,\varepsilon}(\tilde{C}^{A}_{n,\varepsilon}-\varphi_{\varepsilon}(e^{x}-K_{c}))=0\quad\text{in }\ {\Omega}^{n}_{T_{c}},\\ \tilde{C}^{A}_{n,\varepsilon}(\zeta,-n)=0\quad\text{and}\quad\tilde{C}^{A}_{n,\varepsilon}(\zeta,n)=e^{n}-K_{c}\quad\text{for }\ \zeta\in[0,T_{c}],\\ \tilde{C}^{A}_{n,\varepsilon}(0,x)=\varphi_{\varepsilon}(e^{x}-K_{c})\quad\text{for }\ x\in(-n,n)\end{cases}$
and
(4.8) $\begin{cases}\partial_{\zeta}\tilde{P}^{A}_{n,\varepsilon}-\mathcal{L}\tilde{P}^{A}_{n,\varepsilon}+\beta_{p,\varepsilon}(\tilde{P}^{A}_{n,\varepsilon}-\varphi_{\varepsilon}(K_{p}-e^{x}))=0\quad\text{in }\ {\Omega}^{n}_{T_{p}},\\ \tilde{P}^{A}_{n,\varepsilon}(\zeta,-n)=K_{p}-e^{-n}\quad\text{and}\quad\tilde{P}^{A}_{n,\varepsilon}(\zeta,n)=0\quad\text{for }\ \zeta\in[0,T_{p}],\\ \tilde{P}^{A}_{n,\varepsilon}(0,x)=\varphi_{\varepsilon}(K_{p}-e^{x})\quad\text{for }\ x\in(-n,n).\end{cases}$
where $\beta_{c,\varepsilon}$ and $\beta_{p,\varepsilon}$ are appropriate penalty functions, ${\Omega}^{n}_{T_{c}}:=(0,T_{c})\times(-n,n)$ and ${\Omega}^{n}_{T_{p}}:=(0,T_{p})\times(-n,n)$. The existence and uniqueness of the solutions to (4.7) and (4.8) is guaranteed by the Schauder’s fixed-point theorem. Then $\tilde{C}^{A}_{n}:=\lim_{\varepsilon\to 0^{+}}\tilde{C}^{A}_{n,\varepsilon}$ and $\tilde{P}^{A}_{n}:=\lim_{\varepsilon\to 0^{+}}\tilde{P}^{A}_{n,\varepsilon}$ are well-defined and solutions to the following obstacle problems:
$\begin{cases}\partial_{\zeta}\tilde{C}^{A}_{n}(\zeta,x)-\mathcal{L}\tilde{C}^{A}_{n}(\zeta,x)\geqslant 0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{c}}\ \ \text{with }\ \tilde{C}^{A}_{n}(\zeta,x)=(e^{x}-K_{c})^{+},\\ \partial_{\zeta}\tilde{C}^{A}_{n}(\zeta,x)-\mathcal{L}\tilde{C}^{A}_{n}(\zeta,x)=0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{c}}\ \ \text{with }\ \tilde{C}^{A}_{n}(\zeta,x)>(e^{x}-K_{c})^{+},\\ \tilde{C}^{A}_{n}(\zeta,-n)=0\quad\text{and}\quad\tilde{C}^{A}_{n}(\zeta,n)=e^{n}-K_{c}\quad\text{for }\ \zeta\in[0,T_{c}],\\ \tilde{C}^{A}_{n}(0,x)=(e^{x}-K_{c})^{+}\quad\text{for }\ x\in(-n,n)\end{cases}$
and
$\begin{cases}\partial_{\zeta}\tilde{P}^{A}_{n}(\zeta,x)-\mathcal{L}\tilde{P}^{A}_{n}(\zeta,x)\geqslant 0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{p}}\ \ \text{with }\ \tilde{P}^{A}_{n}(\zeta,x)=(K_{p}-e^{x})^{+},\\ \partial_{\zeta}\tilde{P}^{A}_{n}(\zeta,x)-\mathcal{L}\tilde{P}^{A}_{n}(\zeta,x)=0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{p}}\ \ \text{with }\ \tilde{P}^{A}_{n}(\zeta,x)>(K_{p}-e^{x})^{+},\\ \tilde{P}^{A}_{n}(\zeta,-n)=K_{p}-e^{-n}\quad\text{and}\quad\tilde{P}^{A}_{n,\varepsilon}(\zeta,n)=0\quad\text{for }\ \zeta\in[0,T_{p}],\\ \tilde{P}^{A}_{n}(0,x)=(K_{p}-e^{x})^{+}\quad\text{for }\ x\in(-n,n).\end{cases}$
Observe that $e^{x}+2$ satisfies
$\begin{cases}\partial_{\zeta}(e^{x}+2)-\mathcal{L}(e^{x}+2)+\beta_{c,\varepsilon}(e^{x}+2-\varphi_{\varepsilon}(e^{x}-K_{c}))>0\quad\text{for }\ (\zeta,x)\in\Omega^{n}_{T_{c}},\\ e^{-n}+2>0=\tilde{C}^{A}_{n,\varepsilon}(\zeta,-n)\quad\text{and}\quad e^{n}+2>e^{n}-K_{c}=\tilde{C}^{A}_{n,\varepsilon}(\zeta,n)\quad\text{for }\ \zeta\in[0,T_{c}],\\ e^{x}+2>\varphi_{\varepsilon}(e^{x}-K_{c})=\tilde{C}^{A}_{n,\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n)\end{cases}$
since $\beta_{c,\varepsilon}\in C^{\infty}(\mathbb{R})$ is a penalty function satisfying $\beta_{c,\varepsilon}(t)=0$ for all $t\geqslant\varepsilon$ and $e^{x}+2>\varphi_{\varepsilon}(e^{x}-K_{c})+\varepsilon$ holds for all $\varepsilon<1.$ By the comparison principle [6, 52p],
(4.9) $\tilde{C}^{A}_{n,\varepsilon}\leqslant e^{x}+2\ \ \text{for all}\ \ \varepsilon\in(0,1),\ \ \text{hence}\ \ \tilde{C}^{A}_{n}\leqslant e^{x}+2\quad\text{in}\ \ \Omega^{n}_{T_{c}}.$
Similarly, $K_{p}+2$ satisfies
$\begin{cases}\partial_{\zeta}(K_{p}+2)-\mathcal{L}(K_{p}+2)+\beta_{p,\varepsilon}(K_{p}+2-\varphi_{\varepsilon}(K_{p}-e^{x}))>0\quad\text{for }\ (\zeta,x)\in\Omega^{n}_{T_{p}},\\ K_{p}+2>K_{p}-e^{-n}=\tilde{P}^{A}_{n,\varepsilon}(\zeta,-n)\quad\text{and}\quad K_{p}+2>0=\tilde{P}^{A}_{n,\varepsilon}(\zeta,n)\quad\text{for }\ \zeta\in[0,T_{p}],\\ K_{p}+2>\varphi_{\varepsilon}(K_{p}-e^{x})=\tilde{P}^{A}_{n,\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
Again by the comparison principle [6, 52p],
(4.10) $\tilde{P}^{A}_{n,\varepsilon}\leqslant K_{p}+2\ \ \text{for all}\ \ \varepsilon\in(0,1),\ \ \text{hence}\ \ \tilde{P}^{A}_{n}\leqslant K_{p}+2\quad\text{in}\ \ \Omega^{n}_{T_{p}}.$
Moreover, for each $R>0$, ${\lim_{n\to\infty}}\tilde{C}^{A}_{n}$ and ${\lim_{n\to\infty}}\tilde{P}^{A}_{n}$ satisfy (2.5) and (2.6) in the domains $[-R,R]\times(0,T_{c})$ and $[-R,R]\times(0,T_{p})$ respectively. Since the problems (2.5) and (2.6) have unique solutions, taking $n\rightarrow\infty$ in (4.9) and (4.10) yield the results in Part (i).
We next prove Part (ii). Let $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ be the unique solutions to the following equations with Neumann boundary conditions:
(4.11) $\begin{cases}\partial_{\zeta}\widehat{C}^{A}_{n,\varepsilon}-\mathcal{L}\widehat{C}^{A}_{n,\varepsilon}+\beta_{c,\varepsilon}(\widehat{C}^{A}_{n,\varepsilon}-\varphi_{\varepsilon}(e^{x}-K_{c}))=0\quad\text{in }\ {\Omega}^{n}_{T_{c}},\\ \partial_{x}\widehat{C}^{A}_{n,\varepsilon}(\zeta,-n)=0\quad\text{and}\quad\partial_{x}\widehat{C}^{A}_{n,\varepsilon}(\zeta,n)=e^{n}\quad\text{for }\ \zeta\in[0,T_{c}],\\ \widehat{C}^{A}_{n,\varepsilon}(0,x)=\varphi_{\varepsilon}(e^{x}-K_{c})\quad\text{for }\ x\in(-n,n)\end{cases}$
and
(4.12) $\begin{cases}\partial_{\zeta}\widehat{P}_{n,\varepsilon}^{A}-\mathcal{L}\widehat{P}_{n,\varepsilon}^{A}+\beta_{p,\varepsilon}(\widehat{P}^{A}_{n,\varepsilon}-\varphi_{\varepsilon}(K_{p}-e^{x}))=0\quad\text{in }\ {\Omega}^{n}_{T_{p}},\\ \partial_{x}\widehat{P}_{n,\varepsilon}^{A}(\zeta,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}\widehat{P}_{n,\varepsilon}^{A}(\zeta,n)=0\quad\text{for }\ \zeta\in[0,T_{p}],\\ \widehat{P}_{n,\varepsilon}^{A}(0,x)=\varphi_{\varepsilon}(K_{p}-e^{x})\quad\text{for }\ x\in(-n,n).\end{cases}$
Then $\widehat{C}^{A}_{n}:=\lim_{\varepsilon\to 0^{+}}\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n}:=\lim_{\varepsilon\to 0^{+}}\widehat{P}^{A}_{n,\varepsilon}$ are well-defined and solutions to the following obstacle problems:
$\begin{cases}\partial_{\zeta}\widehat{C}^{A}_{n}(\zeta,x)-\mathcal{L}\widehat{C}^{A}_{n}(\zeta,x)\geqslant 0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{c}}\ \ \text{with }\ \widehat{C}^{A}_{n}(\zeta,x)=(e^{x}-K_{c})^{+},\\ \partial_{\zeta}\widehat{C}^{A}_{n}(\zeta,x)-\mathcal{L}\widehat{C}^{A}_{n}(\zeta,x)=0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{c}}\ \ \text{with }\ \widehat{C}^{A}_{n}(\zeta,x)>(e^{x}-K_{c})^{+},\\ \partial_{x}\widehat{C}^{A}_{n}(\zeta,-n)=0\quad\text{and}\quad\partial_{x}\widehat{C}^{A}_{n}(\zeta,n)=e^{n}\quad\text{for }\ \zeta\in[0,T_{c}],\\ \widehat{C}^{A}_{n}(0,x)=(e^{x}-K_{c})^{+}\quad\text{for }\ x\in(-n,n)\end{cases}$
and
$\begin{cases}\partial_{\zeta}\widehat{P}^{A}_{n}(\zeta,x)-\mathcal{L}\widehat{P}^{A}_{n}(\zeta,x)\geqslant 0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{p}}\ \ \text{with }\ \widehat{P}^{A}_{n}(\zeta,x)=(K_{p}-e^{x})^{+},\\ \partial_{\zeta}\widehat{P}^{A}_{n}(\zeta,x)-\mathcal{L}\widehat{P}^{A}_{n}(\zeta,x)=0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{p}}\ \ \text{with }\ \widehat{P}^{A}_{n}(\zeta,x)>(K_{p}-e^{x})^{+},\\ \partial_{x}\widehat{P}^{A}_{n}(\zeta,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}\widehat{P}^{A}_{n,\varepsilon}(\zeta,n)=0\quad\text{for }\ \zeta\in[0,T_{p}],\\ \widehat{P}^{A}_{n}(0,x)=(K_{p}-e^{x})^{+}\quad\text{for }\ x\in(-n,n).\end{cases}$
Note that by $W^{1,2}_{p}$-regularity, $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ are $W^{1,2}_{p}$ functions for all $1<p<\infty$ in the domain $\Omega_{T_{c}}$ and $\Omega_{T_{p}}$, respectively. Thus, the Sobolev embedding theorem implies that $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ are $C^{\frac{\alpha}{2},\alpha}$ functions for some $\alpha\in(0,1)$. Based on the Schauder theory for linear parabolic equations and a bootstrap argument, we deduce that the functions $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ are smooth. Thus, differentiating (4.11) and (4.12) yields that $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}$ and $\partial_{x}\widehat{P}^{A}_{n,\varepsilon}$ satisfies
$\begin{cases}\partial_{\zeta}(\partial_{x}\widehat{C}^{A}_{n,\varepsilon})-\mathcal{L}(\partial_{x}\widehat{C}^{A}_{n,\varepsilon})+\beta^{\prime}_{c,\varepsilon}(\cdots)\partial_{x}\widehat{C}^{A}_{n,\varepsilon}=\beta^{\prime}_{c,\varepsilon}(\cdots)\varphi_{\varepsilon}^{\prime}(e^{x}-K_{c})e^{x}\geqslant 0\quad\text{in }\ {\Omega}^{n}_{T_{c}},\\ \partial_{x}\widehat{C}^{A}_{n,\varepsilon}(\zeta,-n)=0\quad\text{and}\quad\partial_{x}\widehat{C}^{A}_{n,\varepsilon}(\zeta,n)=e^{n}\geqslant 0\quad\text{for }\ \zeta\in[0,T_{c}],\\ \partial_{x}\widehat{C}^{A}_{n,\varepsilon}(0,x)=\varphi^{\prime}_{\varepsilon}(e^{x}-K_{c})e^{x}\geqslant 0\quad\text{for }\ x\in(-n,n)\end{cases}$
and
$\begin{cases}\partial_{\zeta}(\partial_{x}\widehat{P}_{n,\varepsilon}^{A})-\mathcal{L}(\partial_{x}\widehat{P}_{n,\varepsilon}^{A})+\beta^{\prime}_{p,\varepsilon}(\cdots)\partial_{x}\widehat{P}_{n,\varepsilon}^{A}=-\beta^{\prime}_{p,\varepsilon}(\cdots)\varphi_{\varepsilon}^{\prime}(K_{p}-e^{x})e^{x}\leqslant 0\quad\text{in }\ {\Omega}^{n}_{T_{p}},\\ \partial_{x}\widehat{P}_{n,\varepsilon}^{A}(\zeta,-n)=-e^{-n}\leqslant 0\quad\text{and}\quad\partial_{x}\widehat{P}_{n,\varepsilon}^{A}(\zeta,n)=0\quad\text{for }\ \zeta\in[0,T_{p}],\\ \partial_{x}\widehat{P}_{n,\varepsilon}^{A}(0,x)=-\varphi^{\prime}_{\varepsilon}(K_{p}-e^{x})e^{x}\leqslant 0\quad\text{for }\ x\in(-n,n)\end{cases}$
respectively. Applying the maximum principle for each equation, we deduce $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}\geqslant 0$ in ${\Omega}^{n}_{T_{c}}$ and $\partial_{x}\widehat{P}_{n,\varepsilon}^{A}\leqslant 0$ in ${\Omega}^{n}_{T_{p}}$. Furthermore, since $-\mathcal{L}e^{x}+\beta^{\prime}_{c,\varepsilon}(\cdots)e^{x}\geqslant 0$, we obtain
$\begin{cases}\partial_{\zeta}(\partial_{x}\widehat{C}^{A}_{n,\varepsilon}-e^{x})-\mathcal{L}(\partial_{x}\widehat{C}^{A}_{n,\varepsilon}-e^{x})+\beta^{\prime}_{c,\varepsilon}(\cdots)(\partial_{x}\widehat{C}^{A}_{n,\varepsilon}-e^{x})\leqslant 0\quad\text{in }\ {\Omega}^{n}_{T_{c}},\\ \partial_{x}\widehat{C}^{A}_{n,\varepsilon}(\zeta,-n)-e^{-n}=-e^{-n}\leqslant 0\quad\text{and}\quad\partial_{x}\widehat{C}^{A}_{n,\varepsilon}(\zeta,n)-e^{n}=0\quad\text{for }\ \zeta\in[0,T_{c}],\\ \widehat{C}^{A}_{n,\varepsilon}(0,x)-e^{x}=\{\varphi^{\prime}_{\varepsilon}(e^{x}-K_{c})-1\}e^{x}\leqslant 0\quad\text{for }\ x\in(-n,n)\end{cases}$
and
$\begin{cases}\partial_{\zeta}(\partial_{x}\widehat{P}_{n,\varepsilon}^{A}+e^{x})-\mathcal{L}(\partial_{x}\widehat{P}_{n,\varepsilon}^{A}+e^{x})+\beta^{\prime}_{p,\varepsilon}(\cdots)(\partial_{x}\widehat{P}_{n,\varepsilon}^{A}+e^{x})\geqslant 0\quad\text{in }\ {\Omega}^{n}_{T_{p}},\\ \partial_{x}\widehat{P}_{n,\varepsilon}^{A}(\zeta,-n)+e^{-n}=0\quad\text{and}\quad\partial_{x}\widehat{P}_{n,\varepsilon}^{A}(\zeta,n)+e^{n}=e^{n}\geqslant 0\quad\text{for }\ \zeta\in[0,T_{p}],\\ \widehat{P}_{n,\varepsilon}^{A}(0,x)+e^{x}=\{1-\varphi^{\prime}_{\varepsilon}(K_{p}-e^{x})\}e^{x}\geqslant 0\quad\text{for }\ x\in(-n,n).\end{cases}$
Therefore, by the maximum principle, we deduce $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}\leqslant e^{x}$ and $\partial_{x}\widehat{P}_{n,\varepsilon}^{A}\geqslant-e^{x}$. Note that $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}$ and $\partial_{x}\widehat{P}^{A}_{n,\varepsilon}$ converge to $\partial_{x}\widehat{C}^{A}_{n}$ and $\partial_{x}\widehat{P}^{A}_{n}$ uniformly in $\Omega_{T_{c}}^{n}$ and $\Omega_{T_{p}}^{n}$, respectively, as $\varepsilon\rightarrow 0^{+}$. Hence $0\leqslant\partial_{x}\widehat{C}^{A}_{n}\leqslant e^{x}$ in $\Omega_{T_{c}}^{n}$ and $-e^{x}\leqslant\partial_{x}\widehat{P}^{A}_{n}\leqslant 0$ in $\Omega_{T_{p}}^{n}$. Finally, passing $n\to\infty$, we obtain the desired inequalities for $\partial_{x}\widehat{C}^{A}$ and $\partial_{x}\widehat{P}^{A}$.
To prove Part (iii), for each fixed $\delta>0$, we first define $\tilde{C}_{n}^{\delta}$ by
$\tilde{C}_{n}^{\delta}(\zeta,x):=\tilde{C}_{n}^{A}(\zeta+\delta,x)$
for all $(\zeta,x)\in\Omega_{T_{c}-\delta}$. Then, we see that $\tilde{C}_{n}^{\delta}$ satisfies the following obstacle problem:
$\begin{cases}\partial_{\zeta}\tilde{C}_{n}^{\delta}(\zeta,x)-\mathcal{L}\tilde{C}_{n}^{\delta}(\zeta,x)\geqslant 0\quad\text{for}\ \ (\zeta,x)\in{\Omega}^{n}_{T_{c}-\delta}\ \ \text{with }\ \tilde{C}_{n}^{\delta}(\zeta,x)=(e^{x}-K_{c})^{+},\\ \partial_{\zeta}\tilde{C}_{n}^{\delta}(\zeta,x)-\mathcal{L}\tilde{C}_{n}^{\delta}(\zeta,x)=0\quad\text{for}\ \ (\zeta,x)\in{\Omega}^{n}_{T_{c}-\delta}\ \ \text{with }\ \tilde{C}_{n}^{\delta}(\zeta,x)>(e^{x}-K_{c})^{+},\\ \tilde{C}_{n}^{\delta}(\zeta,-n)=0\quad\text{and}\quad\tilde{C}_{n}^{\delta}(\zeta,n)=e^{n}-K_{c}\quad\text{for }\ \zeta\in[0,T_{c}-\delta],\\ \tilde{C}_{n}^{\delta}(0,x)=\tilde{C}^{A}_{n}(\delta,x)\geqslant(e^{x}-K_{c})^{+}\quad\text{for }\ x\in(-n,n).\end{cases}$
Since $\tilde{C}_{n}^{\delta}(0,x)\geqslant(e^{x}-K_{c})^{+}=\tilde{C}^{A}_{n}(0,x)$, by the comparison principle for the variational inequality, (see [7, 80p, problem 5]) we deduce
$\tilde{C}^{A}_{n}(\zeta+\delta,x)=\tilde{C}_{n}^{\delta}(\zeta,x)\geqslant\tilde{C}^{A}_{n}(\zeta,x)$
for all $(\zeta,x)\in\Omega^{n}_{T_{c}-\delta}$. Take $n\rightarrow\infty$ to get
$\tilde{C}^{A}(\zeta+\delta,x)\geqslant\tilde{C}^{A}(\zeta,x)$
for all $(\zeta,x)\in\Omega_{T_{c}-\delta}$. Since $\delta>0$ is arbitrary we conclude $\partial_{\tau}C\geqslant 0$. Similarly, we can obtain $\partial_{\tau}P\geqslant 0$.
The proof for Part (iv) can be found in [22]. ∎
Conditions of Schauder fixed point theorem
We show that the function of the operator $\mathcal{F}:\mathcal{D}\to\mathcal{B}$ in the proof of Theorem (3.1) satisfies
  • (1)
    $\mathcal{F}(\mathcal{D})\subset\mathcal{D}$;
  • (2)
    $\mathcal{F}$ is continuous;
  • (3)
    $\mathcal{F}(\mathcal{D})$ is precompact in $\mathcal{B}$.
(1) Fix any $w\in\mathcal{D}$ and consider $u=\mathcal{F}(w).$ Note that $J_{\varepsilon}(\tau,x)\geqslant 0,$ $-\beta_{\varepsilon}(\cdots)\geqslant 0$ and $\ \nu\cdot D_{x}u(\tau,x)\leqslant 0$ for each $(\tau,x)$ at the lateral boundary and the inward pointing normal vector $\nu$ at $(\tau,x).$ Then by the comparison principle [14, 13p], we have $u\geqslant 0.$
(2) To obtain the continuity of $\mathcal{F},$ it suffices to prove the following:
[…]
$\text{Then}\ \displaystyle{\lim_{j\to\infty}\mathcal{F}(w_{j})}=\mathcal{F}(w).$
Let $u_{j}:=\mathcal{F}(w_{j})\ \text{for each }j\in\mathbb{N}\ \text{and}\ u:=\mathcal{F}(w).$ By subtracting the two equations that $u$ and $u_{j}$ satisfy respectively, we have
$\begin{cases}\partial_{\tau}(u-u_{j})-\mathcal{L}(u-u_{j})=-\beta_{\varepsilon}^{\prime}(\cdots)\ (w-w_{j})\quad\text{in }\ \Omega_{T}^{n},\\ \partial_{x}(u-u_{j})(\tau,-n)=0\quad\text{and}\quad\partial_{x}(u-u_{j})(\tau,n)=0\quad\text{for }\ \tau\in[0,T],\\ (u-u_{j})(0,x)=0\quad\text{for }\ x\in(-n,n).\end{cases}$
Then the $W^{1,2}_{p}$ estimate combined with the embedding theorem implies that $\|u-u_{j}\|_{C(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}\|\beta_{\varepsilon}^{\prime}(\cdots)(w-w_{j})\|_{L^{\infty}(\Omega_{T}^{n})}$ for some constant $\mathcal{K}>0$. Therefore, it follows that ${\lim_{j\to\infty}u_{j}}=u.$
(3) Let $\{u_{j}\}_{j=1}^{\infty}$ be a sequence in the closure of $\mathcal{F}(\mathcal{D})$ with respect to the $\|\cdot\|_{C(\overline{\Omega_{T}^{n}})}$ norm.
Case 1. For each $u_{j}\in\mathcal{F}(\mathcal{D})$, there exists a sequence of functions $\{w_{j}\}_{j=1}^{\infty}$ such that $\mathcal{F}(w_{j})=u_{j}$. By applying the $W^{1,2}_{p}$ estimate on $u_{j}$, we obtain
$\|u_{j}\|_{W^{1,2}_{p}(\Omega^{n}_{T})}$ $\leqslant\mathcal{K}(\|J_{\varepsilon}\|_{W^{2}_{p}((-n,n))}+\|\beta_{\varepsilon}(w_{j}-J_{\varepsilon})\|_{L^{p}(\Omega^{n}_{T})}+\|e^{n}+e^{-n}\|_{W^{1}_{p}((0,T))})$
$\leqslant\mathcal{K}(\|J_{\varepsilon}\|_{W^{2}_{p}((-n,n))}+\|\beta_{\varepsilon}(-J_{\varepsilon})\|_{L^{p}(\Omega^{n}_{T})}+\|e^{n}+e^{-n}\|_{W^{1}_{p}((0,T))})$
for some constant $\mathcal{K}>0$ that does not depend on $j\in\mathbb{N}$ and $\alpha\in(0,1)$. Note that the last inequality is due to the monotonicity of $\beta_{\varepsilon}$ and $w\geqslant 0$. Thus, combined with the embedding theorem it can be seen that
$\|u_{j}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}^{\prime}$
for some constant $\mathcal{K}^{\prime}$ that does not depend on $j\in\mathbb{N}$ and $\alpha\in(0,1)$.
Case 2. For each fixed $u_{i}\notin\mathcal{F}(\mathcal{D})$, there exists a sequence $\{u^{(i)}_{l}\}_{l=1}^{\infty}\subset\mathcal{F}(\mathcal{D})$ such that
$u^{(i)}_{l}\rightarrow u_{i}\quad\text{in }\ C(\overline{\Omega_{T}^{n}})\ \text{ as }\ l\rightarrow\infty.$
Hence, there exists $\{w^{(i)}_{l}\}_{l=1}^{\infty}$ in $\mathcal{D}$ such that $\mathcal{F}(w^{(i)}_{l})=u^{(i)}_{l}$. Again, by the $W^{1,2}_{p}$ estimate, the embedding theorem, and the monotonicity of $\beta_{\varepsilon}$, we deduce $\|u_{l}^{(i)}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}^{\prime}$ for some $\alpha\in(0,1)$. Take $l\rightarrow\infty$ to get $\|u_{i}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}^{\prime}$.
Combining both cases, we conclude
$\|u_{j}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}^{\prime}$
for all $j\in\mathbb{N}$. By the Arzela-Ascoli theorem there exists a subsequence $\{u_{j_{k}}\}_{k=1}^{\infty}\subset\{u_{j}\}_{j=1}^{\infty}$ and $u\in C(\overline{\Omega_{T}^{n}})$ such that
$u_{j_{k}}\rightarrow u\quad\text{in }\ C(\overline{\Omega_{T}^{n}})\quad\text{as }\ k\rightarrow\infty.$
In particular, $u$ is contained in the closure of $\mathcal{F}(\mathcal{D})$ by the assumption. This proves that $\mathcal{F}(\mathcal{D})$ is precompact in $\mathcal{B}=C(\overline{\Omega_{T}^{n}}).$
Acknowledgment
J. Jeon was supported by NRF grant funded by MSIT (RS-2023-00212648). J.Ok was supported by NRF grant funded by MSIT (NRF-2022R1C1C1004523).