### v3

# The Obstacle Problem Arising from the American Chooser Option

Abstract.
We study the obstacle problem associated with the American chooser option. The obstacle is given by the maximum of an American call option and an American put option, which, in turn, can be expressed as the maximum of the solutions to the corresponding obstacle problems. This structure makes the obstacle problem particularly challenging and non-trivial. Using theoretical analysis, we overcome these difficulties and establish the existence and uniqueness of a strong solution. Furthermore, we rigorously prove the monotonicity and smoothness of the free boundary arising from the obstacle problem.
Key words and phrases:
American chooser option, Free boundary problem, Optimal stopping, Parabolic obstacle problem
2020 Mathematics Subject Classification:
35R35, 35K85, 60G40, 91G80
1. Introduction
In this paper, our primary objective is to analyze the following (lower) obstacle problem derived from the American chooser option:
[…]
where the differential operator $\mathfrak{L}$ is defined as
$\mathfrak{L}:=\frac{\sigma^{2}}{2}s^{2}\partial_{ss}+(r-q)s\partial_{s}-r,\quad\text{with}\quad r,\;\sigma>0,\;q\geqslant 0,$
$D_{\eta}:=\{(t,s)\;:\;0<t<\eta,\ 0<s<\infty\}$ for $\eta>0$, and the functions $C^{A}$ and $P^{A}$ satisfy the following obstacle problems, respectively:
[…]
To ensure well-posedness, we assume:
$0<T<\min\{T_{c},T_{p}\},\quad 0<K_{p}<K_{c}.$
The existence and uniqueness of strong solutions to (1.2) and (1.3) are well-established, as shown in [26, 27].
The American chooser option allows the holder to exercise it as either an American call or an American put option, depending on which is more valuable at the time of exercise. This flexibility makes it particularly useful in uncertain market conditions where large price movements are anticipated. Consequently, it shares similarities with the American strangle option and includes it as a special case.
From a mathematical perspective, pricing an American chooser option leads to a parabolic obstacle problem, where the obstacle function is determined as the maximum of two separate solutions—one corresponding to the American call and the other to the American put. Unlike standard American options, in which the obstacle function is explicitly defined and well-behaved, the American chooser option introduces additional complexity due to the implicit and non-standard nature of the obstacle. This fundamental distinction makes the problem significantly more challenging to analyze within a mathematical framework.
Standard American options have been extensively studied in the literature (e.g., [28, 29, 27, 12]). Most of these studies however rely on obstacle problems where the obstacle function has an explicit, closed-form representation. In contrast, the American chooser option involves an obstacle function that depends on the interaction between two separate obstacle problems. This interaction gives rise to a new level of mathematical difficulty, as the obstacle is no longer explicitly known but rather emerges as the maximum of two distinct solutions. More precisely, unlike the existing literature, our obstacle is given by the maximum of two strong solutions to separate obstacle problems. Therefore, one cannot directly apply the operator to the obstacle and compute it in a pointwise manner as in standard problems with an explicit payoff, and this feature makes the analysis highly challenging.
Earlier studies on the American chooser option, such as [7, 21], adopted probabilistic approaches. In contrast, we present a rigorous analysis within a PDE framework, which, to the best of our knowledge, has not been previously explored. Our work focuses on the associated obstacle problem and develops a detailed PDE-based methodology. To address the challenges mentioned above, we carefully analyze the structure and behavior of the solutions to the two underlying obstacle problems, and as a result, we are able to construct a framework to study the American chooser option.
Specifically, following the methodology in [12], we introduce a penalized problem to approximate the original obstacle problem. Given the irregular and implicit nature of the obstacle function in this setting, a regularized version of the problem is considered. By leveraging the properties of the obstacle function and employing advanced PDE techniques, including the comparison principle, we demonstrate that the penalized problem yields a uniformly bounded solution. Importantly, while the obstacle in [12] is explicitly defined, the obstacle in our case depends on two interacting obstacle problems. Despite this complexity, we show that strong solutions can still be constructed, enabling the application of the comparison principle and rigorous analysis.
A crucial part of our analysis concerns the structure of the free boundaries. Owing to the nature of the American chooser option, the associated obstacle problem gives rise to two time-dependent free boundaries. More precisely, the two components of the obstacle coincide with the relevant solutions on two separate and disjoint regions, similarly to [12]. However, unlike [12], the obstacle is not explicitly known pointwise throughout the domain. The key observation is that these regions are contained in the respective exercise regions of the American put and call options. As a consequence, contact occurs only along the smooth payoff parts $e^{x}-K_{c}$ and $K_{p}-e^{x}$ of the obstacle, which allows us to employ standard arguments to establish the monotonicity and smoothness of the free boundaries.
Our paper is structured as follows. Section 2 provides the financial background and formal definition of the American chooser option and demonstrates the properties of the American call and put options. Section 3 establishes the existence and uniqueness of the obstacle problem solution using the penalty method. Section 4 examines the monotonicity and smoothness of the two free boundaries that arise in this framework.
2. Preliminaries
2.1. Notations
Throughout the paper, for each $\varepsilon>0$, $\varphi_{\varepsilon}\in C^{\infty}(\mathbb{R})$ is a function that satisfies
(2.1) $\begin{cases}\varphi_{\varepsilon}\geqslant 0,\quad 0\leqslant\varphi^{\prime}_{\varepsilon}\leqslant 1,\quad\varphi^{\prime\prime}_{\varepsilon}\geqslant 0,\\ \varphi_{\varepsilon}(\lambda)=\lambda\ \ \text{if}\ \ \lambda\geqslant\varepsilon,\quad\varphi_{\varepsilon}(\lambda)=0\ \ \text{if}\ \ \lambda\leqslant-\varepsilon,\\ \varphi_{\varepsilon}(\lambda)\leqslant(\lambda+\varepsilon)^{+}\ \ \text{for}\ \ \lambda\in\mathbb{R},\\ \displaystyle{\lim_{\varepsilon\to 0}\varphi_{\varepsilon}(\lambda)}=\lambda^{+}\ \ \text{uniformly for }\ \lambda\in\mathbb{R}.\end{cases}$
Let $\mathcal{D}\subset\mathbb{R}^{d+1}$ be a bounded parabolic domain, where $d\in\mathbb{N}$ denotes the dimension of the spatial variable $x$ and $t$ represents the time variable.
  • $C^{k+\frac{\alpha}{2},2k+\alpha}(\mathcal{D})$, $\alpha\in(0,1)$, $k\in\mathbb{N}$ is the Banach space under the following norm for $V$:
    $\begin{split}&\|V\|_{C^{k+\frac{\alpha}{2},2k+\alpha}(\mathcal{D})}\\ &\quad:=\sum_{i+2j=0,2,\cdots,2k}\limits\sup_{(t_{1},x_{1}),(t_{2},x_{2})\in\mathcal{D}}\{|D^{i}_{x}D^{j}_{t}V(t_{1},x_{1})|+\frac{|D^{i}_{x}D^{j}_{t}V(t_{1},x_{1})-D^{i}_{x}D^{j}_{t}V(t_{2},x_{2})|}{{|t_{1}-t_{2}|}^{\frac{\alpha}{2}}+{|x_{1}-x_{2}|}^{\alpha}}\}.\end{split}$
  • $L^{p}(\mathcal{D}),p\geqslant 1$ is the completion of $C^{\infty}(\mathcal{D})$ under the following norm for $V$:
    $\|V\|_{L^{p}(\mathcal{D})}:=\left(\iint_{\mathcal{D}}{|V(t,x)|}^{p}\,dx\,dt\right)^{\frac{1}{p}}.$
  • $W^{1,2}_{p}(\mathcal{D}),$ $p\geqslant 1$ is the completion of $C^{\infty}(\mathcal{D})$ under the following norm for $V$:
    $\|V\|_{W^{1,2}_{p}(\mathcal{D})}:=\left(\iint_{\mathcal{D}}\big\{|V|^{p}+|\partial_{t}V|^{p}+|\partial_{x}V|^{p}+|\partial^{2}_{x}V|^{p}\big\}\,dx\,dt\right)^{\frac{1}{p}}.$
Let $(t_{0},x_{0})\in\mathcal{D}$ and $r>0$.
  • $Q((t_{0},x_{0}),r)$ is the cylinder such that
    $Q((t_{0},x_{0}),r):=\{(t,x)\in\mathcal{D}:\text{max}\{|x-x_{0}|,|t-t_{0}|^{\frac{1}{2}}\}<r,t<t_{0}\}.$
  • $\mathcal{P}\mathcal{D}$ is the parabolic boundary of $\mathcal{D}$ which is defined to be the set of all points $(t_{0},x_{0})\in\partial{\mathcal{D}}$ such that for any $r>0,$ the cylinder $Q((t_{0},x_{0}),r)$ contains points not in $\mathcal{D}$}.
  • $\mathcal{BD}$ is the set of all points $(t,x)\in\mathcal{P}\mathcal{D}$ such that there is a positive $r>0$ with
    $Q((t+r^{2},x),r)\subset\mathcal{D}.$
2.2. Financial Background: American Chooser Option
Let $(\Omega,\mathcal{G},\mathbb{P})$ be a filtered probability space with a filtration $\{\mathcal{G}_{t}\}_{t\geqslant 0}$ satisfying the usual conditions. Under the risk-neutral measure $\mathbb{Q}$, the stock price $S_{t}$ follows a geometric Brownian motion (GBM) given by
$dS_{t}=(r-q)S_{t}dt+\sigma S_{t}dW_{t}^{\mathbb{Q}},\quad S_{0}>0,$
where $r>0$ is the constant risk-free interest rate, $q\geqslant 0$ is the continuous dividend rate, $\sigma>0$ is the volatility of the stock, and $W_{t}^{\mathbb{Q}}$ is a standard Brownian motion under the risk-neutral measure $\mathbb{Q}$. Throughout this paper, we assume $q>0$ for simplicity. The case $q=0$ leads to a degenerate situation with only one free boundary, but similar results can be obtained (see [12]).
To define the American chooser option, we first introduce the American call and put options. These two options serve as the fundamental building blocks for the American chooser option. An American call option with strike price $K_{c}$ and expiration time $T_{c}$ gives the holder the right to buy the underlying asset at price $K_{c}$ at any time $\theta\in\mathcal{U}_{t,T_{c}}$, where $\mathcal{U}_{t,T}$ is the set of all $\mathcal{G}$-stopping times taking values in $[t,T]$. The value $C^{A}(t,S_{t})$ of the American call option at time $t$ is given by
(2.2) $C^{A}(t,S_{t})=\sup_{\theta\in\mathcal{U}_{t,T_{c}}}\mathbb{E}^{\mathbb{Q}}\left[e^{-r(\theta-t)}(S_{\theta}-K_{c})^{+}\mid\mathcal{G}_{t}\right].$
Similarly, an American put option with the strike price $K_{p}<K_{c}$ and expiration time $T_{p}$ grants the holder the right to sell the underlying asset at price $K_{p}$ at any time $\tau\in[0,T_{p}]$. The value $P^{A}(t,S_{t})$ of the American put option at time $t$ is given by
(2.3) $P^{A}(t,S_{t})=\sup_{\theta\in\mathcal{U}_{t,T_{p}}}\mathbb{E}^{\mathbb{Q}}\left[e^{-r(\theta-t)}(K_{p}-S_{\theta})^{+}\mid\mathcal{G}_{t}\right].$
If $C^{A}$ and $P^{A}$ are strong solutions to the obstacle problems (1.2) and (1.3), respectively, then one can show, by applying Itô’s lemma for Sobolev spaces (see [14]), that they solve the corresponding optimal stopping problems (2.2) and (2.3); see, for instance, [5, Appendix B].
Under these circumstances, the value of the American chooser option, denoted by $V^{\rm ch}$, is defined as the solution to the following optimal stopping problem:
(2.4) $V^{\rm ch}(t,S_{t})=\sup_{\theta\in\mathcal{U}_{t,T}}\mathbb{E}\left[e^{-r(\theta-t)}\max\left\{C^{A}(\theta,S_{\theta}),P^{A}(\theta,S_{\theta})\right\}\mid\mathcal{G}_{t}\right].$
Using dynamic programming and Itô’s lemma, we can easily derive the obstacle problem (1.1) from (2.4) (see Appendix A.).
2.3. Change of variables
We transform them into forward non-degenerate problems and denote the solutions by $\widehat{C}^{A}$ and $\widehat{P}^{A}$, respectively. More precisely, we introduce the following transformation:
$\widehat{C}^{A}(\zeta,x)=C^{A}(T_{c}-\zeta,e^{x})\quad\text{and}\quad\widehat{P}^{A}(\zeta,x)=P^{A}(T_{p}-\zeta,e^{x}).$
Moreover, we define the domain $\Omega_{\eta}$ for a given constant $\eta>0$ as
$\Omega_{\eta}:=\{(\zeta,x)\mid 0<\zeta<\eta,\;x\in\mathbb{R}\}.$
Then, $\widehat{C}^{A}$ and $\widehat{P}^{A}$ satisfy the following obstacle problems, respectively:
(2.5) $\begin{cases}\partial_{\zeta}\widehat{C}^{A}(\zeta,x)-\mathcal{L}\widehat{C}^{A}(\zeta,x)\geqslant 0\quad\text{for }\ (\zeta,x)\in\Omega_{T_{c}}\ \text{ with }\ \widehat{C}^{A}(\zeta,x)=(e^{x}-K_{c})^{+},\\ \partial_{\zeta}\widehat{C}^{A}(\zeta,x)-\mathcal{L}\widehat{C}^{A}(\zeta,x)=0\quad\text{for }\ (\zeta,x)\in\Omega_{T_{c}}\ \text{ with }\ \widehat{C}^{A}(\zeta,x)>(e^{x}-K_{c})^{+},\\ \widehat{C}^{A}(0,x)=(e^{x}-K_{c})^{+}\quad\text{for }\ x\in\mathbb{R},\end{cases}$
(2.6) $\begin{cases}\partial_{\zeta}\widehat{P}^{A}(\zeta,x)-\mathcal{L}\widehat{P}^{A}(\zeta,x)\geqslant 0\quad\text{for }\ (\zeta,x)\in\Omega_{T_{p}}\ \text{ with }\ \widehat{P}^{A}(\zeta,x)=(K_{p}-e^{x})^{+},\\ \partial_{\zeta}\widehat{P}^{A}(\zeta,x)-\mathcal{L}\widehat{P}^{A}(\zeta,x)=0\quad\text{for }\ (\zeta,x)\in\Omega_{T_{p}}\ \text{ with }\ \widehat{P}^{A}(\zeta,x)>(K_{p}-e^{x})^{+},\\ \widehat{P}^{A}(0,x)=(K_{p}-e^{x})^{+}\quad\text{for }\ x\in\mathbb{R},\end{cases}$
where the differential operator $\mathcal{L}$ is given by
$\mathcal{L}:=\frac{\sigma^{2}}{2}\partial_{xx}+(r-q-\frac{\sigma^{2}}{2})\partial_{x}-r.$
According to the vast literature on American options (see, for instance, [1, 12, 13, 19, 24, 26, 27]), the obstacle problems (2.5) and (2.6) have unique strong solutions, $\widehat{C}^{A}\in W^{1,2}_{p\textup{,}\,\text{loc}}(\Omega_{T_{c}})\cap C(\overline{\Omega_{T_{c}}})$ and $\widehat{P}^{A}\in W^{1,2}_{p\textup{,}\,\text{loc}}(\Omega_{T_{p}})\cap C(\overline{\Omega_{T_{p}}})$, respectively.
We again perform the change of variables by setting
$V(\tau,x):=V^{\rm ch}(T-\tau,e^{x}).$
Then, $V(\tau,x)$ satisfies the following forward non-degenerate obstacle problem:
(2.7) $\begin{cases}\partial_{\tau}V(\tau,x)-\mathcal{L}V(\tau,x)\geqslant 0\quad\text{for }\ (\tau,x)\in\Omega_{T}\ \text{ with }\ V(\tau,x)=\max\{C(\tau,x),\;P(\tau,x)\},\\ \partial_{\tau}V(\tau,x)-\mathcal{L}V(\tau,x)=0\quad\text{for }\ (\tau,x)\in\Omega_{T}\ \text{ with }\ V(\tau,x)>\max\{C(\tau,x),\;P(\tau,x)\},\\ V(0,x)=\max\{C(0,x),P(0,x)\}\quad\text{for }\ x\in\mathbb{R},\end{cases}$
where
(2.8) $\begin{cases}C(\tau,x):=\widehat{C}^{A}(T_{c}-T+\tau,x)=C^{A}(T-\tau,e^{x}),\\ P(\tau,x):=\widehat{P}^{A}(T_{p}-T+\tau,x)=P^{A}(T-\tau,e^{x}),\end{cases}\quad(\tau,x)\in\overline{\Omega_{T}}.$
2.4. Properties of American Call and Put Options
The payoff function of the American chooser option takes the maximum form of the American call option $C^{A}$ and the American put option $P^{A}$. Therefore, the properties of $C^{A}$ and $P^{A}$ are essential for the analysis of the American chooser option. Since the properties of $C^{A}$ and $P^{A}$ naturally extend to $\widehat{C}^{A}$ and $\widehat{P}^{A}$, we briefly summarize the well-known properties of the American call and American put options in terms of $\widehat{C}^{A}$ and $\widehat{P}^{A}$ for convenience.
In each obstacle problem (2.5) and (2.6), it is well known that contact with the respective obstacles occurs only at their positive parts. Thus, we define the exercise regions, denoted by $\mathcal{E}_{C}$ and $\mathcal{E}_{P}$, and the continuation regions, denoted by $\mathcal{C}_{C}$ and $\mathcal{C}_{P}$, as follows:
$\mathcal{E}_{C}:=\{(\zeta,x)\in\Omega_{T_{c}}\,:\,\widehat{C}^{A}(\zeta,x)=e^{x}-K_{c}\},\quad\mathcal{C}_{C}:=\{(\zeta,x)\in\Omega_{T_{c}}\,:\,\widehat{C}^{A}(\zeta,x)>(e^{x}-K_{c})^{+}\},$
$\mathcal{E}_{P}:=\{(\zeta,x)\in\Omega_{T_{p}}\,:\,\widehat{P}^{A}(\zeta,x)=K_{p}-e^{x}\},\quad\mathcal{C}_{P}:=\{(\zeta,x)\in\Omega_{T_{p}}\,:\,\widehat{P}^{A}(\zeta,x)>(K_{p}-e^{x})^{+}\}.$
Then, the free boundaries $\hat{x}_{c}(\tau)$ for the American call and $\hat{x}_{p}(\tau)$ for the American put are well-defined as follows:
$\hat{x}_{c}(\tau):=\partial\mathcal{E}_{C}=\inf\{x\in\mathbb{R}\,:\,(\tau,x)\in\mathcal{E}_{C}\}\quad\text{and}\quad\hat{x}_{p}(\tau):=\partial\mathcal{E}_{P}=\sup\{x\in\mathbb{R}\,:\,(\tau,x)\in\mathcal{E}_{P}\}.$
It is well known that $\hat{x}_{c}(\tau)$ and $\hat{x}_{p}(\tau)$ satisfy the following properties (see [2, 4, 6, 20, 26, 27]):
  • $\hat{x}_{p}(\tau)$ is a smooth and strictly decreasing function for $\tau\in(0,T_{p}]$.
  • $\hat{x}_{c}(\tau)$ is a smooth and strictly increasing function for $\tau\in(0,T_{c}]$.
  • As $\tau$ approaches 0, the limiting behaviors of $\hat{x}_{c}(\tau)$ and $\hat{x}_{p}(\tau)$ are given by
    $\lim_{\tau\to 0^{+}}\hat{x}_{c}(\tau)=\ln{\left(\max\left\{1,\tfrac{r}{q}\right\}K_{c}\right)}\quad\text{and}\quad\lim_{\tau\to 0^{+}}\hat{x}_{p}(\tau)=\ln{\left(\min\left\{1,\tfrac{r}{q}\right\}K_{p}\right)},$
    where $q>0$. If $q=0$, there does not exist the free boundary $\hat{x}_{c}(\tau)$ and $\lim_{\tau\to 0^{+}}\hat{x}_{p}(\tau)=\ln{K_{p}}$.
Additionally, the following property can also be obtained by [21].
  • There exists a unique $\bar{x}>0$ such that
    (2.9) $\widehat{C}^{A}(T_{c}-T,\bar{x})=\widehat{P}^{A}(T_{p}-T,\bar{x}),\quad\text{or equivalently}\quad C(0,\bar{x})=P(0,\bar{x}).$
According to the properties of the free boundaries of the American call and put options described above, since $K_{c}>K_{p}$, the following holds, regardless of the relationship between $T_{c}$ and $T_{p}$:
$\hat{x}_{c}(\tau)>\hat{x}_{p}(\tau),\quad\quad\forall\;\tau\geqslant 0.$
In other words, $\mathcal{E}_{C}$ and $\mathcal{E}_{P}$ are always disjoint. Figure 1 illustrates the behaviors of the free boundaries $\hat{x}_{c}(\tau)$ and $\hat{x}_{p}(\tau)$.
(a) the free boundary $\hat{x}_{c}(\tau)$
(b) the free boundary $\hat{x}_{p}(\tau)$
Figure 1. The free boundaries $\hat{x}_{c}(\tau)$ and $\hat{x}_{p}(\tau)$.
In terms of $C$ and $P$, we also define $x_{c}(\tau)$ and $x_{p}(\tau)$ as follows:
$x_{c}(\tau):=\hat{x}_{c}(T_{c}-T+\tau),\quad\text{and}\quad x_{p}(\tau):=\hat{x}_{p}(T_{p}-T+\tau).$
For further discussion, it is necessary to verify various properties of C{C} and P{P}. Since C{C} and P{P} are restricted functions of $\widehat{C}^{A}$ and $\widehat{P}^{A}$, it is natural to analyze $\widehat{C}^{A}$ and $\widehat{P}^{A}$ over their original whole domains and then conclude that the same properties hold for C{C} and P{P}, respectively. The following results are essentially well-known and their proofs are provided in the Appendix.
Lemma 2.1.
Let $\widehat{C}^{A}$ and $\widehat{P}^{A}$ be the solution to (2.5) and (2.6), respectively, and define for each $\lambda\in\mathbb{R}$ that $\Omega_{\lambda}:=(0,\lambda)\times\mathbb{R}$. Then the following properties hold.
  1. (i)
    $0\leqslant\widehat{C}^{A}\leqslant e^{x}+2$ in $\Omega_{T_{c}}$,  $0\leqslant\widehat{P}^{A}\leqslant K_{p}+2$ in $\Omega_{T_{p}}$.
  2. (ii)
    $0\leqslant\partial_{x}\widehat{C}^{A}\leqslant e^{x}$ in $\Omega_{T_{c}}$,  $e^{x}\leqslant\partial_{x}\widehat{P}^{A}\leqslant 0$ in $\Omega_{T_{p}}$.
  3. (iii)
    $\partial_{\zeta}\widehat{C}^{A}\geqslant 0$ in $\Omega_{T_{c}}$,  $\partial_{\zeta}\widehat{P}^{A}\geqslant 0$ in $\Omega_{T_{p}}$.
  4. (iv)
    For each $\tau\in(0,T]$, there exists $x_{\tau}\in\mathbb{R}$ such that $C(\tau,x)>P(\tau,x)$ for all $x>x_{\tau}$.
Moreover, by (2.8), the corresponding statements hold for $C$ and $P$ in $\Omega_{T}$.
3. Existence and uniqueness of solution
In this section, we aim to prove the existence and uniqueness of the solution to the obstacle problem (2.7).
3.1. Existence and uniqueness in a bounded region
We first consider the following obstacle problem in the bounded region $\Omega_{T}^{n}:=(0,T)\times(-n,n)$ for each $n\in\mathbb{N}$ with Neumann boundary condition:
(3.1) $\begin{cases}\partial_{\tau}V_{n}(\tau,x)-\mathcal{L}V_{n}(\tau,x)\geqslant 0\quad\text{for}\ \ (\tau,x)\in\Omega_{T}^{n}\ \ \text{with}\ \ V_{n}(\tau,x)=J(\tau,x),\\ \partial_{\tau}V_{n}(\tau,x)-\mathcal{L}V_{n}(\tau,x)=0\quad\text{for}\ \ (\tau,x)\in\Omega_{T}^{n}\ \ \text{with}\ \ V_{n}(\tau,x)>J(\tau,x),\\ \partial_{x}V_{n}(\tau,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}V_{n}(\tau,n)=e^{n}\quad\text{for }\ \tau\in[0,T],\\ V_{n}(0,x)=J(0,x)\quad\text{for }\ x\in(-n,n),\end{cases}$
where
$J(\tau,x):=\max\{C(\tau,x),P(\tau,x)\}\quad\text{ for\ all }\ (\tau,x)\in\Omega_{T}^{n}.$
We prove the existence and uniqueness of the solution to (3.1) in Theorem 3.3. To this end, we use the so called penalty method.
Define a penalty function $\beta_{\varepsilon}(\cdot)\in C^{\infty}(\mathbb{R})$ with $\varepsilon>0$ satisfying
(3.2) $\begin{cases}\beta_{\varepsilon}(\lambda)\leqslant 0,\ \beta^{\prime}_{\varepsilon}(\lambda)\geqslant 0\ \ \text{and}\ \ \beta^{\prime\prime}_{\varepsilon}(\lambda)\leqslant 0\ \ \text{for \ all}\ \ \lambda\in\mathbb{R},\\ \beta_{\varepsilon}(\lambda)=0\ \ \text{if}\ \ \lambda\geqslant\varepsilon,\ \beta_{\varepsilon}(0)=-\mathcal{K}_{0}\ \ \text{where}\ \ \mathcal{K}_{0}:=2\{(q+r)e^{n}+2rK_{c}+5r\},\\ \displaystyle{\lim_{\varepsilon\to 0}\beta_{\varepsilon}(\lambda)}=0\ \ \text{if}\ \ \lambda>0,\quad\displaystyle{\lim_{\varepsilon\to 0}\beta_{\varepsilon}(\lambda)}=-\infty\ \ \text{if}\ \ \lambda<0\end{cases}$
and $\varphi_{\varepsilon}(\cdot)\in C^{{\infty}}(\mathbb{R})$ satisfying (2.1). We then consider the following penalized problem;
(3.3) $\begin{cases}\partial_{\tau}V_{n,\varepsilon}-\mathcal{L}V_{n,\varepsilon}+\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})=0\quad\text{in }\ \Omega^{n}_{T},\\ \partial_{x}V_{n,\varepsilon}(\tau,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}V_{n,\varepsilon}(\tau,n)=e^{n}\quad\text{for }\ \tau\in[0,T],\\ V_{n,\varepsilon}(0,x)=J_{\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n),\end{cases}$
where
(3.4) $J_{\varepsilon}:=\varphi_{\varepsilon}(C-P)+P.$
Theorem 3.1.
For each fixed $n\in\mathbb{N}\setminus\{0\}$ with $n>\max\{|{x_{c}}(0)|,|{x_{p}}(0)|\}$, there exists a solution $V_{n,\varepsilon}\in W^{1,2}_{p}(\Omega_{T}^{n})\cap C(\overline{\Omega_{T}^{n}})$ to the problem (3.3), where $1<p<\infty.$
Proof.
We apply the Schauder fixed point theorem [11, 280p] to the following setting. Set $\mathcal{X}:=C(\overline{\Omega_{T}^{n}})$ and $\mathcal{A}:=\{w\in\mathcal{X}:w\geqslant 0\}$. Then $\mathcal{A}$ is a closed convex subset of the Banach space $\mathcal{X}.\$For each $w\in\mathcal{A}$, let $u\in W^{1,2}_{p}(\Omega_{T}^{n})$ be the solution to
(3.5) $\begin{cases}\partial_{\tau}u-\mathcal{L}u+\beta_{\varepsilon}(w-J_{\varepsilon})=0\quad\text{in }\ \Omega^{n}_{T},\\ \partial_{x}u(\tau,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}u(\tau,n)=e^{n}\quad\text{for }\ \tau\in[0,T],\\ u(0,x)=J_{\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
Note that the well-posedness of the solution to (3.5) and the relevant estimate in $W^{1,2}_{p}(\Omega_{T}^{n})$ for $1<p<\infty$ can be found in [15, Theorem 9.1, p.341]:
$\|u\|_{W^{1,2}_{p}(\Omega_{T}^{n})}\leqslant\mathcal{K}\left(T(e^{n}+e^{-n})+\|J_{\varepsilon}\|_{W^{2}_{p}((-n,n))}+\|\beta_{\varepsilon}(w-J_{\varepsilon})\|_{L^{p}(\Omega_{T}^{n})}\right)$
for some constant $\mathcal{K}>0$. Moreover, the parabolic Sobolev embedding theorem [15, Chapter II. Lemma 3.3] yields that $u$ is Hölder continuous in $\overline{\Omega_{T}^{n}}$. Thus $u\in\mathcal{X}.$ Now we define an operator $\mathcal{F}:\mathcal{A}\rightarrow\mathcal{X}$ by $\mathcal{F}(w):=u$. In order to use the Schauder fixed point theorem in $\mathcal{X},$ it suffices to show the following three properties:
  • (1)
    $\mathcal{F}(\mathcal{A})\subset\mathcal{A}$;
  • (2)
    $\mathcal{F}$ is continuous;
  • (3)
    $\mathcal{F}(\mathcal{A})$ is precompact in $\mathcal{X}$.
We shall prove it in Appendix D. Utilizing the Schauder fixed point theorem, we obtain the solution $V_{n,\varepsilon}$ of the problem (3.3). In particular, $V_{n,\varepsilon}\in W^{1,2}_{p}(\Omega^{n}_{T})\cap C(\overline{\Omega^{n}_{T}})$ for each $1<p<\infty.$
Next, we investigate the convergence of $V_{n,\varepsilon}$ as $\varepsilon\to 0^{+}$. We establish that the limit exists and is a solution to the obstacle problem (3.1).
Lemma 3.2.
Let $V_{n,\varepsilon}$ be the solution of (3.3). Then for each $n\in\mathbb{N},$ there exists some constant $\mathcal{K}=\mathcal{K}(n)>0$ independent of $\varepsilon\in(0,1)$ such that
$-\mathcal{K}\leqslant\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\leqslant 0\quad\text{and}\quad 0\leqslant V_{n,\varepsilon}\leqslant\mathcal{K}\quad\text{in }\ \Omega_{T}^{n}.$
Proof.
From Lemma 2.1 (ii), we note
$0\leqslant\partial_{x}C\leqslant e^{x}\ \ \text{and}\ \ -e^{x}\leqslant\partial_{x}P\leqslant 0\quad\text{in}\ \ \Omega^{n}_{T}.$
Next, observe that $\partial_{x}J_{\varepsilon}=\{1-\varphi^{\prime}_{\varepsilon}(C-P)\}\partial_{x}P+\varphi^{\prime}_{\varepsilon}(C-P)\partial_{x}C$ and $0\leqslant\varphi^{\prime}_{\varepsilon}(\cdot)\leqslant 1$. Combining these estimates, we deduce
(3.6) $e^{x}\leqslant\partial_{x}J_{\varepsilon}\leqslant e^{x}\quad\text{in}\ \ \Omega^{n}_{T}.$
Moreover, by Lemma 2.1 (i), we have
$|C-P|\leqslant C+P\leqslant e^{n}+K_{p}+4\quad\text{in}\ \ \Omega_{T}^{n}$
and since $C,P\in W^{1,2}_{p,\text{loc}}(\Omega_{T})$, it follows that almost everywhere in $\Omega_{T}^{n}$,
$\partial_{\tau}C-\mathcal{L}C=(qe^{x}-rK_{c})\mathbf{1}_{\{C=e^{x}-K_{c}\}}\quad\text{and}\quad\partial_{\tau}P-\mathcal{L}P=(-qe^{x}+rK_{p})\mathbf{1}_{\{P=K_{p}-e^{x}\}}.$
Direct computation and the choice of $\mathcal{K}_{0}$ yield the following in $\Omega_{T}^{n}$:
$\partial_{\tau}J_{\varepsilon}-\mathcal{L}J_{\varepsilon}$ $=\varphi^{\prime}_{\varepsilon}(C-P)\left(\partial_{\tau}C-\mathcal{L}C\right)+\left(1-\varphi^{\prime}_{\varepsilon}(C-P)\right)\left(\partial_{\tau}P-\mathcal{L}P\right)$
$\quad-\frac{\sigma^{2}}{2}\varphi^{\prime\prime}_{\varepsilon}(C-P)\left[\partial_{x}(C-P)\right]^{2}-r\varphi^{\prime}_{\varepsilon}(C-P)(C-P)+r\varphi_{\varepsilon}(C-P)$
$\leqslant\varphi^{\prime}_{\varepsilon}(\cdot)|qe^{n}-rK_{c}|+(1-\varphi^{\prime}_{\varepsilon}(\cdot))|-qe^{n}+rK_{p}|$
$\quad+2r\left(\max_{\Omega_{T}^{n}}|C-P|+1\right)$
$\leqslant 2(qe^{n}+rK_{c})+2r(e^{n}+K_{p}+5)$
$\leqslant 2(q+r)e^{n}+4rK_{c}+10r$
$=\mathcal{K}_{0}=-\beta_{\varepsilon}(0)$
for sufficiently small $0<\varepsilon<1$. Consequently, we deduce
$\begin{cases}\partial_{\tau}J_{\varepsilon}-\mathcal{L}J_{\varepsilon}+\beta_{\varepsilon}(0)\leqslant 0\quad\text{in }\ \Omega^{n}_{T},\\ \partial_{x}J_{\varepsilon}(\tau,-n)\geqslant-e^{-n}=\partial_{x}V_{n,\varepsilon}(\tau,-n)\quad\text{and}\\ \partial_{x}J_{\varepsilon}(\tau,n)\leqslant e^{n}=\partial_{x}V_{n,\varepsilon}(\tau,n)\quad\text{for }\ \tau\in[0,T],\\ J_{\varepsilon}(0,x)=V_{n,\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
Hence, $J_{\varepsilon}$ is a subsolution to (3.3) and by the comparison principle [8, Theorem 17, p.53]
$V_{n,\varepsilon}\geqslant J_{\varepsilon}\geqslant 0\quad\text{in }\ \Omega_{T}^{n}.$
Moreover, the monotonicity of $\beta_{\varepsilon}$ implies
$\mathcal{K}_{0}=\beta_{\varepsilon}(0)\leqslant\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\leqslant 0.$
We next show that $V_{n,\varepsilon}$ is bounded from above. To this end, define
$\overline{V}(\tau,x):=e^{x}+e^{-n}x^{2}+\mathcal{K}_{1},$
where $\mathcal{K}_{1}$ is a constant satisfying
$\mathcal{K}_{1}\geqslant\max\left\{\bigg(1-\frac{q}{r}-\frac{\sigma^{2}}{2r}\right)^{2}+\frac{\sigma^{2}}{r},\;K_{p}+6\bigg\}.$
This implies $\overline{V}\geqslant J_{\varepsilon}+\varepsilon$ and hence $\beta_{\varepsilon}(\overline{V}-J_{\varepsilon})=0$ in $\Omega_{T}^{n}.$ Moreover in $\Omega_{T}^{n}$,
$\partial_{\tau}\overline{V}-\mathcal{L}\overline{V}=qe^{x}+e^{-n}\{rx^{2}-(2r-2q-\sigma^{2})x-\sigma^{2}\}+r\mathcal{K}_{1}\geqslant 0.$
Thus, $\overline{V}$ satisfies
$\begin{cases}\partial_{\tau}\overline{V}-\mathcal{L}\overline{V}+\beta_{\varepsilon}(\overline{V}-J_{\varepsilon})\geqslant 0\quad\text{in }\ \Omega^{n}_{T},\\ \partial_{x}\overline{V}(\tau,-n)=(1-2n)e^{-n}\leqslant\partial_{x}V_{n,\varepsilon}(\tau,-n)\quad\text{and}\\ \partial_{x}\overline{V}(\tau,n)=2ne^{-n}+e^{n}\geqslant\partial_{x}V_{n,\varepsilon}(\tau,n)\quad\text{for }\ \tau\in[0,T],\\ \overline{V}(0,x)\geqslant V_{n,\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
By the comparison principle [8, Theorem 17, p.53],
$V_{n,\varepsilon}\leqslant\overline{V}\leqslant e^{n}+e^{-n}n^{2}+\mathcal{K}_{1}\quad\text{in}\ \ \Omega^{n}_{T}.\qed$
We now establish the existence of the solution $V_{n}$ to problem (3.1). To this end, let us recall the unique point $\bar{x}$ satisfying $C(0,\bar{x})=P(0,\bar{x})$ from (2.9). For each $\varrho>0$, we define the ball $B_{\varrho}(0,\bar{x}):=\{(\tau,x)\in\mathbb{R}^{2}:|(\tau,x)-(0,\bar{x})|<\varrho\}$.
Theorem 3.3.
For each fixed $n\in\mathbb{N}$ with $n>\max\{|{x_{c}}(0)|,|{x_{p}}(0)|\}$ and $\varrho>0$, there exists a solution $V_{n}\in W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))$ to the problem (3.1), where $1<p<\infty$. Moreover, $V_{n}\in C(\overline{\Omega^{n}_{T}})$.
Proof.
Since $J(0,x)=\max\{C(0,x),P(0,x)\}$ is not differentiable at the point $(0,\bar{x})$, we consider the following $W^{1,2}_{p}$-estimate in the domain $\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})$ for $1<p<\infty$ and $\varrho>0$:
$\|V_{n,\varepsilon}\|_{W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))}$ $\leqslant\mathcal{K}\Big(\|V_{n,\varepsilon}\|_{L^{\infty}(\Omega^{n}_{T})}+\|\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\|_{L^{\infty}(\Omega^{n}_{T})}+T(e^{n}+e^{-n})$
$\quad\quad+\|\varphi_{\varepsilon}(C-P)+P\|_{W^{2}_{p}((-n,n)\setminus(\bar{x}-\varrho/2,\bar{x}+\varrho/2))}\Big)$
$\leqslant\mathcal{K}^{\prime}.$
By the Sobolev embedding theorem, we deduce that for every $(\tau_{0},x_{0})\in\overline{\Omega_{T}^{n}}\setminus\{(0,\bar{x})\}$, $V_{n,\varepsilon}$ is $C^{\frac{\alpha}{2},\alpha}$ in some neighborhood of $(\tau_{0},x_{0})$ in $\overline{\Omega_{T}^{n}}$. Note that the constant $\mathcal{K}^{\prime}$ is independent of $\varepsilon\in(0,\varrho/2)$ due to Lemma 3.2 and the property of $\varphi_{\varepsilon}$ in (2.1). Therefore, for each large $n\in\mathbb{N}$, $V_{n,\varepsilon}$ is bounded in $W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))\cap C(\overline{\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})})$ for $\varepsilon\in(0,\varrho/2)$, and hence there exists $V_{n}\in W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))\cap C(\overline{\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})})$ for all $1<p<\infty$ and $\varrho>0$ such that
(3.7) $V_{n,\varepsilon}\rightharpoonup V_{n}\quad$ $\text{in}\quad W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})),$
(3.8) $V_{n,\varepsilon}\rightarrow V_{n}\quad$ $\text{in}\quad C(\overline{\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})}),$
up to a subsequence, as $\varepsilon\rightarrow 0^{+}$.
Now, we prove that $V_{n}$ satisfies (3.1). Observe from the inequality $V_{n,\varepsilon}\geqslant J_{\varepsilon}$, the uniform convergence $V_{n,\varepsilon}\to V_{n}$ and $J_{\varepsilon}\to J$ in $\Omega_{T}^{n}$, as $\varepsilon\to 0^{+}$, yields $V_{n}\geqslant J$ in $\Omega_{T}^{n}$. Fix any $\psi\in C_{c}^{\infty}(\{V_{n}>J\}\cap\Omega^{n}_{T})$. Then, there exist small $\varrho,\delta>0$ such that
$\text{supp}(\psi)\subset\{V_{n}>J\}\cap\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})\quad\text{and}\quad\min\limits_{\text{supp}(\psi)}(V_{n}-J)>2\delta.$
Choose $\varepsilon_{0}>0$ small so that for every $\varepsilon\in(0,\varepsilon_{0}]$,
$|(V_{n,\varepsilon}-J_{\varepsilon})-(V_{n}-J)|<\delta\quad\text{and}\quad\varepsilon<\delta/2.$
Then, for such small $\varepsilon>0$, it follows that $V_{n,\varepsilon}-J_{\varepsilon}>\delta>\varepsilon$, and this implies $\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})=0$ in $\text{supp}(\psi)$. Consequently,
$\iint_{\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})}\big(\partial_{\tau}V_{n,\varepsilon}-\mathcal{L}V_{n,\varepsilon}\big)\psi\,d\tau\,dx=0.$
By the weak convergence (3.7), taking $\varepsilon\rightarrow 0^{+}$ yields
$\iint_{\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})}\big(\partial_{\tau}V_{n}-\mathcal{L}V_{n}\big)\psi\,d\tau\,dx=0.$
Since $\psi\in C_{c}^{\infty}(\{V_{n}>J\}\cap\Omega^{n}_{T})$ is arbitrary, we deduce that $\partial_{\tau}V_{n}-\mathcal{L}V_{n}=0$ a.e. in $\{V_{n}>J\}\cap\Omega^{n}_{T}.$
To prove $\partial_{\tau}V_{n}-\mathcal{L}V_{n}\geqslant 0$ a.e. in $\Omega^{n}_{T}$, let us fix a non-negative $\psi\in C_{c}^{\infty}(\Omega^{n}_{T})$. Then since $\beta_{\varepsilon}(\cdot)\leqslant 0$, $V_{n,\varepsilon}$ satisfies that for all small $\varepsilon>0$,
$\iint_{\Omega^{n}_{T}}({\partial_{\tau}V_{n,\varepsilon}-\mathcal{L}V_{n,\varepsilon}})\psi\,d\tau\,dx=\iint_{\Omega^{n}_{T}}-\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\psi\,d\tau\,dx\geqslant 0.$
By the weak convergence (3.7), taking $\varepsilon\rightarrow 0^{+}$ yields
$\iint_{\Omega^{n}_{T}}({\partial_{\tau}V_{n}-\mathcal{L}V_{n}})\psi\,d\tau\,dx\geqslant 0.$
Since $\psi$ is an arbitrary non-negative test function in $C_{c}^{\infty}(\Omega_{T}^{n})$, we deduce that $\partial_{\tau}V_{n}-\mathcal{L}V_{n}\geqslant 0$ a.e. in $\Omega_{T}^{n}$. Combining all, we conclude that $V_{n}$ is a solution of (3.1).
Finally, by the convergence (3.8) for each $\varrho>0$ and the continuity of the initial datum $V_{n}(0,x)=\max\{C(0,x),P(0,x)\}$, we obtain that $V_{n}\in C(\overline{\Omega_{T}^{n}})$.
3.2. Solution to the obstacle problem
Finally, we prove the existence and uniqueness of the solution to the obstacle problem (2.7).
Theorem 3.4.
There exists a unique solution $V\in W^{1,2}_{p,\mathrm{loc}}(\Omega_{T})\cap C(\overline{\Omega_{T}})$ to (2.7), where $1<p<\infty$.
Proof.
For each $n\in\mathbb{N}$, let $V_{n}$ be the solution to the obstacle problem (3.1) constructed in the previous subsection. By the $W^{1,2}_{p}$-regularity of $V_{n}$, we have
$\begin{cases}\partial_{\tau}V_{n}-\mathcal{L}V_{n}=f_{n}\quad\text{a.e. in }\ \Omega_{T}^{n},\\ \partial_{x}V_{n}(\tau,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}V_{n}(\tau,n)=e^{n}\quad\text{for }\ \tau\in[0,T],\\ V_{n}(0,x)=\max\{C(0,x),P(0,x)\}\quad\text{for }\ x\in(-n,n),\end{cases}$
where
$f_{n}(\tau,x)=(qe^{x}-rK_{c})\mathbf{1}_{\{V_{n}=e^{x}-K_{c}\}\cap\{C>P\}}+(-qe^{x}+rK_{p})\mathbf{1}_{\{V_{n}=K_{p}-e^{x}\}\cap\{P>C\}}.$
Let us fix some $m\in\mathbb{N}$. Then for each $n>2m$, $f_{n}$ is uniformly bounded in $\Omega_{T}^{m}$, with a bound depending only on $m$ and independent of $n$. Then it follows from the interior $W^{1,2}_{p}$-estimate for $1<p<\infty$ in [15, p.355] that for each small $\varrho>0,$
$\|V_{n}\|_{W^{1,2}_{p}(\Omega_{T}^{m}\setminus B_{\varrho}(0,\bar{x}))}$ $\leqslant\mathcal{K}\Big(\|V_{n}\|_{L^{\infty}(\Omega_{T}^{2m})}+\|C(0,\cdot)\|_{W^{2}_{p}((-2m,2m))}$
$\quad+\|P(0,\cdot)\|_{W^{2}_{p}((-2m,2m))}+\|f_{n}\|_{L^{\infty}(\Omega_{T}^{2m})}\Big)$
$\leqslant\mathcal{K}^{\prime},$
for some constant $\mathcal{K}^{\prime}$ independent of $n$. Letting $n\to\infty$, we deduce that, up to a subsequence, there exists a function $V^{m}$ such that
$V_{n}\rightharpoonup V^{m}\quad\text{in }W^{1,2}_{p}(\Omega_{T}^{m}\setminus B_{\varrho}(0,\bar{x}))$
for all $1<p<\infty$ and $\varrho>0$. In addition, the Sobolev embedding theorem [15, Chapter II. Lemma 3.3] yields
$V_{n_{k}}\rightarrow V^{m}\ \text{ in }\ C(\overline{\Omega^{m}_{T}\setminus B_{\varrho}(0,\bar{x})}).$
Combined with the continuity of $J(0,x)=\max\{C(0,x),P(0,x)\}$, it follows that $V^{m}\in C(\overline{\Omega_{T}^{m}})$ for all $m\in\mathbb{N}$. By a standard diagonal argument, we can find a subsequence $\{V_{n_{k}}\}_{k=1}^{\infty}$ such that, as $k\rightarrow\infty$,
(3.9) $V_{n_{k}}\rightharpoonup V^{m}\ \$$\text{in }\ W^{1,2}_{p}(\Omega^{m}_{T}\setminus B_{\varrho}(0,\bar{x}))\quad\text{for\ all }\ 1<p<\infty\ \text{ and }\ \varrho>0$
for each large $m\in\mathbb{N}.$ Thus, $V^{m+1}=V^{m}$ in $\Omega^{m}_{T}$ so we can define $V:=V^{m}$ in $\overline{\Omega_{T}^{m}}$ for all large $m\in\mathbb{N}$.
We now proceed to prove that $V$ is a solution to (2.7). To this end, let us fix $m\in\mathbb{N}$ and show that $V^{m}$ satisfies
$\begin{cases}\partial_{\tau}V^{m}-\mathcal{L}V^{m}\geqslant 0&\text{for }(\tau,x)\in\Omega^{m}_{T}\text{ with }V^{m}(\tau,x)=J(\tau,x),\\ \partial_{\tau}V^{m}-\mathcal{L}V^{m}=0&\text{for }(\tau,x)\in\Omega^{m}_{T}\text{ with }V^{m}(\tau,x)>J(\tau,x),\\ V^{m}(0,x)=J(0,x)&\text{for }x\in(-m,m).\end{cases}$
Recall that for each $n\geqslant m$, $\partial_{\tau}V_{n}-\mathcal{L}V_{n}\geqslant 0$ a.e. in $\Omega_{T}^{m}$. Since $V_{n}\rightharpoonup V^{m}$ weakly in $W^{1,2}_{p}(\Omega_{T}^{m}\setminus B_{\varrho}(0,\bar{x}))$ for all $\varrho>0$, we observe that for any non-negative test function $\psi\in C_{c}^{\infty}(\Omega_{T}^{m})$,
$\iint_{\Omega_{T}^{m}}(\partial_{\tau}V^{m}-\mathcal{L}V^{m})\psi\,d\tau dx=\lim_{n\to\infty}\iint_{\Omega_{T}^{m}}(\partial_{\tau}V_{n}-\mathcal{L}V_{n})\psi\,d\tau dx\geqslant 0.$
Since $\psi\geqslant 0$ is arbitrary, it follows that $\partial_{\tau}V^{m}-\mathcal{L}V^{m}\geqslant 0$ a.e. in $\Omega_{T}^{m}$. Furthermore, the Sobolev embedding theorem implies that $V_{n}\to V^{m}$ in $C(\Omega_{T}^{m})$ as $n\to\infty$. Since $V_{n}\geqslant J$ in $\Omega_{T}^{m}$, passing to the limit yields
$V^{m}\geqslant J\quad\text{in }\Omega^{m}_{T}.$
Consequently, it remains to show that $\partial_{\tau}V^{m}-\mathcal{L}V^{m}=0$ on the set $\{V^{m}>J\}$.
Fix any $\psi\in C_{c}^{\infty}(\{V^{m}>J\}\cap\Omega^{m}_{T})$ and $\varrho>0$ such that $\text{supp}(\psi)\subset\{V^{m}>J\}\cap\Omega_{T}^{m}\setminus B_{\varrho}(0,\bar{x})$. Then, there exists $\delta>0$ such that
$\min\limits_{\text{supp}(\psi)}(V^{m}-J)>\delta.$
Let us fix $n\in\mathbb{N}$ and define $S_{n}^{(1)}$ and $S_{n}^{(2)}$ by
$S_{n}^{(1)}:=\{(\tau,x)\in\Omega^{m}_{T}:|(V_{n}-J)-(V^{m}-J)|<\delta/2\},$
$S_{n}^{(2)}:=\{(\tau,x)\in\Omega^{m}_{T}:|(V_{n}-J)-(V^{m}-J)|\geqslant\delta/2\}.$
Note that $V_{n}-J>\frac{\delta}{2}$ in $S_{n}^{(1)}$, and this implies
$S_{n}^{(1)}\subset\{(\tau,x)\in\Omega_{T}^{n}:V_{n}(\tau,x)>J(\tau,x)\}\cap\Omega^{m}_{T}.$
Next, recall that $\partial_{\tau}V_{n}-\mathcal{L}V_{n}=0$ holds almost everywhere in $\{(\tau,x)\in\Omega_{T}^{n}:V_{n}(\tau,x)>J(\tau,x)\}$. This leads to the conclusion that
$\iint_{S_{n}^{(1)}}({\partial_{\tau}V_{n}-\mathcal{L}V_{n}})\psi\,d\tau dx=0.$
Furthermore, the uniform convergence of $V_{n}-J$ to $V^{m}-J$ ensures that $V_{n}-J$ converges in measure as $n\to\infty$. In other words, the measure of the set $S_{n}^{(2)}$ tends to zero as $n\to\infty$. Together with the uniform estimate $\|V_{n_{k}}\|_{W^{2,1}_{p}(\Omega^{m}_{T}\setminus B_{\varrho}(0,\bar{x}))}\leqslant\mathcal{K}^{\prime}$, we deduce that
$\lim_{k\to\infty}\iint_{S_{n_{k}}^{(2)}}\left(\partial_{\tau}V_{n_{k}}-\mathcal{L}V_{n_{k}}\right)\psi\,d\tau dx=0.$
Combining the above two identities, we obtain
$\iint_{\Omega^{m}_{T}}({\partial_{\tau}V^{m}-\mathcal{L}V^{m}})\psi\,d\tau dx$ $=\displaystyle{\lim_{k\to\infty}}\iint_{S_{n_{k}}^{(1)}\cup S_{n_{k}}^{(2)}}({\partial_{\tau}V_{n_{k}}-\mathcal{L}V_{n_{k}}})\psi\,d\tau dx=0$
and thus, $\partial_{\tau}V^{m}-\mathcal{L}V^{m}=0$ a.e. in $\{V^{m}>J\}$. Since $m\in\mathbb{N}$ is arbitrary, we conclude that $V$ is a solution to (2.7).
For the uniqueness of a solution $V$ to (2.7), we notice that by [22, Remark 3], since $|e^{x}-K|<\max\{K,1\}e^{|x|}$ for $K>0$ and $x\in\mathbb{R}$, $|\max\{C,P\}|\leqslant M_{1}e^{|x|}$ for some $M_{1}>0$ and hence $|V|\leqslant M_{2}e^{|x|}$ for some $M_{2}>0$. Therefore, by [22, Theorem 3] we obtain the uniqueness of $V$. ∎
3.3. Estimate for the derivatives of V
Lemma 3.5.
Let $V$ be a solution to (2.7). Then
$\partial_{\tau}V\geqslant 0\quad\text{a.e.\ in }\ \Omega_{T}\quad\text{and}\quad-e^{x}\leqslant\partial_{x}V\leqslant e^{x}\quad\text{in }\ \Omega_{T}.$
Proof.
To prove the first inequality, fix some small $\delta>0$ and consider $V_{n,\varepsilon}(\tau+\delta,x)$ for all $(\tau,x)\in\Omega^{n}_{T-\delta}.$ Observe that $V_{n,\varepsilon}(\tau+\delta,x)$ satisfies
(3.10) $\begin{cases}\partial_{\tau}V_{n,\varepsilon}(\tau+\delta,x)-\mathcal{L}V_{n,\varepsilon}(\tau+\delta,x)\\ \quad\quad+\beta_{\varepsilon}(V_{n,\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau+\delta,x))=0\quad\text{for }\ (\tau,x)\in\Omega^{n}_{T-\delta},\\ \partial_{x}V_{n,\varepsilon}(\tau+\delta,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}V_{n,\varepsilon}(\tau+\delta,n)=e^{n}\quad\text{for }\ \tau\in[0,T-\delta],\end{cases}$
Subtracting (3.3) from (3.10) and the mean value theorem yield
$\begin{cases}\partial_{\tau}\{V_{n,\varepsilon}(\tau+\delta,x)-V_{n,\varepsilon}(\tau,x)\}-\mathcal{L}\{V_{n,\varepsilon}(\tau+\delta,x)-V_{n,\varepsilon}(\tau,x)\}\\ \ \ +\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})\{V_{n,\varepsilon}(\tau+\delta,x)-V_{n,\varepsilon}(\tau,x)\}=\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})\{J_{\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau,x)\}\ \ \text{for }(\tau,x)\in\Omega^{n}_{T-\delta},\\ \partial_{x}\{V_{n,\varepsilon}(\tau+\delta,-n)-V_{n,\varepsilon}(\tau,-n)\}=0\quad\text{and}\\ \qquad\qquad\qquad\partial_{x}\{V_{n,\varepsilon}(\tau+\delta,n)-V_{n,\varepsilon}(\tau,n)\}=0\quad\text{for }\ \tau\in[0,T-\delta],\\ V_{n,\varepsilon}(\delta,x)-V_{n,\varepsilon}(0,x)=V_{n,\varepsilon}(\delta,x)-J_{\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n),\end{cases}$
where $\gamma_{\tau,x}>0$ is a positive number in between $V_{n,\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau+\delta,x)\geqslant 0$ and $V_{n,\varepsilon}(\tau,x)-J_{\varepsilon}(\tau,x)\geqslant 0$, determined by the mean value theorem. In the case that the latter two values are equal, we choose $\gamma_{\tau,x}$ to be large so that $\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})=0$ (see (3.2)). Then from definition (3.2) it follows that $\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})$ is nonnegative and bounded for $(\tau,x)\in\Omega^{n}_{T-\delta}$. In addition, by (3.4), (2.1) and Lemma 2.1 (iii), we have
$(J_{\varepsilon})_{\tau}=\varphi^{\prime}_{\varepsilon}(C-P)(C_{\tau}-P_{\tau})+P_{\tau}\geqslant 0.$
Hence, the right-hand side of the first equation,
$\beta^{\prime}_{\varepsilon}(\gamma_{\tau,x})\{J_{\varepsilon}(\tau+\delta,x)-J_{\varepsilon}(\tau,x)\},$
is nonnegative for all $(\tau,x)\in\Omega^{n}_{T-\delta}$. Moreover, since $V_{n,\varepsilon}(\delta,x)\geqslant J_{\varepsilon}(\delta,x)$ and $V_{n,\varepsilon}(0,x)=J_{\varepsilon}(0,x)$ for all $x\in(-n,n)$, we deduce
$V_{n,\varepsilon}(\delta,x)-V_{n,\varepsilon}(0,x)\geqslant J_{\varepsilon}(\delta,x)-J_{\varepsilon}(0,x)\geqslant 0\quad\text{for }\ x\in(-n,n).$
Therefore, by the comparison principle,
$V_{n,\varepsilon}(\tau+\delta,x)\geqslant V_{n,\varepsilon}(\tau,x)\quad\text{for all}\ \ (\tau,x)\in\Omega^{n}_{T-\delta}.$
Take $\varepsilon\rightarrow 0$ and $n\rightarrow\infty$, the almost everywhere differentiability of $V$ with respect to $\tau$ yields the first estimate.
Next, from the $W^{1,2}_{p}$-estimate and the Sobolev embedding theorem for the solution $V_{n,\varepsilon}$ of (3.3), we obtain that $V_{n,\varepsilon}\in C^{\frac{\alpha}{2},\alpha}(\overline{\Omega^{n}_{T}})$ for some $\alpha\in(0,1)$. Thus, the right hand side $\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})$ belongs to $C^{\frac{\alpha}{2},\alpha}(\Omega^{n}_{T})$. By standard Schauder theory, we deduce that $V_{n,\varepsilon}\in C^{1+\frac{\alpha}{2},2+\alpha}(\Omega_{T}^{n})$. This enhances the regularity of the right-hand side, and thus, a subsequent application of Schauder theory yields $V_{n,\varepsilon}\in C^{2+\frac{\alpha}{2},4+\alpha}(\Omega_{T}^{n})$. Iterating this bootstrap argument, we conclude that $V_{n,\varepsilon}$ is smooth in $\Omega_{T}^{n}$. This smoothness justifies differentiating the equation in (3.3) with respect to $x$. Consequently, we see that $\partial_{x}V_{n,\varepsilon}$ satisfies
$\begin{cases}\partial_{\tau}(\partial_{x}V_{n,\varepsilon})-\mathcal{L}(\partial_{x}V_{n,\varepsilon})+\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})(\partial_{x}V_{n,\varepsilon})=\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\partial_{x}J_{\varepsilon}\quad\text{in }\ \Omega_{T}^{n},\\ \partial_{x}V_{n,\varepsilon}(\tau,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}V_{n,\varepsilon}(\tau,n)=e^{n}\quad\text{for }\ \tau\in[0,T],\\ \partial_{x}V_{n,\varepsilon}(0,x)=\partial_{x}J_{\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
Define a linear operator $\tilde{\mathcal{L}}$ by
$\tilde{\mathcal{L}}:=\mathcal{L}-\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})=\frac{\sigma^{2}}{2}\partial_{xx}+(r-q-\frac{\sigma^{2}}{2})\partial_{x}-(r+\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon}))$
Then, $\partial_{x}V_{n,\varepsilon}$ satisfies the equation $(\partial_{\tau}-\tilde{\mathcal{L}})(\partial_{x}V_{n,\varepsilon})=\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\partial_{x}J_{\varepsilon}$ in $\Omega^{n}_{T}$. Since $\mathcal{L}e^{x}=-qe^{x}$, with (3.6), we see that $\partial_{x}V_{n,\varepsilon}-e^{x}$ satisfies
$\begin{cases}\partial_{\tau}(\partial_{x}V_{n,\varepsilon}-e^{x})-\tilde{\mathcal{L}}(\partial_{x}V_{n,\varepsilon}-e^{x})=\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})(\partial_{x}J_{\varepsilon}-e^{x})-qe^{x}\leqslant 0\quad\text{in }\ \Omega_{T}^{n},\\ \partial_{x}V_{n,\varepsilon}(\tau,-n)-e^{-n}=-2e^{-n}\quad\text{and}\quad\partial_{x}V_{n,\varepsilon}(\tau,n)-e^{n}=0\quad\text{for }\ \tau\in[0,T],\\ \partial_{x}V_{n,\varepsilon}(0,x)-e^{x}=\partial_{x}J_{\varepsilon}(0,x)-e^{x}\leqslant 0\quad\text{for }\ x\in(-n,n).\end{cases}$
Note that the zeroth-order coefficient of $\tilde{\mathcal{L}}$ is nonnegative. Thus, by the maximum principle, we deduce $\partial_{x}V_{n,\varepsilon}\leqslant e^{x}$ in $\Omega_{T}^{n}$. Similarly, $\partial_{x}V_{n,\varepsilon}+e^{x}$ satisfies
$\begin{cases}\partial_{\tau}(\partial_{x}V_{n,\varepsilon}+e^{x})-\tilde{\mathcal{L}}(\partial_{x}V_{n,\varepsilon}+e^{x})=\beta^{\prime}_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})(\partial_{x}J_{\varepsilon}+e^{x})+qe^{x}\geqslant 0\quad\text{in }\ \Omega_{T}^{n},\\ \partial_{x}V_{n,\varepsilon}(\tau,-n)+e^{-n}=0\quad\text{and}\quad\partial_{x}V_{n,\varepsilon}(\tau,n)+e^{n}=2e^{n}\quad\text{for }\ \tau\in[0,T],\\ \partial_{x}V_{n,\varepsilon}(0,x)+e^{x}=\partial_{x}J_{\varepsilon}(0,x)+e^{x}\geqslant 0\quad\text{for }\ x\in(-n,n)\end{cases}$
and again by the maximum principle, it follows that $\partial_{x}V_{n,\varepsilon}\geqslant-e^{x}$ in $\Omega^{n}_{T}$. Note each weak convergence (3.7), (3.9) with $p>3$ and the Sobolev embedding theorem [15, Chapter II. Lemma 3.3] to see that $\partial_{x}V_{n,\varepsilon}\to\partial_{x}V$ uniformly as $\varepsilon\rightarrow 0^{+}$ and $n\rightarrow\infty$. Thus, taking $\varepsilon\to 0^{+}$ and $n\to\infty$ yields the result.
4. Analysis of the free boundary
4.1. Representation of the free boundary
Let us define the exercise region $\mathcal{E}$ and the continuation region $\mathcal{C}$ for the solution $V$ of (2.7) as
$\mathcal{E}:=$ $\{(\tau,x)\in\Omega_{T}:V(\tau,x)=J(\tau,x)\}\quad\text{and}\quad\mathcal{C}:=\{(\tau,x)\in\Omega_{T}:V(\tau,x)>J(\tau,x)\}.$
Furthermore, let us define $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$ as
$\mathcal{E}_{P}^{\rm ch}:=$ $\{(\tau,x)\in\Omega_{T}:V(\tau,x)=P(\tau,x)\}\quad\text{and}\quad\mathcal{E}_{C}^{\rm ch}:=$ $\{(\tau,x)\in\Omega_{T}:V(\tau,x)=C(\tau,x)\}.$
In the following lemma, we prove that $V$ contacts the obstacle exclusively along its smooth portions; that is, $V$ coincides with either $e^{x}-K_{c}$ or $K_{p}-e^{x}$ in the exercise region $\mathcal{E}$. This ensures that standard arguments remain applicable within our framework.
Furthermore, we establish that $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$ are mutually disjoint. Consequently, the free boundary $\partial\mathcal{E}$ decomposes into the disjoint union of $\partial\mathcal{E}_{P}^{\rm ch}$ and $\partial\mathcal{E}_{C}^{\rm ch}$. Therefore, it suffices to analyze $\partial\mathcal{E}_{P}^{\rm ch}$ and $\partial\mathcal{E}_{C}^{\rm ch}$ separately.
Lemma 4.1.
Let $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$ be defined as above. Then,
(4.1) $\mathcal{E}_{P}^{\rm ch}\subset\{P=K_{p}-e^{x}\}\quad\text{and}\quad\mathcal{E}_{C}^{\rm ch}\subset\{C=e^{x}-K_{c}\}.$
In particular, we have
(4.2) $\mathcal{E}_{P}^{\rm ch}=\{V=K_{p}-e^{x}\}\quad\text{and}\quad\mathcal{E}_{C}^{\rm ch}=\{V=e^{x}-K_{c}\},$
which further implies that $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$ are mutually disjoint.
Proof.
Suppose that there exists a point $(\tau_{0},x_{0})\in\{P>K_{p}-e^{x}\}\cap\mathcal{E}_{P}^{\rm ch}$. Since $\{P>K_{p}-e^{x}\}$ is open, there exists some $\delta>0$ such that
$Q:=(\tau_{0}-\delta,\tau_{0}]\times(x_{0}-\delta,\infty)\subset\{P>K_{p}-e^{x}\}$
(see Figure 1 (B)). Then, by the above inclusion, it follows that $\partial_{\tau}P-\mathcal{L}P=0$ in $Q$ and thus,
$\partial_{\tau}(V-P)-\mathcal{L}(V-P)\geqslant 0\quad\text{a.e. in }\ Q.$
Moreover, since $(\tau_{0},x_{0})\in\mathcal{E}_{P}^{\rm ch}$, VPV-P attains its minimum value 0 at the interior point $(\tau_{0},x_{0})\in Q$. Thus, by the strong maximum principle [18, Corollary 2.4], we deduce that $V\equiv P$ in $Q$. This is a contradiction since $C(\tau_{0},x)>P(\tau_{0},x)$ for all $x>x_{\tau_{0}}$ (see Lemma 2.1 (iv)) and $V\geqslant C$ in $\Omega_{T}$. Therefore,
$\{P>K_{p}-e^{x}\}\cap\mathcal{E}_{P}^{\rm ch}=\emptyset.$
In other words, $\mathcal{E}_{P}^{\rm ch}$ is contained in the set $\mathcal{E}_{P}=\{P=K_{p}-e^{x}\}$. Using a similar argument with $V$ and $C$, we can also obtain $\mathcal{E}_{C}^{\rm ch}\subset\mathcal{E}_{C}$.
Since $V\geqslant P\geqslant K_{p}-e^{x}$ and $V\geqslant C\geqslant e^{x}-K_{c}$, we have
$\{V=K_{p}-e^{x}\}\subset\mathcal{E}_{P}^{\rm ch}\quad\text{and}\quad\{V=e^{x}-K_{c}\}\subset\mathcal{E}_{C}^{\rm ch}.$
Combined with (4.1), these inclusions yield (4.2). Furthermore, since $\mathcal{E}_{P}$ and $\mathcal{E}_{C}$ are disjoint (see Subsection 2.4), the inclusions $\mathcal{E}_{P}^{\rm ch}\subset\mathcal{E}_{P}$ and $\mathcal{E}_{C}^{\rm ch}\subset\mathcal{E}_{C}$ immediately imply the disjointness of $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$. ∎
Lemma 4.2.
Let $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$ be defined as above. Then, $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$ are nonempty.
Proof.
We first observe from the increasing property of $x_{c}(\tau)$ and the inequality $x_{p}(\tau)<x_{c}(\tau)$ that $C(\tau,x)>e^{x}-K_{c}$ for all $(\tau,x)\in(0,T)\times(-\infty,x_{p}(0))$. Suppose, by contradiction, that $\mathcal{E}_{P}^{\rm ch}=\emptyset$. Choose $x_{0}\in\mathbb{R}$ and $\delta>0$ such that $x_{0}+3\delta<\min\{x_{p}(0),\bar{x}\}$, and set
$Q:=(0,\delta)\times(x_{0}-2\delta,x_{0}-\delta).$
Since $Q\subset\{C>e^{x}-K_{c}\}$ and the set $\{V=C\}$ is identical to $\{V=e^{x}-K_{c}\}$ by Lemma 4.1, it follows that $Q$ is disjoint from $\{V=C\}$. Furthermore, under the assumption that $\mathcal{E}_{P}^{\rm ch}=\emptyset$, the set $Q$ is necessarily contained within the continuation region $\mathcal{C}$. Consequently, we obtain $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$.
Define the bottom boundary of $Q$ as
$\mathcal{B}_{Q}:=\{0\}\times(x_{0}-2\delta,x_{0}-\delta).$
Then from the initial condition $V(0,x)=\max\{C(0,x),P(0,x)\}$ and $x_{0}<\min\{x_{p}(0),\bar{x}\}$, it follows that $V(0,x)=K_{p}-e^{x}$ on $\mathcal{B}_{Q}$. Moreover, since $V$ satisfies the equation $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$, we apply the result in [8, Theorem 19, p. 321] to $V-(K_{p}-e^{x})$ to deduce that $V$ is smooth in $Q$ and up to the initial boundary $\mathcal{B}_{Q}$. Therefore, passing to the limit as $\tau\to 0^{+}$ in $(\partial_{\tau}-\mathcal{L})(V-(K_{p}-e^{x}))=qe^{x}-rK_{p}$, we obtain for sufficiently small $\delta>0$ that
$\partial_{\tau}V\leqslant rK_{p}(e^{-\delta}-1)<0\quad\text{on }\mathcal{B}_{Q}.$
This contradicts the first estimate in Lemma 3.5. Hence, $\mathcal{E}_{P}^{\rm ch}\neq\emptyset$.
The same argument yields $\mathcal{E}_{C}^{\rm ch}\neq\emptyset$. ∎
To proceed further, recall from Subsection 2.4 that the $x$-coordinates of points in $\mathcal{E}_{C}$ and $\mathcal{E}_{P}$ satisfy $x>\ln K_{c}$ and $x<\ln K_{p}$, respectively. Combined with Lemma 4.1, the same bounds hold for points in $\mathcal{E}_{C}^{\rm ch}$ and $\mathcal{E}_{P}^{\rm ch}$, respectively. Therefore, using Lemma 3.5 and Lemma 4.1, we parameterize the free boundaries by
(4.3) $x_{p}^{\rm ch}(\tau)$ $:=\sup\limits_{x<\ln K_{p}}\ \{x:V(\tau,x)=K_{p}-e^{x}\}\quad\tau\in(0,T),$
(4.4) $x_{c}^{\rm ch}(\tau)$ $:=\inf\limits_{x>\ln K_{c}}\ \{x:V(\tau,x)=e^{x}-K_{c}\}\quad\tau\in(0,T).$
Remark 4.3.
Since $\mathcal{E}_{P}^{\rm ch}$ and $\mathcal{E}_{C}^{\rm ch}$ are nonempty, the set $\{x:V(\tau_{1},x)=K_{p}-e^{x}\}$ and $\{x:V(\tau_{2},x)=e^{x}-K_{c}\}$ are nonempty for some $\tau_{1},\ \tau_{2}\in(0,T)$. Moreover, the monotonicity of $V$ with respect to $\tau$, obtained in Lemma 3.5, implies that $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$ are monotone decreasing and increasing respectively.
Figure 2. solution $V(\tau,x)$ and obstacle $\max\{C(\tau,x),P(\tau,x)\}.$
Figure 2 illustrates the solution $V(\tau,x)$ of (2.7) and the obstacle $J(\tau,x)=\max\{C(\tau,x),P(\tau,x)\}$. From the figure, we observe that $V(\tau,x)$ and the obstacle meet in the regions where $C(\tau,x)=e^{x}-K_{c}$ and $P(\tau,x)=K_{p}-e^{x}$, which was proven in Lemma 4.1. Therefore, this confirms that
(4.5) $x_{p}^{\rm ch}(\tau)\leqslant x_{p}(\tau)<x_{c}(\tau)\leqslant x_{c}^{\rm ch}(\tau),\quad\quad\forall\;\tau\in[0,T].$
4.2. Properties of the free boundary
At the end of the previous subsection, we represented each free boundary by viewing their $x$-coordinate as the supremum and infimum of the points where the contact between the solution and the obstacle occurs. To see that such representations correspond to the actual free boundary, we need to establish that the parametrizations in (4.3) and (4.4) are continuous.
Lemma $4.$
Let $x_{p}^{\rm ch}(\tau),$ $x_{c}^{\rm ch}(\tau)$ be defined as in (4.3) and (4.4) respectively. Then $x_{p}^{\rm ch}$ and $x_{c}^{\rm ch}$ are continuous on $(0,T)$.
Proof.
Suppose that $x_{p}^{\rm ch}$ is not continuous at some point $\tau_{0}\in(0,T).$ Then by the monotonicity of $x_{p}^{\rm ch}(\tau),$ either ${\lim_{\tau\to{\tau_{0}}^{+}}x_{p}^{\rm ch}(\tau)}>{x_{p}^{\rm ch}}(\tau_{0})$ or ${\lim_{\tau\to{\tau_{0}}^{-}}x_{p}^{\rm ch}(\tau)}<{x_{p}^{\rm ch}}(\tau_{0})$. Let us consider the first case. Then, there exist some $\varepsilon>0$ and $\delta>0$ such that for all $\tau$ in the interval $(\tau_{0},\tau_{0}+\delta)$ we have
$x_{p}^{\rm ch}(\tau)-x_{p}^{\rm ch}(\tau_{0})\geqslant\varepsilon.$
Consider the cylinder
$Q:=(\tau_{0},\tau_{0}+\delta)\times\big(x_{p}^{\rm ch}(\tau_{0})-\tfrac{\varepsilon}{2},x_{p}^{\rm ch}(\tau_{0})\big).$
By construction, it can be seen that $Q$ is contained in the set $\{V>K_{p}-e^{x}\}$ and disjoint from $\{V=e^{x}-K_{c}\}$. Hence, $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$. Moreover, since $V=K_{p}-e^{x}$ on the bottom boundary $\mathcal{B}_{Q}:=\{\tau_{0}\}\times(x_{p}^{\rm ch}(\tau_{0})-\frac{\varepsilon}{2},x_{p}^{\rm ch}(\tau_{0}))$ of $Q$, we apply the result in [8, Theorem 19, p. 321] to $V-(K_{p}-e^{x})$ to deduce that $V$ is smooth in $Q$ up to the initial boundary $\mathcal{B}_{Q}$. Next, recall from the monotonicity of $x_{p}^{\rm ch}$ that $x_{p}^{\rm ch}(\tau_{0})<x_{p}^{\rm ch}(0)\leqslant\ln{(\frac{r}{q}K_{p})}$ and this implies
$(\partial_{\tau}-\mathcal{L})\{V-(K_{p}-e^{x})\}=qe^{x}-rK_{p}\leqslant qe^{x_{p}^{\rm ch}(\tau_{0})}-rK_{p}<0\quad\text{in}\ \ Q.$
Using the previous regularity result, letting $\tau\to\tau_{0}^{+}$ yields
$\partial_{\tau}V\leqslant qe^{x_{p}^{\rm ch}(\tau_{0})}-rK_{p}<0\quad\text{on }\ \mathcal{B}_{Q}.$
This contradicts the first estimate in Lemma 3.5.
The same argument applies to the other discontinuity case
$\lim_{\tau\to\tau_{0}^{-}}x_{p}^{\rm ch}(\tau)<x_{p}^{\rm ch}(\tau_{0}),$
which also leads to a contradiction. Therefore, $x_{p}^{\rm ch}$ is continuous on $(0,T)$. The continuity of $x_{c}^{\rm ch}$ follows analogously. ∎
In the following theorems, we will explore additional important properties of $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$.
Theorem 4.5.
Let $x_{p}^{\rm ch}(\tau)$ be defined as in (4.3). Then,
  • i)
    $x_{p}^{\rm ch}$ is strictly decreasing on $(0,T)$.
  • ii)
    ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}=\min\{\bar{x},{x_{p}}(0)\}$, where $\bar{x}$ is the point stated in (2.9).
Proof of i).
From Lemma 3.5, we can easily verify that $x_{p}^{\rm ch}(\tau)$ is monotone decreasing. Assume, for contradiction, that $x_{p}^{\rm ch}(\tau)$ is not strictly decreasing. Then there exists some $\tau_{1},\tau_{2}>0$ with $\tau_{1}<\tau_{2}$ such that $x_{p}^{\rm ch}(\tau_{1})=x_{p}^{\rm ch}(\tau_{2})=:x_{0}.$ Moreover, from that $x_{c}^{\rm ch}(0)\geqslant x_{c}(0)>\ln K_{c}$ (see subsection 2.4 and (4.5)), we see the cylinder $Q:=(\tau_{1},\tau_{2}-\delta)\times(x_{0},\ln{K_{c}})$ is contained in the continuation region $\mathcal{C}$. This implies $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$. For each sufficiently small $\delta>0$, we define by
$V_{\delta}(\tau,x):=V(\tau+\delta,x)-V(\tau,x)\quad\text{for all }\ (\tau,x)\in\Omega_{T-\delta}.$
Then, $V_{\delta}$ satisfies $\partial_{\tau}V_{\delta}-\mathcal{L}V_{\delta}=0$ in $Q$. Note that there exists a point $(\tau^{\prime},x^{\prime})\in Q$ such that $V_{\delta}(\tau^{\prime},x^{\prime})=0.$ Otherwise, we see from the parabolic Hopf lemma that
(4.6) $\partial_{x}V_{\delta}(\tau,x_{0})>0\quad\text{for \ all}\ \ \tau\in(\tau_{1},\tau_{2}-\delta).$
But $V(\tau,x_{0})=K_{p}-e^{x_{0}}$ and the continuity of $\partial_{x}V$, induced by the Sobolev embedding theorem, implies that $\partial_{x}V_{\delta}(\tau,x_{0})=0$ for all $\tau\in(\tau_{1},\tau_{2}-\delta)$, which contradicts (4.6).
Now, by the estimate in Lemma 3.5 and the continuity of $V$, it follows that $V_{\delta}(\tau,x)\geqslant 0$ for all $(\tau,x)\in Q$. Thus, $V_{\delta}$ attains its minimum as zero at $(\tau^{\prime},x^{\prime})\in Q$. Furthermore, since $\partial_{\tau}V_{\delta}-\mathcal{L}V_{\delta}=0$ in $Q$, we deduce from the Schauder theory that $V_{\delta}$ is smooth in $Q$. Applying the strong maximum principle [17, Theorem 2.7] to $V_{\delta}$ yields
(4.7) $V_{\delta}=0\quad\text{in }\ \{(\tau,x)\in Q:\tau\leqslant\tau^{\prime}\}.$
Define the cylinder $Q^{\prime}:=(0,\tau^{\prime})\times(\ln K_{p},\ln K_{c})$. In view of (4.3), (4.4), and the monotonicity of each free boundaries obtained by the estimates in Lemma 3.5, we find that $Q^{\prime}\subset\mathcal{C}\cap\Omega_{T-\delta}$, and this implies $\partial_{\tau}V_{\delta}-\mathcal{L}V_{\delta}=0$ in $Q^{\prime}$. Moreover, since (4.7) ensures that $V_{\delta}\equiv 0$ in the non-empty set $Q\cap Q^{\prime}$, applying the strong maximum principle again yields $V_{\delta}\equiv 0$ in $Q^{\prime}$. Since $\delta>0$ is arbitrary, we conclude that $V$ is constant with respect to $\tau$ in $Q^{\prime}$. Considering the initial condition $V(0,x)=\max\{C(0,x),P(0,x)\}$ and the regularity $V\in C(\overline{\Omega_{T}})$, it follows that
(4.8) $V(\tau,x)=\max\{C(0,x),P(0,x)\}\quad\text{for all }(\tau,x)\in Q^{\prime}.$
Since $C$ and $P$ are non-decreasing with respect to $\tau$ by Lemma 2.1 (iii), we have
$\max\{C(0,x),P(0,x)\}\leqslant\max\{C(\tau,x),P(\tau,x)\}\quad\text{for all }(\tau,x)\in Q^{\prime}.$
However, the inclusion $Q^{\prime}\subset\mathcal{C}$ yields
$V(\tau,x)>\max\{C(\tau,x),P(\tau,x)\}\geqslant\max\{C(0,x),P(0,x)\}\quad\text{for all }(\tau,x)\in Q^{\prime},$
and this contradicts (4.8) and completes the proof. ∎
Proof of ii).
We define $x_{p}^{\rm ch}(0)$ by
$x_{p}^{\rm ch}(0):=\sup\{x\in\mathbb{R}:V(0,x)=K_{p}-e^{x}\}.$
Since $V(0,x)=\max\{C(0,x),P(0,x)\}$ for all $x\in\mathbb{R}$, it is straightforward to verify that
$x_{p}^{\rm ch}(0)=\min\{\bar{x},x_{p}(0)\}.$
We first observe that $\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)$ exists, since $x_{p}^{\rm ch}(\tau)$ is monotonically decreasing and bounded from above.
Suppose ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}>x_{p}^{\rm ch}(0)$. Then there exists some $\tau_{p}>0$ such that $x_{p}^{\rm ch}(\tau_{p})>x_{p}^{\rm ch}(0)$. Hence,
$K_{p}-e^{x_{p}^{\rm ch}(\tau_{p})}=V\big(\tau_{p},x_{p}^{\rm ch}(\tau_{p})\big)\geqslant J\big(\tau_{p},x_{p}^{\rm ch}(\tau_{p})\big)\geqslant J\big(0,x_{p}^{\rm ch}(\tau_{p})\big)>K_{p}-e^{x_{p}^{\rm ch}(\tau_{p})},$
where $J=\max\{C,P\}$ and the second inequality follows from the result $\partial_{\tau}C\geqslant 0$ and $\partial_{\tau}P\geqslant 0$ in $\Omega_{T}$ (see Lemma 2.1 iii)). Therefore, a contradiction occurs.
On the other hand, if ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}$ is strictly less than $x_{p}^{\rm ch}(0)$, there exist some $\varepsilon>0$ and $\delta>0$ such that for all $\tau\in(0,\delta)$,
$x^{\rm ch}_{p}(0)-x^{\rm ch}_{p}(\tau)\geqslant\varepsilon.$
Now, consider a cylinder $Q:=(0,\delta)\times(x^{\rm ch}_{p}(0)-\frac{\varepsilon}{2},x^{\rm ch}_{p}(0))$. By construction, we see that $Q$ is contained in the continuation region $\mathcal{C}$ and thus, $\partial_{\tau}V-\mathcal{L}V=0$ in $Q$. Moreover, since $x^{\rm ch}_{p}(0)<\ln({\frac{r}{q}}K_{p})$, it follows that
$(\partial_{\tau}-\mathcal{L})(V-(K_{p}-e^{x}))\leqslant qe^{x^{\rm ch}_{p}(0)}-rK_{p}<0\quad\text{in}\ \ Q.$
Note that since $V=K_{p}-e^{x}$ on the line segment $\mathcal{B}_{Q}:=\{0\}\times(x^{\rm ch}_{p}(0)-\frac{\varepsilon}{2},x^{\rm ch}_{p}(0))$, we apply the result in [8, Theorem 19, p. 321] to $V-(K_{p}-e^{x})$ to deduce that $V$ is smooth in $Q$ and up to the initial boundary $\mathcal{B}_{Q}$. Thus, letting $\tau\to 0^{+}$ in the above inequality yields $\partial_{\tau}V<0$ on $\mathcal{B}_{Q}$. This contradicts the first estimate in Lemma 3.5. Therefore, we conclude ${\lim_{\tau\to 0^{+}}x_{p}^{\rm ch}(\tau)}=x_{p}^{\rm ch}(0)$. ∎
By applying the same argument, analogous results for $x_{c}^{\rm ch}(\tau)$ can be obtained. Hence we omit the proof.
Theorem 4.6.
Let $x_{c}^{\rm ch}(\tau)$ be defined as in (4.4). Then,
  • i)
    $x_{c}^{\rm ch}$ is strictly increasing on $(0,T)$.
  • ii)
    ${\lim_{\tau\to 0^{+}}x_{c}^{\rm ch}(\tau)}=\max\{\bar{x},{x_{c}}(0)\}$, where $\bar{x}$ is the point stated in (2.9).
Figure 3. The free boundaries $x_{c}^{\rm ch}(\tau)$ and ${x}_{p}^{\rm ch}(\tau)$.
Figure 3 illustrates the behavior of the two free boundaries, $x_{c}^{\rm ch}(\tau)$ and $x_{p}^{\rm ch}(\tau)$, for $V(\tau,x)$. In this figure, we can observe that, as shown in Theorems 4.5 and 4.6, $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$ are strictly decreasing and strictly increasing in $\tau\in[0,T]$, respectively. Moreover, we can also verify the inequality (4.5), which was discussed earlier.
4.3. Regularity of the free boundary
We conclude this section by proving that the free boundaries $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$ are smooth. To reach such a conclusion, we first establish the $C^{1}$-regularity of the free boundaries, which allows us to reformulate the obstacle problem as a Stefan-type free boundary problem (see [28]). The proof of the $C^{1}$-regularity relies fundamentally on [9, Theorem 4.6].
In our framework, recall that the continuation region $\mathcal{C}$ can be represented as
(4.9) $\mathcal{C}=\{(\tau,x)\in\Omega_{T}:x_{p}^{\rm ch}(\tau)<x<x_{c}^{\rm ch}(\tau)\},$
and we have verified that the free boundaries $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$ are monotonically decreasing and increasing curves, respectively. In [9], conditions (3.1) and (3.2) are introduced to guarantee the monotonicity of the free boundaries and the configuration of the continuation region described as in (4.9). Since our framework already exhibits these properties, we can directly apply the arguments from the proof of [9, Theorem 4.6] to our problem (2.7) to deduce the following proposition:
Proposition 4.7.
Let $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$ be defined as in (4.3) and (4.4). Then, $x_{p}^{\rm ch},x_{c}^{\rm ch}\in C^{1}((0,T))$. Moreover, $\partial_{\tau}V$ and $\partial_{\tau x}V$ are continuous up to the free boundaries $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$.
Once the $C^{1}$-regularity of the free boundary $x_{p}^{\rm ch}(\tau)$ is established, we reformulate the obstacle problem as a Stefan-type problem to obtain higher regularity for $x_{p}^{\rm ch}(\tau)$. Specifically, since $V$ contacts with $K_{p}-e^{x}$ on the free boundary $x_{p}^{\rm ch}(\tau)$, we denote $W:=V-(K_{p}-e^{x})$ and $Z:=\partial_{\tau}W$ and restrict our analysis to the region
$\mathcal{C}_{p}:=\{(\tau,x)\in\Omega_{T}:x_{p}^{\rm ch}(\tau)<x<\ln K_{c}\}.$
By Proposition 4.7, the derivatives $\partial_{\tau}W$ and $\partial_{\tau x}W$ are continuous up to the free boundary $x_{p}^{\rm ch}(\tau)$. Using the conditions $W(\tau,x_{p}^{\rm ch}(\tau))=0$ and $\partial_{x}W(\tau,x_{p}^{\rm ch}(\tau))=0$, we differentiate the former with respect to $\tau$ to obtain
$\partial_{\tau}W(\tau,x_{p}^{\rm ch}(\tau))+\partial_{x}W(\tau,x_{p}^{\rm ch}(\tau))\left[\frac{d}{d\tau}x_{p}^{\rm ch}(\tau)\right]=0.$
Since $\partial_{x}W(\tau,x_{p}^{\rm ch}(\tau))=0$, this yields $\partial_{\tau}W(\tau,x_{p}^{\rm ch}(\tau))=0$. Next, evaluating the governing equation $\partial_{\tau}W-\mathcal{L}W=qe^{x}-rK_{p}$ and utilizing the fact that $W=\partial_{x}W=\partial_{\tau}W=0$ at the free boundary, we find that $\partial_{xx}W$ is also continuous up to the boundary and satisfies
$\frac{\sigma^{2}}{2}\partial_{xx}W(\tau,x_{p}^{\rm ch}(\tau))=qe^{x_{p}^{\rm ch}(\tau)}-rK_{p}.$
Furthermore, differentiating the second condition $\partial_{x}W(\tau,x_{p}^{\rm ch}(\tau))=0$ with respect to $\tau$ gives
$\partial_{\tau x}W(\tau,x_{p}^{\rm ch}(\tau))+\partial_{xx}W(\tau,x_{p}^{\rm ch}(\tau))\left[\frac{d}{d\tau}x_{p}^{\rm ch}(\tau)\right]=0.$
Substituting the expression for $\partial_{xx}W$ into this equation, and recalling that $Z=\partial_{\tau}W$ (which implies $\partial_{x}Z=\partial_{\tau x}W$), we deduce the following Stefan-type condition:
(4.10) $\frac{d}{d\tau}x_{p}^{\rm ch}(\tau)=\frac{\sigma^{2}}{2}\frac{\partial_{\tau x}W(\tau,x_{p}^{\rm ch}(\tau))}{qe^{x_{p}^{\rm ch}(\tau)}-rK_{p}}=\frac{\sigma^{2}}{2}\frac{\partial_{x}Z(\tau,x_{p}^{\rm ch}(\tau))}{qe^{x_{p}^{\rm ch}(\tau)}-rK_{p}},\quad\tau\in[0,T].$
The denominator in (4.10) is strictly negative (and thus non-zero), since $x_{p}^{\rm ch}(\tau)\leqslant x_{p}(0)=\hat{x}_{p}(T_{p}-T)<\ln\big(\frac{r}{q}K_{p}\big)$ for all $\tau\in[0,T]$.
For any sufficiently small $\delta\in(0,1)$, the function $Z$ satisfies
$\begin{cases}\partial_{\tau}Z-\mathcal{L}Z=0\quad&\text{in }\{(\tau,x)\in\mathcal{C}_{p}:\delta<\tau<T\},\\ Z(\tau,x_{p}^{\rm ch}(\tau))=0\quad\text{and}\quad Z(\tau,\ln K_{c})=\partial_{\tau}V(\tau,\ln K_{c})\quad&\text{for }\tau\in(\delta,T),\\ Z(\delta,x)=\partial_{\tau}V(\delta,x)\quad&\text{for }x_{p}^{\rm ch}(\delta)<x<\ln K_{c}.\end{cases}$
Here, we need to consider the positive constant $\delta$ since $V$ is not sufficiently regular at the point $(0,\bar{x})$. Then, applying the standard bootstrap argument in [23], to the above equation with considering (4.10), we conclude that the free boundary $x_{p}^{\rm ch}(\tau)$ is $C^{\infty}$ on $(2\delta,T)$ for every $\delta>0$ and hence on $(0,T)$. The same argument implies that the free boundary $x_{c}^{\rm ch}(\tau)$ is also $C^{\infty}$ on $(0,T)$. Therefore, we have obtained the following theorem.
Theorem 4.8.
Let $x_{p}^{\rm ch}(\tau)$ and $x_{c}^{\rm ch}(\tau)$ be defined as in (4.3) and (4.4). Then, $x_{p}^{\rm ch},\,x_{c}^{\rm ch}\in C^{\infty}((0,T))$.
Remark 4.9.
Lemma 4.1 is a key observation for the analysis of the free boundaries. It shows that the solution $V$ coincides with the smooth payoff functions $e^{x}-K_{c}$ and $K_{p}-e^{x}$ on the exercise region $\mathcal{E}$. Consequently, near the free boundaries the obstacle behaves like a smooth function, and the Stefan-type condition (4.10) is well-defined.
In particular, the denominator in (4.10), which involves the difference between the spatial derivatives of $V$ and the obstacle, is non-vanishing. This allows us to apply the standard regularity arguments for Stefan-type free boundary problems.
If the contact occurred with the continuation region of $C$ or $P$ instead of their payoff parts $e^{x}-K_{c}$ and $K_{p}-e^{x}$, the obstacle would no longer be smooth and the denominator in (4.10) might vanish. In that case, the above argument would fail and the smoothness of the free boundaries would not follow directly.
Appendix
In the Appendix, we derive the formulation of the problem (1.1) from (2.4) and verify that the solution of (1.1) satisfies (2.4). Furthermore, we provide the proof of Lemma 2.1, and show that $\mathcal{F}$ in the proof of Theorem 3.1 satisfies the conditions of the Schauder fixed point theorem.
A. Formulation of the model
Recall that $V^{\rm ch}$ satisfies that for each $t\in(0,T)$,
$V^{\rm ch}(t,S_{t})=\sup_{\theta\in\mathcal{U}_{t,T}}\mathbb{E}\left[e^{-r(\theta-t)}\max\left\{C^{A}(\theta,S_{\theta}),P^{A}(\theta,S_{\theta})\right\}\mid\mathcal{G}_{t}\right],$
where $\mathcal{U}_{t,T}$ is the set of all stopping times taking values in $[t,T]$ and $S_{t}$ is the stock price that follows a geometric Brownian motion (GBM) given by
$dS_{t}=(r-q)S_{t}dt+\sigma S_{t}dW_{t}^{\mathbb{Q}},\quad S_{0}>0.$
For each $t\in(0,T)$, since $S_{t}$ is $\mathcal{G}_{t}$-measurable and both $s\to C^{A}(t,s)$ and $s\to P^{A}(t,s)$ are Borel measurable, it follows that
$V^{\rm ch}(t,S_{t})$ $=\sup_{\theta\in\mathcal{U}_{t,T}}\mathbb{E}\left[e^{-r(\theta-t)}\max\left\{C^{A}(\theta,S_{\theta}),P^{A}(\theta,S_{\theta})\right\}\mid\mathcal{G}_{t}\right]$
$\geqslant\mathbb{E}\left[\max\left\{C^{A}(t,S_{t}),P^{A}(t,S_{t})\right\}\mid\mathcal{G}_{t}\right]$
$=\max\left\{C^{A}(t,S_{t}),P^{A}(t,S_{t})\right\}$
Similarly, since $\mathcal{U}_{T,T}=\{T\}$,
$V^{\rm ch}(T,S_{T})=\max\{C^{A}(T,S_{T}),P^{A}(T,S_{T})\}$
and this implies $V^{\rm ch}(T,s)=\max\{C^{A}(T,s),P^{A}(T,s)\}$ for all $s\in(0,\infty)$. For each $t\in(0,T)$ and small $h>0$, since $\mathcal{G}_{t}\subset\mathcal{G}_{t+h}$ we deduce by the tower property for the conditional expectation that
$e^{-rt}V^{\rm ch}(t,S_{t})$ $=\sup_{\theta\in\mathcal{U}_{t,T}}\mathbb{E}\left[e^{-r\theta}\max\left\{C^{A}(\theta,S_{\theta}),P^{A}(\theta,S_{\theta})\right\}\mid\mathcal{G}_{t}\right]$
$=\sup_{\theta\in\mathcal{U}_{t,T}}\mathbb{E}\left[\mathbb{E}\left[e^{-r\theta}\max\left\{C^{A}(\theta,S_{\theta}),P^{A}(\theta,S_{\theta})\right\}\mid\mathcal{G}_{t+h}\right]\mid\mathcal{G}_{t}\right]$
$\geqslant\mathbb{E}\left[\sup_{\theta\in\mathcal{U}_{t+h,T}}\mathbb{E}\left[e^{-r\theta}\max\left\{C^{A}(\theta,S_{\theta}),P^{A}(\theta,S_{\theta})\right\}\mid\mathcal{G}_{t+h}\right]\mid\mathcal{G}_{t}\right]$
$=\mathbb{E}\left[e^{-r(t+h)}V^{\rm ch}(t+h,S_{t+h})\mid\mathcal{G}_{t}\right].$
This yields $Z_{t}:=e^{-rt}V^{\rm ch}(t,S_{t})$ is a supermartingale under $\mathbb{P}$.
Applying Itô’s formula onto $Z_{t}$, we obtain
$Z_{t+h}-Z_{t}$ $=\int_{t}^{t+h}e^{-ru}\Big\{\partial_{t}V^{\rm ch}(u,S_{u})+\frac{\sigma^{2}}{2}S_{u}^{2}\partial_{ss}V^{\rm ch}(u,S_{u})$
$\quad+(r-q)S_{u}\partial_{s}V^{\rm ch}(u,S_{u})-rV^{\rm ch}(u,S_{u})\Big\}du+\int_{t}^{t+h}e^{-ru}\sigma S_{u}\partial_{s}V^{\rm ch}(u,S_{u})dW_{u}.$
Moreover, using that $Z_{t}$ is $\mathcal{G}_{t}$-measurable and a supermartingale,
$\mathbb{E}\Big[\int_{t}^{t+h}e^{-ru}\big\{\partial_{t}V^{\rm ch}(u,S_{u})+\frac{\sigma^{2}}{2}S_{u}^{2}\partial_{ss}V^{\rm ch}(u,S_{u})$
$\quad+(r-q)S_{u}\partial_{s}V^{\rm ch}(u,S_{u})-rV^{\rm ch}(u,S_{u})\big\}du\mid\mathcal{G}_{t}\Big]=\mathbb{E}\left[Z_{t+h}-Z_{t}\mid\mathcal{G}_{t}\right]\leqslant 0.$
Taking $h\to 0^{+}$, the Lebesgue differentiation theorem and the dominated convergence theorem yields
$\partial_{t}V^{\rm ch}(t,S_{t})+\frac{\sigma^{2}}{2}S_{t}^{2}\partial_{ss}V^{\rm ch}(t,S_{t})+(r-q)S_{t}\partial_{s}V^{\rm ch}(t,S_{t})-rV^{\rm ch}(t,S_{t})\leqslant 0\quad\text{a.s. in }\ \Omega.$
Since $S_{t}$ has a strictly positive density on $(0,\infty)$, we conclude
$\partial_{t}V^{\rm ch}(t,s)+\mathfrak{L}V^{\rm ch}(t,s)\leqslant 0\quad\text{for a.e. }\ (t,s)\in(0,T)\times(0,\infty).$
If $V^{\rm ch}(t,S_{t})>\max\{C^{A}(t,S_{t}),P^{A}(t,S_{t})\}$, there exists some small $h>0$ and a stopping time $\theta_{t}$ such that $\theta_{t}\in\mathcal{U}_{t+h,T}$ and
$V^{\rm ch}(t,S_{t})=\mathbb{E}\left[e^{-r(\theta_{t}-t)}\max\left\{C^{A}(\theta_{t},S_{\theta_{t}}),P^{A}(\theta_{t},S_{\theta_{t}})\right\}\mid\mathcal{G}_{t}\right].$
Then, $V^{\rm ch}(t,S_{t})$ satisfies
$e^{-rt}V^{\rm ch}(t,S_{t})$ $=\mathbb{E}\left[e^{-r\theta_{t}}\max\left\{C^{A}(\theta_{t},S_{\theta_{t}}),P^{A}(\theta_{t},S_{\theta_{t}})\right\}\mid\mathcal{G}_{t}\right]$
$=\mathbb{E}\left[\mathbb{E}\left[e^{-r\theta_{t}}\max\left\{C^{A}(\theta_{t},S_{\theta_{t}}),P^{A}(\theta_{t},S_{\theta_{t}})\right\}\mid\mathcal{G}_{t+h}\right]\mid\mathcal{G}_{t}\right]$
$=\mathbb{E}\left[\sup_{\theta\in\mathcal{U}_{t+h,T}}\mathbb{E}\left[e^{-r\theta}\max\left\{C^{A}(\theta,S_{\theta}),P^{A}(\theta,S_{\theta})\right\}\mid\mathcal{G}_{t+h}\right]\mid\mathcal{G}_{t}\right]$
$=\mathbb{E}\left[e^{-r(t+h)}V^{\rm ch}(t+h,S_{t+h})\mid\mathcal{G}_{t}\right].$
This implies that $e^{-r(t\wedge\theta_{t})}V^{\rm ch}(t\wedge\theta_{t},S_{t\wedge\theta_{t}})$ is a martingale under $\mathbb{P}$. Therefore, if $V^{\rm ch}(t,S_{t})>\max\{C^{A}(t,S_{t}),P^{A}(t,S_{t})\}$, by applying the Itô’s formula, we can similarly obtain
$\partial_{t}V^{\rm ch}(t,S_{t})+\frac{\sigma^{2}}{2}S_{t}^{2}\partial_{ss}V^{\rm ch}(t,S_{t})+(r-q)S_{t}\partial_{s}V^{\rm ch}(t,S_{t})-rV^{\rm ch}(t,S_{t})=0.$
Therefore, we deduce
$\partial_{t}V^{\rm ch}(t,s)+\mathfrak{L}V^{\rm ch}(t,s)=0$
for almost every $(t,s)\in(0,T)\times(0,\infty)$ satisfying $V^{\rm ch}(t,s)>\max\{C^{A}(t,s),P^{A}(t,s)\}$.
Combining all, our desired formulation for $V^{\rm ch}$ would be as follows:
[…]
B. Verification
Since $V^{\rm ch}\in W^{1,2}_{p,{\rm loc}}((0,T)\times(0,\infty))$, we apply the Itô’s lemma for the Sobolev space (see [14, Theorem 1, p.122]) to see that for any $\theta\in\mathcal{U}_{t,T}$,
$e^{-r(\theta-t)}$ $V^{\rm ch}(\theta,S_{\theta})$
$=$ $\ V^{\rm ch}(t,S_{t})+\int_{t}^{\theta}e^{-r(u-t)}\Big(\partial_{t}V^{\rm ch}(u,S_{u})+\frac{\sigma^{2}}{2}S^{2}_{u}\partial_{ss}V^{\rm ch}(u,S_{u})$
$+(r-q)S_{u}\partial_{s}V^{\rm ch}(u,S_{u})-rV^{\rm ch}(u,S_{u})\Big)du+\int_{t}^{\theta}\sigma S_{u}\partial_{s}V^{\rm ch}(u,S_{u})dW_{u}$
$\leqslant$ $\ V^{\rm ch}(t,S_{t})+\int_{t}^{\theta}\sigma S_{u}\partial_{s}V^{\rm ch}(u,S_{u})dW_{u}$
since $\partial_{t}V^{\rm ch}(t,s)+\mathfrak{L}V^{\rm ch}(t,s)\leqslant 0$ for almost every $(t,s)\in(0,T)\times(0,\infty)$. Note that the second estimate in Lemma 3.5 implies that $1\leqslant\partial_{s}V^{\rm ch}(t,s)\leqslant 1$ for all $(t,s)\in(0,T)\times(0,\infty)$. This yields that
(4.12) $\mathbb{E}\left[\int_{t}^{\theta}\Big(e^{-r(u-t)}\sigma S_{u}\partial_{s}V^{\rm ch}(u,S_{u})\Big)^{2}du\right]\leqslant\mathbb{E}\left[\int_{t}^{\theta}\Big(e^{-r(u-t)}\sigma S_{u}\Big)^{2}du\right]<\infty.$
Denote by
$M_{v}:=\int_{t}^{v}\sigma S_{u}\partial_{s}V^{\rm ch}(u,S_{u})dW_{u},\quad v\in[t,T].$
Then, by the result (4.12), we obtain that $(M_{v})_{v\in[t,T]}$ is a $\mathcal{G}_{v}$-martingale. Thus, taking conditional expectation $\mathbb{E}[\cdot\mid\mathcal{G}_{t}]$ on both sides, we get
$V^{\rm ch}(t,S_{t})$ $\geqslant\mathbb{E}\left[e^{-r(\theta-t)}V^{\rm ch}(\theta,S_{\theta})\mid\mathcal{G}_{t}\right]$
$\geqslant\mathbb{E}\left[e^{-r(\theta-t)}\max\{C^{A}(\theta,S_{\theta}),P^{A}(\theta,S_{\theta})\}\mid\mathcal{G}_{t}\right]$
where the last inequality follows from the fact that $V^{\rm ch}(t,S_{t})\geqslant\max\{C^{A}(t,S_{t}),P^{A}(t,S_{t})\}$ holds for all $t\in(0,T)$. Since $\theta\in\mathcal{U}_{t,T}$ is arbitrary, we deduce
$V^{\rm ch}(t,S_{t})\geqslant\sup_{\theta\in\mathcal{U}_{t,T}}\mathbb{E}\left[e^{-r(\theta-t)}\max\{C^{A}(\theta,S_{\theta}),P^{A}(\theta,S_{\theta})\}\mid\mathcal{G}_{t}\right].$
Define $\theta^{*}$ as
$\theta^{*}:=\inf\{u\in[t,T]:V^{\rm ch}(u,S_{u})=\max\{C^{A}(u,S_{u}),P^{A}(u,S_{u})\}\}\wedge T.$
Then, if $(t,S_{t})$ lies in the region $\{V^{\rm ch}=\max\{C^{A},P^{A}\}\}$ we have $\theta^{*}=t$ and thus,
$V^{\rm ch}(t,S_{t})=\max\{C^{A}(t,S_{t}),P^{A}(t,S_{t})\}=\mathbb{E}\left[e^{-r(\theta^{*}-t)}\max\{C^{A}(\theta^{*},S_{\theta^{*}}),P^{A}(\theta^{*},S_{\theta^{*}})\}\mid\mathcal{G}_{t}\right].$
If $(t,S_{t})$ lies in the region $\{V^{\rm ch}>\max\{C^{A},P^{A}\}\}$, we have
(4.13) $\begin{cases}\partial_{t}V^{\rm ch}(u,S_{u})+\frac{\sigma^{2}}{2}S^{2}_{u}\partial_{ss}V^{\rm ch}(u,S_{u})+(r-q)S_{u}\partial_{s}V^{\rm ch}(u,S_{u})-rV^{\rm ch}(u,S_{u})=0,\\ \qquad\text{for all }\ u\in(t,\theta^{*}),\\ V^{\rm ch}(\theta^{*},S_{\theta^{*}})=\max\{C^{A}(\theta^{*},S_{\theta^{*}}),P^{A}(\theta^{*},S_{\theta^{*}})\}.\end{cases}$
By (4.11) and (4.13), we obtain that
$V^{\rm ch}(t,S_{t})$ $=\mathbb{E}\left[e^{-r(\theta^{*}-t)}V^{\rm ch}(\theta^{*},S_{\theta^{*}})\mid\mathcal{G}_{t}\right]$
$=\mathbb{E}\left[e^{-r(\theta^{*}-t)}\max\{C^{A}(\theta^{*},S_{\theta^{*}}),P^{A}(\theta^{*},S_{\theta^{*}})\}\mid\mathcal{G}_{t}\right].$
Combining all, we conclude
$V^{\rm ch}(t,S_{t})=\sup_{\theta\in\mathcal{U}_{t,T}}\mathbb{E}\left[e^{-r(\theta-t)}\max\{C^{A}(\theta,S_{\theta}),P^{A}(\theta,S_{\theta})\}\mid\mathcal{G}_{t}\right].$
C. Proof of Lemma 2.1
Similar results and approaches used in the proof can be found in previous research; see, e.g., [26, 27]. Therefore, we shall omit the details of approximations.
Proof of Part (i)
Let $\tilde{C}^{A}_{n,\varepsilon}$ and $\tilde{P}^{A}_{n,\varepsilon}$ be a solutions of the following equations with Dirichlet boundary conditions:
(4.14) $\begin{cases}\partial_{\zeta}\tilde{C}^{A}_{n,\varepsilon}-\mathcal{L}\tilde{C}^{A}_{n,\varepsilon}+\beta_{c,\varepsilon}(\tilde{C}^{A}_{n,\varepsilon}-\varphi_{\varepsilon}(e^{x}-K_{c}))=0\quad\text{in }\ {\Omega}^{n}_{T_{c}},\\ \tilde{C}^{A}_{n,\varepsilon}(\zeta,-n)=0\quad\text{and}\quad\tilde{C}^{A}_{n,\varepsilon}(\zeta,n)=e^{n}-K_{c}\quad\text{for }\ \zeta\in[0,T_{c}],\\ \tilde{C}^{A}_{n,\varepsilon}(0,x)=\varphi_{\varepsilon}(e^{x}-K_{c})\quad\text{for }\ x\in(-n,n)\end{cases}$
and
(4.15) $\begin{cases}\partial_{\zeta}\tilde{P}^{A}_{n,\varepsilon}-\mathcal{L}\tilde{P}^{A}_{n,\varepsilon}+\beta_{p,\varepsilon}(\tilde{P}^{A}_{n,\varepsilon}-\varphi_{\varepsilon}(K_{p}-e^{x}))=0\quad\text{in }\ {\Omega}^{n}_{T_{p}},\\ \tilde{P}^{A}_{n,\varepsilon}(\zeta,-n)=K_{p}-e^{-n}\quad\text{and}\quad\tilde{P}^{A}_{n,\varepsilon}(\zeta,n)=0\quad\text{for }\ \zeta\in[0,T_{p}],\\ \tilde{P}^{A}_{n,\varepsilon}(0,x)=\varphi_{\varepsilon}(K_{p}-e^{x})\quad\text{for }\ x\in(-n,n).\end{cases}$
where $\beta_{c,\varepsilon}$ and $\beta_{p,\varepsilon}$ are appropriate penalty functions, ${\Omega}^{n}_{T_{c}}:=(0,T_{c})\times(-n,n)$ and ${\Omega}^{n}_{T_{p}}:=(0,T_{p})\times(-n,n)$. It can be seen that the limits $\tilde{C}^{A}_{n}:=\displaystyle\lim_{\varepsilon\to 0^{+}}\tilde{C}^{A}_{n,\varepsilon}$ and $\tilde{P}^{A}_{n}:=\displaystyle\lim_{\varepsilon\to 0^{+}}\tilde{P}^{A}_{n,\varepsilon}$ are well-defined and satisfy the following obstacle problems, respectively:
$\begin{cases}\partial_{\zeta}\tilde{C}^{A}_{n}(\zeta,x)-\mathcal{L}\tilde{C}^{A}_{n}(\zeta,x)\geqslant 0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{c}}\ \ \text{with }\ \tilde{C}^{A}_{n}(\zeta,x)=(e^{x}-K_{c})^{+},\\ \partial_{\zeta}\tilde{C}^{A}_{n}(\zeta,x)-\mathcal{L}\tilde{C}^{A}_{n}(\zeta,x)=0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{c}}\ \ \text{with }\ \tilde{C}^{A}_{n}(\zeta,x)>(e^{x}-K_{c})^{+},\\ \tilde{C}^{A}_{n}(\zeta,-n)=0\quad\text{and}\quad\tilde{C}^{A}_{n}(\zeta,n)=e^{n}-K_{c}\quad\text{for }\ \zeta\in[0,T_{c}],\\ \tilde{C}^{A}_{n}(0,x)=(e^{x}-K_{c})^{+}\quad\text{for }\ x\in(-n,n)\end{cases}$
and
$\begin{cases}\partial_{\zeta}\tilde{P}^{A}_{n}(\zeta,x)-\mathcal{L}\tilde{P}^{A}_{n}(\zeta,x)\geqslant 0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{p}}\ \ \text{with }\ \tilde{P}^{A}_{n}(\zeta,x)=(K_{p}-e^{x})^{+},\\ \partial_{\zeta}\tilde{P}^{A}_{n}(\zeta,x)-\mathcal{L}\tilde{P}^{A}_{n}(\zeta,x)=0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{p}}\ \ \text{with }\ \tilde{P}^{A}_{n}(\zeta,x)>(K_{p}-e^{x})^{+},\\ \tilde{P}^{A}_{n}(\zeta,-n)=K_{p}-e^{-n}\quad\text{and}\quad\tilde{P}^{A}_{n,\varepsilon}(\zeta,n)=0\quad\text{for }\ \zeta\in[0,T_{p}],\\ \tilde{P}^{A}_{n}(0,x)=(K_{p}-e^{x})^{+}\quad\text{for }\ x\in(-n,n).\end{cases}$
Observe that $e^{x}+2$ satisfies
$\begin{cases}\partial_{\zeta}(e^{x}+2)-\mathcal{L}(e^{x}+2)+\beta_{c,\varepsilon}(e^{x}+2-\varphi_{\varepsilon}(e^{x}-K_{c}))>0\quad\text{for }\ (\zeta,x)\in\Omega^{n}_{T_{c}},\\ e^{-n}+2>0=\tilde{C}^{A}_{n,\varepsilon}(\zeta,-n)\quad\text{and}\quad e^{n}+2>e^{n}-K_{c}=\tilde{C}^{A}_{n,\varepsilon}(\zeta,n)\quad\text{for }\ \zeta\in[0,T_{c}],\\ e^{x}+2>\varphi_{\varepsilon}(e^{x}-K_{c})=\tilde{C}^{A}_{n,\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n)\end{cases}$
since $\beta_{c,\varepsilon}\in C^{\infty}(\mathbb{R})$ is a penalty function satisfying $\beta_{c,\varepsilon}(\lambda)=0$ for all $\lambda\geqslant\varepsilon$ and $e^{x}+2>\varphi_{\varepsilon}(e^{x}-K_{c})+\varepsilon$ holds for all $\varepsilon<1.$ Therefore, for each $0<\varepsilon<1$, we deduce by the comparison principle [8, 52p] that $\tilde{C}^{A}_{n,\varepsilon}\leqslant e^{x}+2$ in $\Omega^{n}_{T_{c}}$.
Similarly, $K_{p}+2$ satisfies
$\begin{cases}\partial_{\zeta}(K_{p}+2)-\mathcal{L}(K_{p}+2)+\beta_{p,\varepsilon}(K_{p}+2-\varphi_{\varepsilon}(K_{p}-e^{x}))>0\quad\text{for }\ (\zeta,x)\in\Omega^{n}_{T_{p}},\\ K_{p}+2>K_{p}-e^{-n}=\tilde{P}^{A}_{n,\varepsilon}(\zeta,-n)\quad\text{and}\quad K_{p}+2>0=\tilde{P}^{A}_{n,\varepsilon}(\zeta,n)\quad\text{for }\ \zeta\in[0,T_{p}],\\ K_{p}+2>\varphi_{\varepsilon}(K_{p}-e^{x})=\tilde{P}^{A}_{n,\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
By the comparison principle [8, 52p], $\tilde{P}^{A}_{n,\varepsilon}\leqslant K_{p}+2$ in $\Omega^{n}_{T_{p}}$ holds for all $0<\varepsilon<1$. Take $\varepsilon\to 0^{+}$ to get
$\tilde{C}^{A}_{n}\leqslant e^{x}+2\quad\text{in }\ \Omega_{T_{c}}^{n}\quad\text{and}\quad\tilde{P}^{A}_{n}\leqslant K_{p}+2\quad\text{in }\ \Omega_{T_{p}}^{n}.$
It can be seen further that ${\lim_{n\to\infty}}\tilde{C}^{A}_{n}$ and ${\lim_{n\to\infty}}\tilde{P}^{A}_{n}$ satisfy (2.5) and (2.6) on each compact subset of $\Omega_{T_{c}}$ and $\Omega_{T_{p}}$, respectively. Since the solution of each problem (2.5) and (2.6) is unique, we see that $C^{A}=\displaystyle{\lim_{n\to\infty}}\tilde{C}^{A}_{n}$ in $\Omega_{T_{c}}$ and $P^{A}=\displaystyle{\lim_{n\to\infty}}\tilde{P}^{A}_{n}$ in $\Omega_{T_{p}}$. Therefore, taking $n\rightarrow\infty$ yields the results in Part (i).
Proof of Part (ii)
Let $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ be the unique solutions to the following equations with Neumann boundary conditions:
(4.16) $\begin{cases}\partial_{\zeta}\widehat{C}^{A}_{n,\varepsilon}-\mathcal{L}\widehat{C}^{A}_{n,\varepsilon}+\beta_{c,\varepsilon}(\widehat{C}^{A}_{n,\varepsilon}-\varphi_{\varepsilon}(e^{x}-K_{c}))=0\quad\text{in }\ {\Omega}^{n}_{T_{c}},\\ \partial_{x}\widehat{C}^{A}_{n,\varepsilon}(\zeta,-n)=0\quad\text{and}\quad\partial_{x}\widehat{C}^{A}_{n,\varepsilon}(\zeta,n)=e^{n}\quad\text{for }\ \zeta\in[0,T_{c}],\\ \widehat{C}^{A}_{n,\varepsilon}(0,x)=\varphi_{\varepsilon}(e^{x}-K_{c})\quad\text{for }\ x\in(-n,n)\end{cases}$
and
(4.17) $\begin{cases}\partial_{\zeta}\widehat{P}_{n,\varepsilon}^{A}-\mathcal{L}\widehat{P}_{n,\varepsilon}^{A}+\beta_{p,\varepsilon}(\widehat{P}^{A}_{n,\varepsilon}-\varphi_{\varepsilon}(K_{p}-e^{x}))=0\quad\text{in }\ {\Omega}^{n}_{T_{p}},\\ \partial_{x}\widehat{P}_{n,\varepsilon}^{A}(\zeta,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}\widehat{P}_{n,\varepsilon}^{A}(\zeta,n)=0\quad\text{for }\ \zeta\in[0,T_{p}],\\ \widehat{P}_{n,\varepsilon}^{A}(0,x)=\varphi_{\varepsilon}(K_{p}-e^{x})\quad\text{for }\ x\in(-n,n).\end{cases}$
Then $\widehat{C}^{A}_{n}:=\displaystyle\lim_{\varepsilon\to 0^{+}}\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n}:=\displaystyle\lim_{\varepsilon\to 0^{+}}\widehat{P}^{A}_{n,\varepsilon}$ are well-defined and solutions to the following obstacle problems:
$\begin{cases}\partial_{\zeta}\widehat{C}^{A}_{n}(\zeta,x)-\mathcal{L}\widehat{C}^{A}_{n}(\zeta,x)\geqslant 0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{c}}\ \ \text{with }\ \widehat{C}^{A}_{n}(\zeta,x)=(e^{x}-K_{c})^{+},\\ \partial_{\zeta}\widehat{C}^{A}_{n}(\zeta,x)-\mathcal{L}\widehat{C}^{A}_{n}(\zeta,x)=0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{c}}\ \ \text{with }\ \widehat{C}^{A}_{n}(\zeta,x)>(e^{x}-K_{c})^{+},\\ \partial_{x}\widehat{C}^{A}_{n}(\zeta,-n)=0\quad\text{and}\quad\partial_{x}\widehat{C}^{A}_{n}(\zeta,n)=e^{n}\quad\text{for }\ \zeta\in[0,T_{c}],\\ \widehat{C}^{A}_{n}(0,x)=(e^{x}-K_{c})^{+}\quad\text{for }\ x\in(-n,n)\end{cases}$
and
$\begin{cases}\partial_{\zeta}\widehat{P}^{A}_{n}(\zeta,x)-\mathcal{L}\widehat{P}^{A}_{n}(\zeta,x)\geqslant 0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{p}}\ \ \text{with }\ \widehat{P}^{A}_{n}(\zeta,x)=(K_{p}-e^{x})^{+},\\ \partial_{\zeta}\widehat{P}^{A}_{n}(\zeta,x)-\mathcal{L}\widehat{P}^{A}_{n}(\zeta,x)=0\quad\text{for }\ (\zeta,x)\in{\Omega}^{n}_{T_{p}}\ \ \text{with }\ \widehat{P}^{A}_{n}(\zeta,x)>(K_{p}-e^{x})^{+},\\ \partial_{x}\widehat{P}^{A}_{n}(\zeta,-n)=-e^{-n}\quad\text{and}\quad\partial_{x}\widehat{P}^{A}_{n,\varepsilon}(\zeta,n)=0\quad\text{for }\ \zeta\in[0,T_{p}],\\ \widehat{P}^{A}_{n}(0,x)=(K_{p}-e^{x})^{+}\quad\text{for }\ x\in(-n,n).\end{cases}$
Note that $\widehat{C}^{A}_{n}$ and $\widehat{P}^{A}_{n}$ are obtained by letting $\varepsilon\to 0^{+}$, where $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ converge weakly in $W^{1,2}_{p}$ for each $1<p<\infty$ in the domains $\Omega_{T_{c}}$ and $\Omega_{T_{p}}$, respectively. This yields that $\widehat{C}^{A}_{n}\in W^{1,2}_{p}(\Omega_{T_{c}})$ and $\widehat{P}^{A}_{n}\in W^{1,2}_{p}(\Omega_{T_{p}})$ for each $1<p<\infty$ and the Sobolev embedding theorem further implies that $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ are $C^{\frac{\alpha}{2},\alpha}$ functions for some $\alpha\in(0,1)$. Based on the Schauder theory for linear parabolic equations and a bootstrap argument, we deduce that the functions $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ are smooth. Thus, differentiating (4.16) and (4.17) yields that $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}$ and $\partial_{x}\widehat{P}^{A}_{n,\varepsilon}$ satisfies
$\begin{cases}\partial_{\zeta}(\partial_{x}\widehat{C}^{A}_{n,\varepsilon})-\mathcal{L}(\partial_{x}\widehat{C}^{A}_{n,\varepsilon})+\beta^{\prime}_{c,\varepsilon}(\cdots)\partial_{x}\widehat{C}^{A}_{n,\varepsilon}=\beta^{\prime}_{c,\varepsilon}(\cdots)\varphi_{\varepsilon}^{\prime}(e^{x}-K_{c})e^{x}\geqslant 0\quad\text{in }\ {\Omega}^{n}_{T_{c}},\\ \partial_{x}\widehat{C}^{A}_{n,\varepsilon}(\zeta,-n)=0\quad\text{and}\quad\partial_{x}\widehat{C}^{A}_{n,\varepsilon}(\zeta,n)=e^{n}\geqslant 0\quad\text{for }\ \zeta\in[0,T_{c}],\\ \partial_{x}\widehat{C}^{A}_{n,\varepsilon}(0,x)=\varphi^{\prime}_{\varepsilon}(e^{x}-K_{c})e^{x}\geqslant 0\quad\text{for }\ x\in(-n,n)\end{cases}$
and
$\begin{cases}\partial_{\zeta}(\partial_{x}\widehat{P}_{n,\varepsilon}^{A})-\mathcal{L}(\partial_{x}\widehat{P}_{n,\varepsilon}^{A})+\beta^{\prime}_{p,\varepsilon}(\cdots)\partial_{x}\widehat{P}_{n,\varepsilon}^{A}=-\beta^{\prime}_{p,\varepsilon}(\cdots)\varphi_{\varepsilon}^{\prime}(K_{p}-e^{x})e^{x}\leqslant 0\quad\text{in }\ {\Omega}^{n}_{T_{p}},\\ \partial_{x}\widehat{P}_{n,\varepsilon}^{A}(\zeta,-n)=-e^{-n}\leqslant 0\quad\text{and}\quad\partial_{x}\widehat{P}_{n,\varepsilon}^{A}(\zeta,n)=0\quad\text{for }\ \zeta\in[0,T_{p}],\\ \partial_{x}\widehat{P}_{n,\varepsilon}^{A}(0,x)=-\varphi^{\prime}_{\varepsilon}(K_{p}-e^{x})e^{x}\leqslant 0\quad\text{for }\ x\in(-n,n)\end{cases}$
respectively. Applying the maximum principle for each equation, we deduce $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}\geqslant 0$ in ${\Omega}^{n}_{T_{c}}$ and $\partial_{x}\widehat{P}_{n,\varepsilon}^{A}\leqslant 0$ in ${\Omega}^{n}_{T_{p}}$. Furthermore, since $\mathcal{L}e^{x}+\beta^{\prime}_{c,\varepsilon}(\cdots)e^{x}\geqslant 0$, we obtain
$\begin{cases}\partial_{\zeta}(\partial_{x}\widehat{C}^{A}_{n,\varepsilon}-e^{x})-\mathcal{L}(\partial_{x}\widehat{C}^{A}_{n,\varepsilon}-e^{x})+\beta^{\prime}_{c,\varepsilon}(\cdots)(\partial_{x}\widehat{C}^{A}_{n,\varepsilon}-e^{x})\leqslant 0\quad\text{in }\ {\Omega}^{n}_{T_{c}},\\ \partial_{x}\widehat{C}^{A}_{n,\varepsilon}(\zeta,-n)-e^{-n}=-e^{-n}\leqslant 0\quad\text{and}\quad\partial_{x}\widehat{C}^{A}_{n,\varepsilon}(\zeta,n)-e^{n}=0\quad\text{for }\ \zeta\in[0,T_{c}],\\ \widehat{C}^{A}_{n,\varepsilon}(0,x)-e^{x}=\{\varphi^{\prime}_{\varepsilon}(e^{x}-K_{c})-1\}e^{x}\leqslant 0\quad\text{for }\ x\in(-n,n)\end{cases}$
and
$\begin{cases}\partial_{\zeta}(\partial_{x}\widehat{P}_{n,\varepsilon}^{A}+e^{x})-\mathcal{L}(\partial_{x}\widehat{P}_{n,\varepsilon}^{A}+e^{x})+\beta^{\prime}_{p,\varepsilon}(\cdots)(\partial_{x}\widehat{P}_{n,\varepsilon}^{A}+e^{x})\geqslant 0\quad\text{in }\ {\Omega}^{n}_{T_{p}},\\ \partial_{x}\widehat{P}_{n,\varepsilon}^{A}(\zeta,-n)+e^{-n}=0\quad\text{and}\quad\partial_{x}\widehat{P}_{n,\varepsilon}^{A}(\zeta,n)+e^{n}=e^{n}\geqslant 0\quad\text{for }\ \zeta\in[0,T_{p}],\\ \widehat{P}_{n,\varepsilon}^{A}(0,x)+e^{x}=\{1-\varphi^{\prime}_{\varepsilon}(K_{p}-e^{x})\}e^{x}\geqslant 0\quad\text{for }\ x\in(-n,n).\end{cases}$
Therefore, by the maximum principle, we deduce $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}\leqslant e^{x}$ and $\partial_{x}\widehat{P}_{n,\varepsilon}^{A}\geqslant-e^{x}$. Note by the Sobolev embedding theorem [15, Chapter II. Lemma 3.3] that $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}$ and $\partial_{x}\widehat{P}^{A}_{n,\varepsilon}$ converge to $\partial_{x}\widehat{C}^{A}_{n}$ and $\partial_{x}\widehat{P}^{A}_{n}$ uniformly in $\Omega_{T_{c}}^{n}$ and $\Omega_{T_{p}}^{n}$, respectively, as $\varepsilon\rightarrow 0^{+}$. Hence
$0\leqslant\partial_{x}\widehat{C}^{A}_{n}\leqslant e^{x}\quad\text{in }\ \Omega_{T_{c}}^{n}\quad\text{and}\quad-e^{x}\leqslant\partial_{x}\widehat{P}^{A}_{n}\leqslant 0\quad\text{in }\ \Omega_{T_{p}}^{n}.$
Finally, by the uniqueness, passing $n\to\infty$ yields the desired inequalities for $\partial_{x}{C}^{A}$ and $\partial_{x}{P}^{A}$.
Proof of Part (iii)
For each fixed $\delta>0$, we first define $\tilde{C}_{n}^{\delta}$ by
$\tilde{C}_{n}^{\delta}(\zeta,x):=\tilde{C}_{n}^{A}(\zeta+\delta,x)$
for all $(\zeta,x)\in\Omega_{T_{c}-\delta}$. Then, we see that $\tilde{C}_{n}^{\delta}$ satisfies the following obstacle problem:
$\begin{cases}\partial_{\zeta}\tilde{C}_{n}^{\delta}(\zeta,x)-\mathcal{L}\tilde{C}_{n}^{\delta}(\zeta,x)\geqslant 0\quad\text{for}\ \ (\zeta,x)\in{\Omega}^{n}_{T_{c}-\delta}\ \ \text{with }\ \tilde{C}_{n}^{\delta}(\zeta,x)=(e^{x}-K_{c})^{+},\\ \partial_{\zeta}\tilde{C}_{n}^{\delta}(\zeta,x)-\mathcal{L}\tilde{C}_{n}^{\delta}(\zeta,x)=0\quad\text{for}\ \ (\zeta,x)\in{\Omega}^{n}_{T_{c}-\delta}\ \ \text{with }\ \tilde{C}_{n}^{\delta}(\zeta,x)>(e^{x}-K_{c})^{+},\\ \tilde{C}_{n}^{\delta}(\zeta,-n)=0\quad\text{and}\quad\tilde{C}_{n}^{\delta}(\zeta,n)=e^{n}-K_{c}\quad\text{for }\ \zeta\in[0,T_{c}-\delta],\\ \tilde{C}_{n}^{\delta}(0,x)=\tilde{C}^{A}_{n}(\delta,x)\geqslant(e^{x}-K_{c})^{+}\quad\text{for }\ x\in(-n,n).\end{cases}$
Since $\tilde{C}_{n}^{\delta}(0,x)\geqslant(e^{x}-K_{c})^{+}=\tilde{C}^{A}_{n}(0,x)$, by the comparison principle for the variational inequality, (see [25, Theorem 1.2]) we deduce
$\tilde{C}^{A}_{n}(\zeta+\delta,x)=\tilde{C}_{n}^{\delta}(\zeta,x)\geqslant\tilde{C}^{A}_{n}(\zeta,x)\quad\text{for all }\ (\zeta,x)\in\Omega^{n}_{T_{c}-\delta}.$
Take $n\rightarrow\infty$ and the uniqueness result yields that
${C}^{A}(\zeta+\delta,x)\geqslant{C}^{A}(\zeta,x)\quad\text{for all }\ (\zeta,x)\in\Omega_{T_{c}-\delta}.$
Observe that ${C}^{A}$ is differentiable almost everywhere with respect to $\tau$ in $\Omega_{T_{c}}$, and that $\partial_{\tau}C^{A}\geqslant 0$ almost everywhere. In particular, this condition satisfies the hypothesis of [3, Theorem 1.1]. Hence, $\partial_{\tau}C^{A}$ is continuous in $\Omega_{T_{c}}$, and consequently, $\partial_{\tau}C^{A}\geqslant 0$ everywhere in $\Omega_{T_{c}}$.
The same argument applies to $P^{A}$. By following the same procedure, it can be obtained that $\partial_{\tau}P^{A}\geqslant 0$ almost everywhere in $\Omega_{T_{p}}$. Theorem 1.1 in [3] yields that $\partial_{\tau}P^{A}$ is continuous, and hence nonnegative everywhere in $\Omega_{T_{p}}$.
Proof of Part (iv)
The proof for Part (iv) can be found in [21]. ∎
D. Conditions of the operator $\mathcal{F}$ in Theorem 3.1
Let us prove that the operator $\mathcal{F}:\mathcal{A}\to\mathcal{X}$ in the proof of Theorem 3.1 satisfies
  • (1)
    $\mathcal{F}(\mathcal{A})\subset\mathcal{A}$,
  • (2)
    $\mathcal{F}$ is continuous,
  • (3)
    $\mathcal{F}(\mathcal{A})$ is precompact in $\mathcal{X}$.
(1) Fix any $w\in\mathcal{A}$ and consider $u=\mathcal{F}(w).$ Note that $J_{\varepsilon}(\tau,x)\geqslant 0,$ $\beta_{\varepsilon}(\cdots)\geqslant 0$ and $\nu\cdot D_{x}u(\tau,x)\leqslant 0$ for each $(\tau,x)$ at the lateral boundary and the inward pointing normal vector $\nu$ at $(\tau,x).$ Then by the comparison principle [17, 13p], we have $u\geqslant 0.$
(2) To obtain the continuity of $\mathcal{F},$ it suffices to prove the following:
$\text{Let}\ \{w_{j}\}\$$\text{be a sequence in}\ \mathcal{A}\ \ \text{such that}\ \displaystyle{\lim_{j\to\infty}w_{j}}=w.$
$\text{Then}\ \displaystyle{\lim_{j\to\infty}\mathcal{F}(w_{j})}=\mathcal{F}(w).$
Let $u_{j}:=\mathcal{F}(w_{j})\ \text{for each }j\in\mathbb{N}\ \text{and}\ u:=\mathcal{F}(w).$ By subtracting the two equations that $u$ and $u_{j}$ satisfy respectively, we have
$\begin{cases}\partial_{\tau}(u-u_{j})-\mathcal{L}(u-u_{j})=-\beta_{\varepsilon}^{\prime}(\cdots)\ (w-w_{j})\quad\text{in }\ \Omega_{T}^{n},\\ \partial_{x}(u-u_{j})(\tau,-n)=0\quad\text{and}\quad\partial_{x}(u-u_{j})(\tau,n)=0\quad\text{for }\ \tau\in[0,T],\\ (u-u_{j})(0,x)=0\quad\text{for }\ x\in(-n,n).\end{cases}$
Then the $W^{1,2}_{p}$-estimate combined with the embedding theorem implies that
$\|u-u_{j}\|_{C(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}\|\beta_{\varepsilon}^{\prime}(\cdots)(w-w_{j})\|_{L^{\infty}(\Omega_{T}^{n})}$
for some constant $\mathcal{K}>0$. Therefore, it follows that ${\lim_{j\to\infty}u_{j}}=u.$
(3) Let $\{u_{j}\}_{j=1}^{\infty}$ be a sequence in the closure of $\mathcal{F}(\mathcal{A})$ with respect to the $\|\cdot\|_{C(\overline{\Omega_{T}^{n}})}$ norm.
Case 1. For each $u_{j}\in\mathcal{F}(\mathcal{A})$, there exists a sequence of functions $\{w_{j}\}_{j=1}^{\infty}$ such that $\mathcal{F}(w_{j})=u_{j}$. By applying the $W^{1,2}_{p}$ estimate on $u_{j}$, we obtain
$\|u_{j}\|_{W^{1,2}_{p}(\Omega^{n}_{T})}$ $\leqslant\mathcal{K}\big(\|J_{\varepsilon}\|_{W^{2}_{p}((-n,n))}+\|\beta_{\varepsilon}(w_{j}-J_{\varepsilon})\|_{L^{p}(\Omega^{n}_{T})}+(e^{n}+e^{-n})T\big)$
$\leqslant\mathcal{K}\big(\|J_{\varepsilon}\|_{W^{2}_{p}((-n,n))}+\|\beta_{\varepsilon}(-J_{\varepsilon})\|_{L^{p}(\Omega^{n}_{T})}+(e^{n}+e^{-n})T\big)$
for some constant $\mathcal{K}>0$ that does not depend on $j\in\mathbb{N}$ and $\alpha\in(0,1)$. Note that the last inequality is due to the monotonicity of $\beta_{\varepsilon}$ and $w\geqslant 0$. Thus, combined with the embedding theorem it can be seen that
$\|u_{j}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}^{\prime}$
for some constant $\mathcal{K}^{\prime}$ that does not depend on $j\in\mathbb{N}$ and $\alpha\in(0,1)$.
Case 2. For each fixed $u_{i}\notin\mathcal{F}(\mathcal{A})$, there exists a sequence $\{u^{(i)}_{l}\}_{l=1}^{\infty}\subset\mathcal{F}(\mathcal{A})$ such that
$u^{(i)}_{l}\rightarrow u_{i}\quad\text{in }\ C(\overline{\Omega_{T}^{n}})\ \text{ as }\ l\rightarrow\infty.$
Hence, there exists $\{w^{(i)}_{l}\}_{l=1}^{\infty}$ in $\mathcal{A}$ such that $\mathcal{F}(w^{(i)}_{l})=u^{(i)}_{l}$. Again, by the $W^{1,2}_{p}$ estimate, the embedding theorem, and the monotonicity of $\beta_{\varepsilon}$, we deduce $\|u_{l}^{(i)}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}^{\prime}$ for some $\alpha\in(0,1)$. Take $l\rightarrow\infty$ to get $\|u_{i}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}^{\prime}$.
Combining both cases, we conclude
$\|u_{j}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}^{\prime}$
for all $j\in\mathbb{N}$.
By the Arzela-Ascoli theorem, there exists a subsequence $\{u_{j_{k}}\}_{k=1}^{\infty}\subset\{u_{j}\}_{j=1}^{\infty}$ and $u\in C(\overline{\Omega_{T}^{n}})$ such that
$u_{j_{k}}\rightarrow u\quad\text{in }\ C(\overline{\Omega_{T}^{n}})\quad\text{as }\ k\rightarrow\infty.$
Observe that $u$ is contained in the closure of $\mathcal{F}(\mathcal{A})$ by the assumption. This proves that $\mathcal{F}(\mathcal{A})$ is precompact in $\mathcal{X}=C(\overline{\Omega_{T}^{n}}).$
Acknowledgment
J. Jeon was supported by NRF grant funded by MSIT (RS-2023-00212648). J.Ok was supported by NRF grant funded by MSIT (NRF-2022R1C1C1004523).